### v1

# LVLM-COUNT: Enhancing the Counting Ability of Large Vision-Language Models

Abstract
Counting is a fundamental skill for various visual tasks in real-life applications, requiring both object recognition and robust counting capabilities. Despite their advanced visual perception, large vision-language models (LVLMs) struggle with counting tasks, especially when the number of objects exceeds those commonly encountered during training. We enhance LVLMs’ counting abilities using a divide-and-conquer approach, breaking counting problems into sub-counting tasks. Unlike prior methods, which do not generalize well to counting datasets on which they have not been trained, our method performs well on new datasets without any additional training or fine-tuning. We demonstrate that our approach enhances counting capabilities across various datasets and benchmarks.
1 Introduction
Counting is a key cognitive task with broad applications in industry, healthcare, and environmental monitoring (De Almeida et al., 2015; Guerrero-Gómez-Olmedo et al., 2015; Paul Cohen et al., 2017; Lempitsky and Zisserman, 2010). It improves manufacturing, inventory, and quality control, ensures safety in medical settings, and helps manage resources in environmental efforts (Wang and Wang, 2011; Zen et al., 2012; Arteta et al., 2016). Recent advancements in prompt-based models enable counting of unlimited object varieties without visual exemplars. Although current trained, and text-prompt-based, counting models by Dai et al. (2024); Amini-Naieni et al. (2024) perform well on the datasets they are trained on, they face the following challenges. First, they require fine-tuning on new datasets. Second, since the concepts in the counting datasets are limited, these models do not generalize well to counting questions that involve complex reasoning. There are also training-free models by Shi et al. (2024), but overall they have weaker performance compared to trained models. On the other hand, large vision-language models (LVLMs), such as GPT-4o (Achiam et al., 2023), which also do not need dataset specific training, show very good performance in counting low numbers of objects, usually less than 20, across most datasets. However, their performance deteriorates for larger numbers of objects, regardless of the dataset.
We enhance the accuracy of LVLMs to count objects in images by leveraging their reasoning power within a divide-and-conquer method. The LVLMs allow us to handle diverse objects and complex counting questions, while our divide-and-conquer method alleviates challenges associated with counting large numbers of objects. Inspired by prior work on the rapid and accurate estimation of small quantities by Chattopadhyay et al. (2017), we divide an image into sub-images, and prompt the LVLM to count the objects of interest in each sub-image. The counts from each sub-image are then aggregated to make the final prediction. Our proposed workflow is illustrated in Figure 1.
Figure 1: Illustration of our proposed pipeline. First, the area of interest, such as a set of zebras, is detected by extracting the object of interest ($E$) from a prompt question ($Q$) using a large language model (LLM) which is the same as LVLM in our work. Then, $E$ and the image are provided as input to an area detection model, such as the one by Liu et al. (2023). Second, any objects corresponding to $E$ are segmented. Third, in the object-aware division step, we use the segmentation masks to divide the areas of objects of interest without cutting through the objects. Finally, the number of objects of interest in each sub-image is computed using an LVLM, and the results are aggregated.
Initially, in our pipeline, the object of interest is extracted from the input question using an LLM. The area containing the object of interest is detected in the image by a grounding model, such as Liu et al. (2023), and then cropped. The cropping step is important since it removes irrelevant context from the image. Secondly, using an object detection model by Liu et al. (2023), and a segmentation model by Kirillov et al. (2023), the segmentation masks of the objects of interest are created. Thirdly, we use a mechanism that divides the image into multiple sub-images without cutting through the objects of interest. We call this mechanism object-aware division. The division positions can be either set manually or determined automatically using unsupervised and non-parametric methods. In the latter case, we automatically select the division positions based on object masks, then we treat the object-aware division as a path-finding problem, avoiding objects as obstacles. A black-white image is built by converting all the masks into black and the rest of the image into white pixels. The binary image is converted into a graph where only white pixels are connected as nodes. Using the $A^{*}$ algorithm (Russell and Norvig, 2016), a path is found from one end to the other end of the image, ensuring objects remain intact. Finally, using an LVLM as a counting tool, the objects of interest in the sub-images are counted and aggregated. Our contributions are summarized below:
  1. 1.
    We propose LVLM-Count that leverages the strengths of large vision-language models (LVLMs) in visual perception for counting. Our method is a prompt-based counting approach that, in addition to simple counting problems, can handle complex cases as well. Our method does not require any additional training or fine-tuning on any dataset. LVLM-Count outperforms existing state-of-the-art models across various datasets and benchmarks.
  2. 2.
    We propose a solution for object-aware division. Accurate division is crucial, as parts of cut objects can lead to over-counting (see Figure 2(a)). To the best of our knowledge, this is the first method to divide images without cutting through objects of interest.
(a) Naive division
(b) Object-aware division
(c) Counting error
Figure 2: Comparison of the naive and the object-aware division. The objects of interest are the circles. In Figure 2(a), we illustrate a naive division of the input image, which is divided into equally sized sub-images with straight lines. In Figure 2(b), we illustrate the object-aware division, which avoids cutting through circles. In Figure 2(c), we illustrate the counting error of GPT-4o for images with randomly positioned circles. The absolute counting error is the absolute difference between the ground truth and the number predicted by GPT-4o. The results are averaged over three trials.
As a minor contribution, we create a new benchmark to address two drawbacks of existing benchmarks. Prior datasets either feature simple counting tasks, e.g., counting “strawberries”, or include complex questions with small numbers of objects. To address both issues, we develop a challenging benchmark for counting emoji icons. The subtle variations within emoji classes make this benchmark uniquely difficult. Since none of the models have been exposed to it, this benchmark serves as a fair test of the performance of counting methods for various attributes and complex concepts.
2 Related Work
Early counting models, referred to as class-specific, targeted counting problems for certain categories (Arteta et al., 2016; Babu Sam et al., 2022; Mundhenk et al., 2016; Xie et al., 2018), such as cars, people, or cells. Later, with the emergence of stronger vision models and large-scale datasets, class-agnostic methods were proposed that could count objects from a wide variety of categories. However, most existing class-agnostic, or open-world, models require visual exemplars of the target objects (Đukić et al., 2023; Gong et al., 2022; Lin et al., 2022; Liu et al., 2022; Lu et al., 2019; Nguyen et al., 2022; Ranjan et al., 2021; Shi et al., 2022; Yang et al., 2021; You et al., 2023). The concept of divide and conquer has also been used in early work(Xiong et al., 2019; Chattopadhyay et al., 2017). However, this early work requires training dedicated models to utilize this concept.
Text-based counting-specific trained models. With the advent of vision-language foundation models such as CLIP and GroundingDINO, text-based open-world methods have been proposed. Leveraging the rich textual and visual feature extraction capabilities of foundation models, obtained through web-scale training, the text-based counting methods by Amini-Naieni et al. (2023); Dai et al. (2024); Kang et al. (2024); Amini-Naieni et al. (2024) have started to demonstrate comparable or superior accuracy. GroundingREC by Dai et al. (2024) is an open-world model built on top of GroundingDINO (Liu et al., 2023), and introduces an additional task called referring expression counting. Concurrently, Amini-Naieni et al. (2024) proposed a method that also builds on GroundingDINO but adds an extra image-text fusion module in the input, enabling the model to accept text and/or visual exemplars to determine the target.
Models without counting-specific training. Shi et al. (2024) introduce TFOC, a counting model that does not require any counting-specific training. Instead, they cast the counting problem as a prompt-based segmentation task, using SAM (Kirillov et al., 2023) to obtain segmentation maps that determine the output number. Another group of models that do not need further training to count are LVLMs. State-of-the-art models such as GPT-4o (Achiam et al., 2023), Gemini 1.5 Pro (Reid et al., 2024), and Claude 3.5 Sonnet (Anthropic, 2024) show strong performance in counting small numbers of objects, although their performance degrades with larger numbers of objects. In comparison to prior work, our method (LVLM-Count) is an open-world, text-based counting approach that does not require any counting-specific training. Although we use pre-trained GroundingDINO and SAM in our method, we differ from prior work by not treating the counting task as a segmentation or detection problem. Instead, we use LVLMs as a counting tool.
3 LVLM-Count
Our proposed method aims to answer counting questions by dividing an image into sub-images while avoiding cuts through objects of interest. LVLM-Count consists of four key stages. First, in the “Area Detection” stage, we localize areas containing relevant objects. Second, in the “Target Segmentation” stage, we identify and segment these local areas. Third, in the “Object-aware Division” stage, we divide the localized areas into sub-images without cutting through the segmented objects. Finally, the LVLM counts the target objects in each sub-image and aggregates the results. Figure 1 illustrates the workflow of our method, which we will detail in the following subsections. Note that in this section, we illustrate each stage using example images. These images are are used solely for illustration purposes. LVLM-Count may not outperform other methods on these examples.
3.1 Area Detection
In this part of the pipeline, we assume that we are given a counting question $Q$ along with an image. The question $Q$ contains an expression $E$ that specifies a set of objects of interest. The expression $E$ distinguishes these objects from objects of other categories or the same category but with different attributes present in the image. By employing an LLM, the expression $E$ is extracted from $Q$. For example, let $Q$ be “How many people are in the boat?”. $Q$ is given to an LLM, which is prompted to return the expression $E$ “people in the boat”, referring to the objects we want to count. After $E$ is extracted, it is given as input to GroundingDINO along with the image. The output of GroundingDINO is a set of bounding boxes that have relevance to $E$ beyond a certain threshold. These bounding boxes often overlap and typically contain repeated objects. Thus, all the overlapping output bounding boxes are merged. After merging, a set of non-overlapping areas of interest may remain. We consider the non-overlapping areas as “detected areas”, which are then cropped to be passed to the next stage. Note that the area detection stage is important as it extracts the area with the most relevant context for the counting question. This process is illustrated in Figure 3.
Figure 3: Illustration of the area detection step of LVLM-Count. For this image, $Q$ is set to “How many people are in the boat”. The LLM that is used in this step returns an $E$ which is “people in the boat”. $E$ and the original image are given as input to GroundingDINO, which returns some bounding boxes (left and upper right images) that are merged to form the final detected area.
3.2 Target Segmentation
The cropped images from the first stage contain objects of interest, and the ultimate goal is to divide them without cutting through those objects. However, a prerequisite for implementing such a mechanism is to first detect and localize the objects of interest. Each cropped image is fed into an open-world detection model along with $E$. The output of the open-world detection model produces a bounding box for each object of interest. The bounding boxes are then given as input to a segmentation model, which returns segmentation masks for the objects within each bounding box. We illustrate an example of this process in Figure 4.
(a) Original image
(b) GroundingDINO output
(c) SAM output
Figure 4: Illustration of the target segmentation step of LVLM-Count. The goal is to produce all the instance masks for $E$ set to “brown egg”. Figure 4(a), together with $E$, is given as input to GroundingDINO, which produces the output shown in Figure 4(b). Figure 4(b) is then given as input to SAM, which produces the output shown in Figure 4(c).
How to determine the bounding boxes. To determine the bounding boxes, we use GroundingDINO and set the bounding box probability threshold to a low value to avoid missing any objects. The bounding boxes alone cannot help with the object-aware division of the cropped images due to their rigid structure, which includes redundant areas in the vicinity of the object and, in the worst case, overlaps with other bounding boxes. Our goal is to precisely locate the pixels of an object of interest.
How to determine the segmentation masks. We use a pre-trained segmentation model, i.e., SAM (Kirillov et al., 2023), for the segmentation task, which accepts bounding boxes as prompts. It produces a mask covering the most prominent object within a bounding box. We run non-maximum suppression on the masks produced by SAM to remove the masks corresponding to less certain bounding boxes that have large overlaps with other masks. The masks produced for each cropped image are then passed to the next stage. For GroundingDINO we set the bounding box threshold to a very low value to avoid missing any objects of interest, especially if the objects belong to a rare category that GroundingDINO has encountered less frequently during pre-training. As shown in Figure 4, a low threshold value can lead to false positives, making the number of masks unreliable for the counting problem. Nonetheless, the false positive masks have no negative effect on the object-aware division in the next stage. The only consequence is that, in addition to the target masks, the division paths will not cut through the false positive masks either. In other words, extra masks are of no concern as long as all the target masks are identified.
3.3 Object-aware Division
In this stage, the cropped image is divided into appropriate sub-images so that no object of interest is cut by the dividing paths. The core idea is that the dividing paths should not intersect the pixels covered by the masks corresponding to the objects of interest. This step consists of two sub-steps. First, we decide the starting and ending points of the paths. Second, we draw the paths. Below, we describe how we approach these two sub-steps.
How to determine the starting and ending points of the paths. We consider the first sub-step as a hyper-parameter, which can be set in various ways depending on the type of dataset. We utilize two approaches in our experiments. The first approach is unsupervised and non-parametric, where we use a clustering method to obtain the start and end points of the paths. We describe the clustering approach in more detail below. The second approach is to simply use a pre-determined number of equidistant points. This is helpful when we know a priori that objects are uniformly distributed across the image. In this case, we fix $k$ equidistant vertical and horizontal coordinates in the image as the points of the dividing paths for the entire dataset. In our experiments, we specify which approach we use for each dataset. Nonetheless, we will illustrate in the experimental results that even if we do not treat this as a hyperparameter and always use the unsupervised and non-parametric approach, LVLM-Count is quite effective.
Figure 5: Illustration of the unsupervised and non-parametric method to obtain the division points ($P^{1}_{s},P^{1}_{e}$), and ($P^{2}_{s},P^{2}_{e}$,). The object of interest is “horse”. Firstly, the target masks are produced. Then, a few pixels are sampled (shown as points inside the segmented objects) from the pixels composing each mask. The samples are projected onto the $x$-axis. The projected points are clustered using mean-shift clustering. The point in the middle of two consecutive clusters is considered a vertical division point. The straight vertical lines are obtained using our path-finding method.
Description of the unsuperivsed and non-parametric approach. A few pixels are sampled from each of the masks. To determine the points of the vertical division paths, the samples taken from the masks are projected onto the $x$-axis. The projected points are automatically clustered using a non-parametric mean-shift algorithm. Once the clusters are identified, the point between the point with the highest $x$ value in one cluster and the point with the lowest $x$ value in the subsequent cluster is considered the $x$-coordinate of a vertical division path. Using this technique, we obtain the appropriate coordinates for the vertical paths, as well as the number of paths. For example, if there is only one cluster, no vertical division is required, and if there are two clusters, one vertical path will divide the image into two parts. We illustrate this approach in Figure 5. Note that we chose the vertical axis purely for convenience in illustration. The same process can be applied along the $y$-axis to obtain the $y$-coordinates of the horizontal division paths.
How to draw the paths. In effect, knowing the $x$-coordinate of a vertical path means that the coordinates of its endpoints are known. In particular, assuming height $h$ for an image crop, we consider $P_{s}=(x,0)$ and $P_{e}=(x,h)$ as the start and end points, respectively. In an ideal case where there are no masks in the path of a straight line connecting the two points, this line will be drawn by connecting all the pixels on the straight path. However, there are potential masks that can be considered obstacles blocking the path. In other words, beginning from $P_{s}$, the line needs to go around these obstacles to reach $P_{e}$. Consequently, we treat this as a 2-dimensional path-finding problem. To solve the problem, we build a 2D binary map, $I_{B}$, where the pixels covered by the masks are turned into black, indicating them as obstacles, and all the other pixels are turned into white, showing they are open for passage. This binary image $I_{B}$ is mapped into a graph $G$, where each white pixel is a node, and it is connected to all of its white neighboring pixels. We use the $A^{*}(G,P_{s},P_{e},g)$ search algorithm to find a path that connects $P_{s}$ to $P_{e}$, where the heuristic $g$ is set to be Manhattan distance. The output of $A^{*}$ is a set of connected pixels that go around the obstacles and connect $P_{s}$ to $P_{e}$, creating an object-aware division path, as shown in Figure 6. The path-finding algorithm is run for all division coordinates. Finally, we draw the image contours based on these division paths and take the area surrounded by each contour as a resulting sub-image.
Figure 6: Illustration of object-aware division. The masks for “donut” are obtained and turned into a black-and-white image. A dividing path is found by connecting $P_{s}$ to $P_{e}$ using the $A^{*}$ search algorithm in a graph that corresponds to the binary image, where the only nodes in the graph are white pixels, which are fully connected. The path is mapped back to the pixel domain.
3.4 Target Counting
At this point, all the sub-images obtained from the cropped areas are gathered. Since these are small partitions of the original image, they might lack the desired resolution. Thus, using a super-resolution model can alleviate this problem for the sub-images. In this work, we use Real-ESRGAN (Wang et al., 2021). Then, question $Q$ and each sub-image are given as input to an LVLM. At the end of the loop, the recorded numbers for the sub-images are aggregated to form the final answer.  [^1] Note that using super-resolution before counting the objects in the sub-images is also the default behavior of our method. For a specific dataset where we observe that the resolution is preserved after division, we deactivate super-resolution, as the cost and usage of LVLM APIs is much higher. Finally, for images with a very large number of objects, sometimes LVLMs refuse to count, citing the large number. In those cases, an “estimate” prompt is given instead of a “count” prompt..
4 Experiments
In this section, we present the performance results of our method on a counting-specific dataset, an open-ended counting benchmark with two types of questions, simple and complex, and a challenging counting benchmark that we propose using emoji icons. We compare the results to state-of-the-art models which have been specifically trained on counting datasets, and we also compare to state-of-the-art models which have not been trained on counting datasets. The code to reproduce the experiments can be found here: https://github.com/mrghofrani/lvlm-count.
4.1 Datasets and benchmarks
FSC-147 (Ranjan et al., 2021). FSC-147 is a counting dataset that contains $6135$ images, spanning $147$ different object categories such as kitchen utensils, office supplies, vehicles, and animals. The number of objects in each image ranges from $7$ to $3731$, with an average of $56$ objects per image. The dataset is split into training, validation, and test sets. A total of $89$ object categories are assigned to the training set, $29$ to the validation set, and $29$ to the test set, with different categories in each split. The training set contains $3659$ images, with the validation and test sets containing $1286$ and $1190$ images, respectively. For each image in the test set, a single category name is given, and the expected output is the number of instances from the category.
Open-ended Counting Benchmark. TallyQA (Acharya et al., 2019) is an open-ended counting dataset that includes complex counting questions involving relationships between objects, attribute identification, reasoning, and more. TallyQA is quite a large dataset, with the train set having $249,318$ questions and the test set having $22,991$ simple and $22,991$ complex counting questions. Please refer to Appendix C and Figure 9 for more information on simple and complex categorization of the questions. The number of objects in each image ranges from 00 to $15$. The dataset has a heavy bias towards a low number of objects (see Figure 8 in the Appendix). To alleviate this bias and create a benchmark for efficiently measuring the simple and complex open-ended capabilities of a counting model, we randomly sample $10$ questions per ground truth count. This results in $155$ simple and $149$ complex open-ended counting questions in total. It is important to note that the bias in TallyQA is pronounced; for most ground truth values greater than $10$, there are fewer than $10$ samples available in the entire test set.
Emoji-Count. Although TallyQA addresses the scarcity of complex counting questions in prior datasets to a certain degree, the range of target objects is limited, spanning from 00 to $15$. To our knowledge, no counting benchmark exists for large numbers involving complex reasoning. To this end, we propose a challenging counting benchmark using emoji icons. From the $1816$ standard emoji icons, we remove those that directly overlap with concepts demonstrated by other icons. We then group the remaining $1197$ icons into $82$ classes. In each class, there are icons from the same or similar object categories, but with subtle differences that require complex reasoning to distinguish. For each of the $82$ classes, an empty $1024\times 1024$ image is first created. This image is filled with six categories chosen randomly from the class, with each category having a random count between $30$ and $50$ in the image. We illustrate an example of this dataset in Figure 7.
(a) Q: How many waning gibbous moons are there in the image? Answer: 13.
(b) Q: How many clocks at time “two-thirty” are there in the image? Answer: 15.
Figure 7: Illustration of two challenging cases from Emoji-Count. Note that, for convenience in visualization, a smaller version of the real images is depicted here. The real images are $1024\times 1024$ and contain many more icons. In Figure 7(a), the class name is “Moon Phase”. To predict the correct answers for the question, the model needs accurate information about the different moon phases in addition to strong counting capability. In Figure 7(b), the class name is “Clock Time”. In this case, the model needs to be able to read the time shown on the clocks in addition to strong counting capability.
4.2 Results
The following discusses the numerical results of our experiments with LVLM-Count on each benchmark described in Section 4.1. For visual examples of LVLM-Count’s performance on each benchmark, see Appendix M. Additionally, for the ablation study and experiments on the PASCAL VOC dataset (Everingham et al., 2015), see Appendix A and Appendix K, respectively.
FSC-147. We compare the performance of LVLM-Count to an extensive list of state-of-the-art counting models on the test set of the FSC-147 dataset. We also include two baselines: i) taking the number of target segmentation masks as the final answer, and ii) giving the output of the target segmentation stage to base GPT-4o and asking it to count the masks. We run different experiments using GPT-4o, Gemini 1.5 Pro, and an open-source model Qwen2 VL 72B AWQ (Yang et al., 2024) as LVLMs. Since objects in the FSC-147 dataset are uniformly distributed across scenes, we deactivate the cluster-based method for finding division points and instead use two vertical and two horizontal object-aware division paths with equidistant start and end points. We also deactivate the super-resolution. We further run an experiment with GPT-4o as the LVLM that uses the cluster-based approach along both horizontal and vertical axes to automatically determine the start and end points of the division paths. This is done to measure the performance of our method when prior knowledge is not used to set the start and end points. The rest of the process follows the procedure described in Section 3. The expression $E$ used in different stages of our method is the category name provided in the test set. A simple $Q$ in the form of “How many $E$ are there? If you don’t see any, say zero.” is built and given as a text prompt to the LVLM during the counting stage. The results are shown in Table 1. For this dataset, the mean absolute error (MAE) and root mean square error (RMSE) are reported. We observe that our method enhances the performance of all three LVLMs in terms of MAE. With our method, all three LVLMs outperform TFOC, which which is a training-free method. Interestingly, although the base Qwen2 VL 72B AWQ is not as powerful as its commercial counterpart, GPT-4o, it performs almost on par with GPT-4o when both use our pipeline. The results also show that even without leveraging prior information to set the start and end points of the division paths, our pipeline remains effective.
Table 1: Evaluation of state-of-the-art models on the test set of the FSC-147 dataset. The column “Trained Model” indicates if a model has been trained on FSC-147. In all tables, the results for base LVLMs and LVLM-Count are reported over three trials. Columns marked with $\Delta$ show the improvement that LVLM-Count brings over the base LVLM that it uses. For accuracy metrics, refer to Table 9. Additionally, MAE analysis across different intervals of ground truth values is provided in Appendix H.
Method Trained Model MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC (Shi et al., 2024) 24.79 - 137.15 - VLCounter (Kang et al., 2024) 17.05 - 106.16 - CounTX(Amini-Naieni et al., 2023) 15.88 - 106.29 - $\text{DAVE}_{\text{prm}}$ (Pelhan et al., 2024) 14.90 - 103.42 - CountGD (Amini-Naieni et al., 2024) 14.76 - 120.42 - GroundingREC (Dai et al., 2024) 10.12 - 107.19 - Number of target segmentaion masks 44.14 - 154.39 - Base GPT-4o counting the target segmentaion masks 38.45 - 38.45 - GPT-4o 23.75 - 137.39 - LVLM-Count (GPT-4o as LVLM, Cluster-based on both axes) 17.67 $\downarrow 6.08$ 90.61 $\downarrow 46.78$ LVLM-Count (GPT-4o as LVLM) 16.23 $\downarrow 7.52$ 76.09 $\downarrow 61.3$ Gemini 1.5 Pro 25.20 - 108.76 - LVLM-Count (Gemini 1.5 Pro as LVLM) 14.85 $\downarrow 10.35$ 85.60 $\downarrow 23.16$ Qwen2 VL 72B AWQ 33.45 - 145.90 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 17.77 $\downarrow 15.68$ 113.06 $\downarrow 32.84$
Table 1 shows that, although we outperform models that have not been trained on FSC-147, the best-performing models are those that have been trained on this dataset. However, in subsequent experiments, we demonstrate that, while the best-performing models on FSC-147 show acceptable generalization on benchmarks they have not been trained on, they lack the generalization ability of our method on new datasets, and they are even outperformed by the base LVLM models.
Open-ended Counting Benchmark. We evaluate the performance of LVLM-Count on the simple and complex questions in the open-ended counting benchmark mentioned above. For LVLM-Count, we compute the start and end points of the object-aware division paths using the unsupervised and parameter-free clustering approach described in Section 3.3 on $x$-axis. This approach is necessary because the images in this benchmark are diverse, and objects are not uniformly distributed across the image. Thus, using a fixed number of object-aware division paths with equidistant start and end points does not perform well. We also conduct an experiment using GPT-4o as the LVLM, where we use the clustering method along both axes to find the start and end points.
Since these are special cases of VQA tasks, we report exact accuracy (EA), which is the standard performance metric for VQA, in addition to MAE and RMSE. We compare our method against the base GPT-4o, Gemini 1.5 Pro, and Qwen2 VL 72B models, as well as three of the best-performing trained models from Table 1, namely GroundingREC, CountGD, and $\text{DAVE}_{\text{prm}}$ [^2] GroundingREC in Table 2 does not use the same model weights as those trained for FSC-147. It was trained on the REC-8K dataset (Dai et al., 2024), which is specifically designed for referring expression counting, making it stronger for complex counting questions. CountGD and $\text{DAVE}_{\text{prm}}$ use the same model weights as those used for FSC-147, as the authors have not provided model weights trained on a dataset with referring expressions., and the only prior training-free model, TFOC. Note that due to extremely poor performance, we do not provide the original question to these models. Instead, we assist them with the $E$ extracted in our pipeline. The results are shown in Table 2. Our method improves both MAE and EA over the base GPT-4o model on both complex and simple benchmarks.
In general, however, there is barely any improvement over base LVLMs on simple questions. This is because the questions are straightforward, and the ground truths are within [0, 15]. Based on our observations from Figure 2(c) and Figure 18, we expect the base LVLMs to perform better in such cases. However, for complex questions, there is consistent improvement in EA across all three LVLMs. Interestingly, despite the considerable gap between the EA of the base Qwen2 VL 72B AWQ and base GPT-4o on complex questions, using LVLM-Count enables this open-source model to achieve a higher EA than the base GPT-4o. Moreover, observe that GroundingREC, CountGD, and $\text{DAVE}_{\text{prm}}$ are outperformed by the LVLM-based models on both simple and complex categories. This is because they have not been specifically trained on TallyQA. The performance gap is more pronounced on complex questions.
Table 2: Evaluation of models on the TallyQA benchmark. Numbers in parentheses under the $\Delta$ signs show the performance difference between LVLM-Count and the base LVLM it uses. Green indicates improvement, while red represents degradation. For more details about the simple and complex types of questions, please refer to Appendix C and Figure 9. For additional accuracy measures, refer to Table 10 and 11.
Method Simple Questions Complex Questions EA (%$\uparrow$($\Delta$) MAE $\downarrow$($\Delta$) RMSE $\downarrow$($\Delta$) EA (%$\uparrow$($\Delta$) MAE $\downarrow$($\Delta$) RMSE $\downarrow$($\Delta$) TFOC (Shi et al., 2024) 6.45 (-) 8.28 (-) 14.79 (-) 1.34 (-) 12.41 (-) 22.80 (-) $\text{DAVE}_{\text{prm}}$ (Pelhan et al., 2024) 5.81 (-) 16.06 (-) 35.26 (-) 3.36 (-) 28.36 (-) 57.56 (-) GroundingREC (Dai et al., 2024) 23.87 (-) 2.90 (-) 4.71 (-) 16.78 (-) 5.83 (-) 10.13 (-) CountGD (Amini-Naieni et al., 2024) 41.94 (-) 2.37 (-) 4.56 (-) 5.37 (-) 9.78 (-) 17.21 (-) Number of the target segmentation masks 27.31 (-) 2.60 (-) 4.10 (-) 15.44 (-) 4.25 (-) 6.78 (-) Base GPT-4o counting the target segmentation masks 25.38 (-) 2.82 (-) 2.82 (-) 11.41 (-) 4.99 (-) 4.99 (-) GPT-4o 44.30 (-) 1.38 (-) 2.35 (-) 29.08 (-) 2.60 (-) 4.74 (-) LVLM-Count (GPT-4o as LVLM, Cluster-based on both axes) 41.29 ($\downarrow 3.01$) 1.82 ($\uparrow 0.44$) 3.33 ($\uparrow 0.98$) 28.41 ($\downarrow 0.67$) 3.18 ($\uparrow 0.58$) 8.91 ($\uparrow 4.17$) LVLM-Count (GPT-4o as LVLM, Cluster-based on $x$-axis) 44.73 ($\uparrow 0.43$) 1.18 ($\downarrow 0.20$) 2.06 ($\downarrow 0.29$) 34.68 ($\uparrow 5.6$) 2.28 ($\downarrow 0.32$) 4.18 ($\downarrow 0.56$) Gemini 1.5 Pro 47.10 (-) 1.08 (-) 1.87 (-) 25.73 (-) 2.13 (-) 3.36 (-) LVLM-Count (Gemini 1.5 Pro as LVLM, Cluster-based on $x$-axis) 45.16 ($\downarrow 1.94$) 1.43 ($\uparrow 0.35$) 4.72 ($\uparrow 2.85$) 26.62 ($\uparrow 0.89$) 2.79 ($\uparrow 0.66$) 4.70 ($\uparrow 1.34$) Qwen2 VL 72B AWQ 49.03 (-) 1.44 (-) 2.74 (-) 24.61 (-) 3.21 (-) 5.35 (-) LVLM-Count (Qwen2 VL 72B AWQ as LVLM, Cluster-based on $x$-axis) 41.72 ($\downarrow 7.31$) 1.66 ($\uparrow 0.22$) 3.41 ($\uparrow 0.67$) 30.65 ($\uparrow 6.04$) 2.47 ($\downarrow 0.74$) 4.35 ($\downarrow 1$)
Emoji-Count. We evaluate the performance of LVLM-Count on the Emoji-Count benchmark. The results are shown in Table 3. For LVLM-Count, we deactivate the cluster-based division points and, similar to the FSC-147 experiment, use two vertical and horizontal object-aware division paths with equidistant start and end points. We also run an experiment with GPT-4o where LVLM-Count uses the clustering approach along both axes to determine the start and end points of the division paths. We report MAE and RMSE and compare against the base LVLMs, TFOC, $\text{DAVE}_{\text{prm}}$, CountGD, and GroundingREC. Note that none of the models have been exposed to this benchmark. Furthermore, it is a challenging benchmark as it requires understanding complex concepts. We observe that prior counting models perform poorly because, for any object of interest in the image, these models tend to count all the objects and cannot distinguish between different icons. Nonetheless, the base LVLMs show reasonable performance, but since the number of objects of interest is large in this dataset, all three base LVLMs are outperformed by LVLM-Count. The results for the clustering approach show that LVLM-Count is quite effective even when no prior knowledge is used to set the start and end points of the division paths.
Table 3: Evaluation of state-of-the-art models on the Emoji-Count benchmark. Columns marked with $\Delta$ show the improvement LVLM-Count brings over the base LVLM it uses. Please refer to Table 12 for more accuracy metrics.
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC (Shi et al., 2024) 64.64 - 87.45 - $\text{DAVE}_{\text{prm}}$ (Pelhan et al., 2024) 198.99 - 208.08 - CountGD (Amini-Naieni et al., 2024) 137.93 - 156.80 - GroundingREC (Dai et al., 2024) 36.16 - 51.88 - Number of the target segmentation masks 82.47 - 107.98 - Base GPT-4o counting the target segmentation masks 107.72 - 162.12 - GPT-4o 22.51 - 35.94 - LVLM-Count (GPT-4o as LVLM, Cluster-based on both axes) 11.10 $\downarrow 11.41$ 24.23 $\downarrow 11.71$ LVLM-Count (GPT-4o as LVLM) 9.82 $\downarrow 12.69$ 19.46 $\downarrow 16.48$ Gemini 1.5 Pro 18.17 - 27.83 - LVLM-Count (Gemini 1.5 Pro as LVLM) 14.35 $\downarrow 3.82$ 23.42 $\downarrow 4.41$ Qwen2 VL 72B AWQ 82.41 - 186.32 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 20.67 $\downarrow 61.73$ 34.48 $\downarrow 151.84$
5 Limitations and Future Work
A limitation of LVLM-Count is the area detection stage. If the cropped areas do not provide enough context, performance may suffer. This is especially true for complex questions that require understanding the relationships between all the objects and the background in an image. This opens up opportunities for future work on improved area detection methods or the use of a context provider to complement the area detection stage. Another limitation arises with images containing thousands of objects. After one iteration of division, a significant number of objects may still remain in the sub-images. A potential solution is to apply additional iterations of division on the sub-images; however, the low resolution of these sub-images may make this infeasible. Developing a solution to maintain resolution is another direction for future work. Finally, in some cases, sub-images do not contain any objects of interest. The LVLM occasionally predicts a non-zero value in such instances. This is a weakness of LVLMs that requires special consideration to improve their performance.
Appendix A Ablation Study
We examine the effect of each of the following stages in our method: i) area detection and ii) object-aware division (note that the object-aware division necessitates the inclusion of target segmentation stage). The experiments are designed to investigate the effect of each stage individually, as well as when the stages are combined in our pipeline. Additionally, we run an experiment for a case where both stages of area-detection and target segmentation are excluded. In this case, the object-aware division can not be performed. Thus, images are divided by equidistant straight lines into subimages. We give the name of naive division to such an approach. Moreover, we run another experiment where area detection is in place but the target segmentation is excluded and naive division is applied on the detected areas We run the ablation scenarios for two LVLMs: GPT-4o, and Gemini 1.5 Pro. Furthermore, the effect of super resolution at the final stage of LVLM-Count is also investigated for the case that GPT-4o is the LVLM. For the experiments, we randomly sample $4$ images from each category in the FSC-147 test set ($29$ categories) and report the performance on the resulting subset containing $116$ samples. In Table 4 we show the results of the ablation experiments.
Table 4: Ablation study for LVLM-Count on 116 samples from the FSC-147 test dataset. Columns marked with $\Delta$ show the performance difference between an entry and the base LVLM it uses. Green indicates improvement, while red represents degradation.
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT-4o 18.58 - 56.84 - GPT-4o + Naive division 33.95 $\uparrow 15.37$ 88.75 $\uparrow 31.91$ GPT-4o + Area Detection + Naive division 34.82 $\uparrow 16.24$ 73.19 $\uparrow 16.35$ GPT-4o + Area detection 19.27 $\uparrow 0.69$ 70.64 $\uparrow 13.8$ GPT-4o + Object-aware division 14.92 $\downarrow 3.66$ 31.56 $\downarrow 25.28$ GPT-4o + Area detection + Object-aware divsion (equiv. to LVLM-Count) 11.03 $\downarrow 7.55$ 21.02 $\downarrow 35.82$ GPT-4o + Area detection + Object-aware divsion + Super resolution 12.30 $\downarrow 6.28$ 28.98 $\downarrow 27.86$ Gemini 1.5 Pro 22.07 - 64.34 - Gemini 1.5 Pro + Naive division 41.31 $\uparrow 19.24$ 76.56 $\uparrow 12.22$ Gemini 1.5 Pro + Area Detection + Naive division 43.71 $\uparrow 21.64$ 75.47 $\uparrow 11.13$ Gemini 1.5 Pro + Area detection 23.04 $\uparrow 0.97$ 61.94 $\downarrow 2.4$ Gemini 1.5 Pro + Object-aware division 15.39 $\downarrow 6.68$ 38.15 $\downarrow 26.19$ Gemini 1.5 Pro + area detection + Object-aware division (equiv. to LVLM-Count) 13.64 $\downarrow 8.43$ 36.56 $\downarrow 27.78$
Appendix B Bias to Small Numbers in Datasets and Performance of LVLMs
In our experiments, LVLMs (GPT-4o) are able to make correct predictions when the number of items to be counted is small, but errors increase as the ground truth number grows. Although, we cannot be certain, one likely reason for this behavior in LVLMs is that, during training, the counting questions these models encounter are heavily biased toward small numbers. As an example, in Figure 8 we show the distribution of ’How many’ questions in some well-known VQA datasets.
(a) TallyQA Dataset [Acharya et al., 2019]
(b) VQAv2 Dataset [Goyal et al., 2017]
Figure 8: Distribution of answers in two VQA datasets.
Appendix C Definition and Example of Simple and Complex Counting Tasks
Acharya et al. [2019] were among the first to formally categorize counting questions into simple and complex types. They applied a linguistic rule: first, they removed any substrings such as “…in the photo?” or “…in the image?”. Then, they used SpaCy to perform part-of-speech tagging on the remaining substring. They classified a question as simple if it contained only one noun, no adverbs, and no adjectives; otherwise, they deemed it complex. This rule classifies questions such as “How many dogs?” as simple and “How many brown dogs?” as complex. Following this rule, they built two splits for the TallyQA dataset: a simple split and a complex split. Since we sample our open-ended counting benchmark from the TallyQA simple and complex splits, we adopt the same classification criteria for this benchmark.
In Figure 9, we provide examples of a simple and a complex VQA question from TallyQA. The VQA question in Figure 9(a) is selected from the simple split of TallyQA. It is a straightforward counting question, merely asking for all instances of the animal. Figure 9(b), on the other hand, is selected from the complex split of TallyQA. In addition to counting, it requires detecting context and distinguishing between consonants and vowels.
(a) How many cows are visible?
(b) How many consonants are there on the green sign?
Figure 9: A simple (left image) and a complex (right image) question from TallyQA.
Appendix D Incorrect Examples in the FSC-147 Dataset
We observed some incorrect instances in the FSC-147 dataset. For example, see Figure 10. In these cases, the category names which are provided are incorrect. For these instances, the extensive knowledge embedded within the LVLMs employed in our approach proves to be a disadvantage. These models detect inconsistencies and provide a count of zero as the output, whereas other methods are misled by superficial similarities and mistakenly count the objects. A more thorough study is required to detect all the incorrect examples in FSC-147.
(a) Category name in FSC-147: apples, Answer in FSC-147: $182$, Ground truth: 00.
(b) Category name in FSC-147: sunglasses, Answer in FSC-147: $81$, Ground truth: 00.
Figure 10: Two erroneous examples from FSC-147. Figure 10(a) shows a number of crabapples while the category name in the FSC-147 is apples. Figure 10(b) features a number of glasses while the category name in FSC-147 is sunglasses.
Appendix E Alleviating the Problem of Inaccurate Detections at the Area Detection and Target Segmentation Stages of LVLM-Count
One of the limitations of LVLM-Count is that if the area detection fails to detect a relevant area, the objects of interest in that area will not be counted. Another limitation arises when the target segmentation fails to segment all the instances of the target category of objects. In such a scenario, non-segmented objects of interest might be cut through by division lines, causing the predicted total number to increase due to being counted multiple times. One simple, yet quite effective, approach that we use to overcome this limitation to a significant degree is setting the detection threshold of GroundingDINO, which is used in both of the mentioned stages, to an extremely low value. This significantly reduces the chance of missing an area or object of interest, although it might lead to some false positives. However, note that false positives do not harm the performance of LVLM-Count, as their only effect is that the object-aware division lines avoid cutting through them.
To demonstrate the effectiveness of this approach, we evaluate it on the Penguin dataset [Penguin Research, 2016]. The goal in this dataset is to count penguins in images. The challenging Penguin dataset consistently exhibits heavy occlusion and complex background patterns that can easily be mistaken for penguins Arteta et al. [2016]. This dataset consists of two splits: the mixed-site split, in which images from the same camera can appear in both the training and testing sets, and the separate-site split, in which images in each set strictly belong to different cameras. Images in this dataset are annotated by multiple annotators, where each annotator might identify a different number of penguins due to the challenges in locating them within the images. Since annotators usually undercount the penguins, similar to Arteta et al. [2016], we take the maximum number of penguins among the annotations as the ground truth and calculate MAE and RMSE with respect to this value. For more details about the dataset and the calculation of metrics, we refer the reader to Arteta et al. [2016].
Since both splits are very large, we randomly select 100 samples from each split. To preserve balance, the probability of selecting an image with a certain ground-truth annotation number is inversely proportional to the frequency of that annotation number in the entire split, excluding images with zero penguins. We set the area detection and target segmentation thresholds to 0.5, and then we run them with both thresholds set to 0.1. The results are presented in Table 5 and Table 6. We observe that for both splits, when the detection threshold is high, the MAE generally increases to a value higher than the MAE of the base LVLMs. However, setting the thresholds to a low value improves the MAE for all three LVLMs.
Table 5: Penguin dataset - Mixed sites. Columns marked with $\Delta$ show the performance difference between LVLM-Count and the base LVLM it uses. Green indicates improvement, while red represents degradation.
Method MAE (Max)$\downarrow$ $\Delta$ RMSE (Max)$\downarrow$ $\Delta$ GPT-4o 26.45 - 37.53 - LVLM-Count(GPT-4o - Thredshold=0.5) 36.06 $\uparrow$ 9.61 49.23 $\uparrow$ 11.7 LVLM-Count(GPT-4o - Thredshold=0.1) 23.54 $\downarrow$ 2.91 42.24 $\uparrow$ 4.71 Gemini 1.5 Pro 37.23 - 49.15 - LVLM-Count(Gemini 1.5 Pro - Threshold=0.5) 42.42 $\uparrow$ 5.19 50.39 $\uparrow$ 1.24 LVLM-Count(Gemini 1.5 Pro - Threshold=0.1) 20.09 $\downarrow$ 17.14 29.95 $\downarrow$ 19.2 Qwen2 VL 72B AWQ 34.25 - 55.70 - LVLM-Count(Qwen2 VL 72B AWQ - Threshold=0.5) 37.13 $\uparrow$ 2.87 53.71 $\downarrow$ 1.99 LVLM-Count(Qwen2 VL 72B AWQ - Threshold=0.1) 18.22 $\downarrow$ 16.03 28.76 $\downarrow$ 26.94
Table 6: Penguin dataset - Separate sites. Columns marked with $\Delta$ show the performance difference between LVLM-Count and the base LVLM it uses. Green indicates improvement, while red represents degradation.
Method MAE (Max)$\downarrow$ $\Delta$ RMSE (Max)$\downarrow$ $\Delta$ GPT4o 36.01 - 46.64 - LVLM-Count(GPT-4o - Thredshold=0.5) 43.25 $\uparrow$ 7.24 52.17 $\uparrow$ 5.53 LVLM-Count(GPT-4o - Thredshold=0.1) 27.11 $\downarrow$ 8.9 36.72 $\downarrow$ 9.92 Gemini 1.5 Pro 46.67 - 57.26 - LVLM-Count(Gemini 1.5 Pro - Threshold=0.5) 33.59 $\uparrow$ 13.08 46.23 $\uparrow$ 11.03 LVLM-Count(Gemini 1.5 Pro - Threshold=0.1) 28.50 $\downarrow$ 18.17 41.39 $\downarrow$ 15.87 Qwen2 VL 72B AWQ 44.96 - 66.78 - LVLM-Count(Qwen2 VL 72B AWQ - Threshold=0.5) 51.17 $\uparrow$ 6.21 75.49 $\uparrow$ 8.71 LVLM-Count(Qwen2 VL 72B AWQ - Threshold=0.1) 24.83 $\downarrow$ 20.13 38.92 $\downarrow$ 27.86
In order to also provide a visual insight into this limitation, we chose an image from the dataset where distinguishing penguins from the background features is challenging for GroundingDINO. Figure 11 shows the performance of our method on this image, with the area detection and target segmentation thresholds both set to 0.5. We can observe that the failure of the target segmentation stage in segmenting all instances of penguins in Figure 11(a) has resulted in division paths cutting through several penguins inside the areas surrounded by yellow shapes in Figure 11(b). Consequently, these penguins are counted twice, causing a larger error.
On the other hand, Figure 12 shows the performance of LVLM-Count when the area detection and target segmentation thresholds are set to 0.1. Figure 12(a) illustrates that, while a low threshold has helped to segment all instances of the penguins, it has also caused some false positive areas to be segmented. Nonetheless, it can be observed in Figure 12(b) that the only effect of false positive segmentation masks is that the division paths avoid cutting through those regions as well. Note that for both images, the cluster-based approach was used to automatically find the start and end points of the division paths based on the arrangement of the segmentation masks in the scene. Comparing Figure 11 and Figure 12, we can clearly see that reducing the detection threshold is a quite successful technique to overcome limitations in the initial stages such as target segmentation.
(a) With a threshold of 0.5, target segmentation stage has failed to segment all penguins
(b) Areas surrounded by yellow shapes contain several penguins bisected by the dividing lines.
Figure 11: In this figure, the detection threshold in the target segmentation stage is set to 0.5. For this example, the cluster-based approach has been used to find the start and end points of the division paths. Ground truth (Max) = 106, LVLM-Count prediction = 131.
(a) With a low threshold such as 0.1, some false positive areas have also been segmented.
(b) The division paths avoid cutting through the false positive areas as well.
Figure 12: In this figure, the detection threshold in the target segmentation stage is set to 0.1. For this example, a cluster-based approach was used to find the start and end points of the division paths. Ground truth (Max) = 106, LVLM-Count prediction = 110
Additionally, Figure 13 shows how using a low detection threshold can help alleviate the limitations in the area detection stage for a challenging image taken in unfavorable weather conditions. Figure 13(a) shows the output of the area detection stage when the threshold is set to 0.5. The output contains two detected areas that, in combination, fail to cover all the penguins, resulting in error. However, Figure 13(b) shows the output of the area detection stage when the threshold is set to 0.1, successfully giving the area containing all the penguins.
(a) The area detection threshold set to 0.5
(b) The area detection threshold set to 0.1
Figure 13: This figure shows that setting the detection threshold in the area detection stage to a very low value (Figure 13(b)) significantly alleviates the area detection limitations at the area detection stage.
Appendix F LVLM-Count’s Power in Handling Multiple Object Categories in the Same Image
LVLM-Count is a highly effective method for handling counting tasks that involve multiple objects in the same image. Its strength in such scenarios stems from the capabilities of LVLMs to answer numerous visual questions about an image and its objects. Depending on the given text prompt, it can count instances of a single object category among others or instances of multiple object categories simultaneously. In this section, we demonstrate how LVLM-Count performs in counting different objects of interest, determined simply by a prompt, using an image with multiple object categories.
The image in Figure 14 contains three object categories: person, cow, and horse. In the top row, the object of interest is “cow.” We prompt LVLM-Count to count the cows. First, the masks are produced through the initial stages of our pipeline, and then the cluster-based approach is used to automatically determine the start and end points of the division paths. It can be observed that horses have also been masked as cows. Nonetheless, this does not negatively impact the final answer; it merely causes the division lines to avoid cutting through the horses as well. The counting in LVLM-Count is performed by an LVLM (GPT-4o in this figure) and does not rely on the masks. We observe that GPT-4o successfully counts the number of cows in the resulting subimages, leading to the correct final answer.
In the middle row of Figure 14, the object of interest is “person.” LVLM-Count again successfully counts the number of people accurately. A more interesting case is the bottom row of Figure 14, where both cows and persons are objects of interest. We prompt LVLM-Count to count the number of “cows and persons.” Similar to the first row of the figure, there are false positive masks here as well. However, LVLM-Count successfully counts the number of instances from both categories combined since the counting is ultimately performed by the LVLM. Note that the number of objects in this image is limited, and GPT-4o might answer these questions correctly without the need for the LVLM-Count pipeline. This image has been chosen to illustrate LVLM-Count’s power in handling multiple objects in a counting task rather than for comparison with the baseline LVLM.
Figure 14: Illustration of the ability of LVLM-Count in counting an object of interest determined my a prompt when multiple object categories exist in a single image. Top row: Object of interest is “person”. Middle row: Object of interest is “cow”. Bottom row: Object of interest is “person and cow”
Appendix G Real-world Application of LVLM-Count
As stated in Section 1, counting has numerous real-world applications, including but not limited to biology, health, industry, warehousing, and environmental monitoring. Below, we demonstrate the performance of LVLM-Count on examples from the following areas: i) biology/health, ii) industry/warehousing, and iii) environmental monitoring. We also compare its results with those of the base LVLM (GPT-4o for the figures in this section). Note that in all examples, the cluster-based approach automatically determines the start and end points of the division paths.
In Figure 15, images of two laboratory samples are analyzed using LVLM-Count. The first row shows an image from a dataset introduced by Lempitsky and Zisserman [2010], which contains simulated bacterial cells from fluorescence-light microscopy, created by Lehmussola et al. [2007]. The second row shows an image from the BM dataset introduced by Kainz et al. [2015], which contains bone marrow samples from eight healthy individuals. The standard staining procedure highlights the nuclei of various cell types in blue, while other cellular components appear in shades of pink and red [Paul Cohen et al., 2017]. As observed, LVLM-Count achieves much higher accuracy in counting bacterial cells and bone marrow nuclei in the top and bottom rows of Figure 15, respectively, compared to the base LVLM, particularly for the bone marrow nuclei.
In Figure 16, two images from industrial scenes are analyzed using LVLM-Count. The top row shows a sectional image of a stockpile of tree logs, and the bottom row shows an image from an industrial area containing barrels of various colors. For the top image, the objects of interest are the tree logs, while for the bottom image, LVLM-Count is tasked with counting the blue barrels. In both cases, LVLM-Count’s predictions are significantly closer to the ground truth values than those of the base LVLM, particularly for the tree logs, where the ground truth number is too large for the base LVLM to estimate accurately.
Figure 17 shows an image sourced from a dataset [Penguin Research, 2016] created as part of an ongoing initiative to monitor the penguin population in Antarctica. This dataset comprises images captured hourly by a network of fixed cameras installed at over 40 locations. Over several years, this effort has accumulated over 500,000 images. Zoologists use these images to identify trends in penguin population sizes at each site, facilitating studies on potential correlations with factors such as climate change. Thus, determining the number of penguins in each image is crucial. Given the challenges of engaging human annotators to process such a vast dataset, automating the counting task is highly desirable [Arteta et al., 2016]. LVLM-Count is prompted to count the number of penguins in the image, and as observed, its predictions are significantly closer to the ground truth than those of the base LVLM.
Figure 15: Performance of LVLM-Count on real-world applications in biology/health. The top row shows an image of simulated bacterial cells from fluorescence-light microscopy [Lempitsky and Zisserman, 2010], with the objects of interest being “bacterial cells.” The bottom row shows an image of bone marrow, with the nuclei of various cell types highlighted in blue [Kainz et al., 2015], and the objects of interest being “bone marrow nuclei.”
Figure 16: Performance of LVLM-Count on real-world applications in industry/warehousing. The top row shows an image of a stockpile of tree logs, with the objects of interest being “tree logs.” The bottom row shows an aerial image of an industrial area containing barrels of various colors, with the objects of interest being “blue barrels.”
Figure 17: Performance of LVLM-Count on real-world applications in environmental monitoring. The image is sourced from [Penguin Research, 2016], an initiative to monitor the penguin population in Antarctica, with the objects of interest being “penguins.”
Appendix H Performance Analysis of LVLM-Count for Different Ground Truth Ranges on FSC-147 Dataset
To further investigate the performance of our pipeline, we divide the ground truth values in the FSC-147 test set into intervals and plot the MAE for the base GPT-4o and Gemini 1.5 Pro models, alongside the results from LVLM-Count using each model, as shown in Figure 18. The first interval contains relatively small ground truth values, a range where LVLMs already perform well. As the ground truth values increase, the base models exhibit increasingly larger errors compared to LVLM-Count, with the margin growing rapidly. This behavior is consistent with our observations of counting errors on the blue circles in Figure 2.
(a) LVLM-Count using GPT-4o
(b) LVLM-Count using Gemini 1.5 Pro
Figure 18: Performance analysis of our method, LVLM-Count, on the FSC-147 test set using GPT-4o (Figure 18(a)) and Gemini 1.5 Pro (Figure 18(b)). In the first interval, both base LVLMs exhibit a lower MAE. However, in intervals with higher ground truth values, LVLM-Count achieves a lower MAE compared to the base LVLMs, and this difference increases rapidly.
Appendix I Illustration of the Workflow for the Zebra Image in Figure 1
In this paper, we choose the most appropriate images for simple and understandable illustrations of the execution of each stage in our pipeline. Inevitably, this led to the selection of different images for each stage. However, in this section, for the sake of consistency, we demonstrate the same concepts illustrated in Figures 3, 4, 5, and 6 for the zebra image used in Figure 1.
The zebra image is passed to the pipeline along with the question $Q=$“how many zebras are in the image?”. First, $E=$“zebra” is extracted using the LLM. Then the zebra image is passed to the area detection stage, where the prompt given to GroundingDINO is “zebras”. The output bounding boxes are merged, and the resulting area is cropped, as illustrated in Figure 19. The cropped area is then passed to the target segmentation stage. At this stage, GroundingDINO detects the objects of interest defined by $E$ as the input prompt. SAM then uses the output bounding boxes to produce segmentation masks for the zebras, as shown in Figure 20.
Figure 19: Illustration of the area detection step of LVLM-Count for the zebra image. For this image, $Q$ is set to “How many zebras are in the image?”. The LLM used in this step returns an $E$, which is “zebra”. The plural form of $E$, “zebras”, and the original image are given as input to GroundingDINO, which returns some bounding boxes (left and upper right images) that are merged to form the final detected area.
(a) GroundingDINO output
(b) SAM output
Figure 20: Illustration of the target segmentation step of LVLM-Count for the zebra image. The goal is to produce all the instance masks for $E$ set to “zebra”. The cropped image from Figure 19, together with $E$, is given as input to GroundingDINO, which produces the output shown in Figure 20(a). Figure 20(a) is then given as input to SAM, which produces the output shown in Figure 20(b).
After the target segmentation stage, the masks are passed to the object-aware division stage. First, the masks are used in the cluster-based approach to find the location of the start and end points of the division paths, i.e., ($P^{1}_{s},P^{1}_{e}$) and ($P^{2}_{s},P^{2}_{e}$). Then these masks are turned into black-and-white images, which, in turn, are mapped to a graph. The division paths are then found by connecting each start point to its corresponding end point by running the $A^{*}$ search algorithm on the graph. The found paths are mapped back into the image domain and drawn in red, as depicted in Figure 21. The image contours are determined based on the drawn red paths, and each contour’s interior is masked out independently to obtain the subimages. Finally, the subimages are given to the LVLM to count the number of zebras in each.
Figure 21: Left: Illustration of the unsupervised and non-parametric method to obtain the division points ($P^{1}_{s},P^{1}_{e}$) and ($P^{2}_{s},P^{2}_{e}$). First, a few pixels are sampled (shown as points inside the segmented objects) from the pixels composing each mask. The samples are projected onto the $x$-axis. The projected points are clustered using mean-shift clustering. The point in the middle of two consecutive clusters is considered a vertical division point. The straight vertical lines are drawn just for better visualization of the division points. Right: Illustration of object-aware division. The masks from Figure 20(b) are turned into a black-and-white image. A dividing path is found by connecting $P_{s}$ to $P_{e}$ using the $A^{*}$ search algorithm in a graph that corresponds to the binary image, where the only nodes in the graph are white pixels, which are fully connected. The path is mapped back to the pixel domain.
Appendix J False Positive Masks at the Target Segmentation Stage
One of the reasons we task an LVLM to count the objects in the subimages instead of using the number of generated masks at the target segmentation stage as the final count of the objects of interest is the existence of false positive masks. The GroundingDINO model is responsible for detecting the objects of interest, determined by expression $E$, and passing the output bounding boxes to SAM for producing segmentation masks. Nonetheless, GroundingDINO is not as strong as an LVLM in understanding expressions extracted from complex questions. Thus, it often returns bounding boxes for all instances of the object category mentioned in the expression, even if those instances do not satisfy other conditions in the expression.
For example, in the top row of Figure 22, $E=$brown egg”. However, all the eggs have been segmented regardless of their color. Thus, counting the masks results in a significant error. Interestingly, as we can see, the false positive masks do not negatively affect LVLM-Count’s final answer, as the counting is done by an LVLM at the final stage, which is much stronger than GroundingDINO at understanding referring expressions. The only effect is that the white eggs have not been cut through by the division lines either. In the bottom row, we have chosen an image from the challenging Emoji-Count benchmark. The image contains icons, all of which have an arrow but point in different directions. However, the objects of interest are only “right arrows curving left.” Similar to the eggs example, taking the masks used for object-aware division results in a significant error.
Figure 22: Top row: The object of interest is “brown egg.” However, all the eggs have been segmented because of the limitation of the GroundingDINO model in understanding complex referring expressions. Regardless, LVLM-Count provides a significantly more accurate number compared to the number of masks. Bottom row: The object of interest is “right arrows curving left.” Similar to the image of the eggs, counting the number of masks results in a very large error, while LVLM-Count provides a much more accurate number.
Appendix K LVLM-Count’s Performance on a Benchmark Sampled from the PASCAL VOC Dataset
The PASCAL VOC dataset [Everingham et al., 2015] is a well-known dataset in the field of computer vision, depicting everyday objects in everyday scenes. It is primarily used for tasks like object detection, classification, and segmentation. To provide more experimental results for the performance of LVLM-Count, we adapt a small subset of this dataset into the form of a counting benchmark. Similar to Chattopadhyay et al. [2017], we choose PASCAL VOC 2007 among other variants. This variant contains a training set of 2501 images, a validation set of 2510 images, and a test set of 4952 images, with 20 object categories that remain consistent across the splits.
Each image includes annotations for instances of the 20 object categories in the dataset. The most frequent count per object category (as one would expect in everyday scenes) is 0. Moreover, there is a clear bias toward lower count values. To address this bias and obtain a balanced benchmark, we first create 20 simple counting questions asking for the number of objects from each of the 20 categories for every image in the test set. Then, we randomly sample five questions for each ground truth count. Note that the ground truth counts for this sampling are based on the original annotations of the dataset. The bias toward lower numbers is so significant that for ground truth counts larger than 16, there are fewer than five questions for the entire test set. In such cases, we simply use the available samples. This process resulted in 102 questions in total. Finally, we manually checked the ground truth counts and corrected them if required.
We call the resulting counting questions the PASCAL VOC counting benchmark. We evaluate the performance of LVLM-Count on this benchmark using three different LVLMs: GPT-4o, Gemini 1.5 Pro, and Qwen2 VL 72B AWQ. Note that we use a clustering approach to automatically find the start and end points of the division paths along the $x$-axis. Table 7 shows the performance of LVLM-Count in comparison with state-of-the-art counting models and the base LVLMs. We observe that LVLM-Count improves upon the base LVLMs it uses and outperforms all prior counting methods, except CountGD. Upon further inspection, we noticed that the 20 object categories in PASCAL VOC have a high overlap with the object categories of the dataset used to train CountGD.
Table 7: Evaluation of state-of-the-art models on the PASCAL VOC counting benchmark. The results for base LVLMs and LVLM-Count are reported over three trials. Columns marked with $\Delta$ show the improvement brought by our method over the base LVLM that it uses.
Method EA (%$\uparrow$ $\Delta$ MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TrainingFree [Shi et al., 2024] 2.94 - 12.03 - 18.18 - GroundingRec [Dai et al., 2024] 19.61 - 5.05 - 8.44 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 5.88 - 12.39 - 22.81 - CountGD [Amini-Naieni et al., 2024] 26.47 - 2.81 - 7.01 - GPT4o 30.39 - 4.46 - 8.35 - LVLM-Count (GPT4o as LVLM) 31.37 $\uparrow$ 0.98 3.55 $\downarrow$ 0.91 7.18 $\downarrow$ 1.17 Gemini 1.5 Pro 33.01 - 3.24 - 6.62 - LVLM-Count (Gemini 1.5 Pro as LVLM) 39.22 $\uparrow$ 6.21 3.00 $\downarrow$ 0.24 6.30 $\downarrow$ 0.32 Qwen2 VL 72B AWQ 24.18 - 4.83 - 8.84 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 29.08 $\uparrow$ 4.90 4.12 $\downarrow$ 0.71 7.59 $\downarrow$ 1.25
Appendix L Report of Various Accuracy Metrics for the Performance of LVLM-Count on the FSC-147 dataset, TallyQA Benchmark, and Emoji-Count Benchmark
This section presents various accuracy measures for the experiments reported in Tables Table 1, 2, and 3. The accuracy metrics are defined in Table 8. The observations for each table can be summarized as follows:
i) FSC-147: For this dataset, similar to the results in Table 1, LVLM-Count achieves higher accuracy metrics compared to the prior training-free method and each of the base LVLMs it uses. However, models specifically trained on this dataset generally achieve higher accuracy.
ii) TallyQA Simple Benchmark: For this benchmark, LVLM-Count achieves higher accuracy compared to all prior counting models. However, it does not surpass the accuracy of the base LVLMs used. This is because the questions are straightforward, and the ground truth values are limited to numbers between 0 and 15—a range where the base LVLMs excel. This observation aligns with those in Figure 2 and Figure 18.
iii) TallyQA Complex Benchmark: For this benchmark, LVLM-Count demonstrates significantly higher accuracies compared to prior counting models and, more importantly, consistent accuracy improvements over the base LVLMs used.
iv) Emoji-Count: This is a challenging benchmark due to high object counts. LVLM-Count achieves substantially higher accuracies than both the base LVLMs and prior counting models, particularly for metrics like Acc±$\pm k$ where $k\geq 1$.
Table 8: Definitions of Various Accuracy Metrics. GT denotes the ground truth number.
Metric Definition
Acc Percentage of answers such that $answer=GT$
Acc±$\pm 1$ Percentage of answers such that $|answer-GT|\leq 1$
Acc±$\pm 3$ Percentage of answers such that $|answer-GT|\leq 3$
Acc±$\pm 5$ Percentage of answers such that $|answer-GT|\leq 5$
Acc±$\pm 10$ Percentage of answers such that $|answer-GT|\leq 10$
Table 9: FSC-147 Dataset. A ($\uparrow$) next to the measured accuracies for LVLM-Count indicates improvement over the base LVLM it uses, while a ($\downarrow$) indicates degradation compared to the corresponding base LVLM.
Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) TFOC [Shi et al., 2024] 9.33 20.17 34.20 44.37 61.60 GroundingRec [Dai et al., 2024] 34.03 51.68 67.65 75.04 85.13 CountGD [Amini-Naieni et al., 2024] 31.85 47.90 63.78 73.03 82.61 GPT4o 12.24 26.22 42.10 52.41 66.58 LVLM-Count (GPT4o as LVLM) 14.26 ($\uparrow$) 28.10 ($\uparrow$) 47.42 ($\uparrow$) 58.01 ($\uparrow$) 72.91 ($\uparrow$) Gemini 1.5 Pro 12.97 26.58 41.57 51.71 63.78 LVLM-Count (Gemini 1.5 Pro as LVLM) 13.92 ($\uparrow$) 27.96 ($\uparrow$) 47.65 ($\uparrow$) 58.85 ($\uparrow$) 75.07 ($\uparrow$) Qwen2 VL 72B AWQ 9.19 20.81 36.83 46.95 62.07 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 9.80 ($\uparrow$) 23.14 ($\uparrow$) 40.95 ($\uparrow$) 51.04 ($\uparrow$) 68.54 ($\uparrow$)
Table 10: TallyQA Simple Benchmark. A ($\uparrow$) next to the measured accuracies for LVLM-Count indicates improvement over the base LVLM it uses, while a ($\downarrow$) indicates degradation compared to the corresponding base LVLM.
Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) TFOC [Shi et al., 2024] 6.45 17.42 34.84 59.35 79.35 GroundingRec [Dai et al., 2024] 23.87 49.03 73.55 86.45 93.55 CountGD [Amini-Naieni et al., 2024] 41.94 63.23 78.06 85.16 94.84 GPT4o 44.30 69.89 87.96 94.41 100.00 LVLM-Count (GPT4o as LVLM) 44.73 ($\uparrow$) 73.33 ($\uparrow$) 91.83 ($\uparrow$) 96.99 ($\uparrow$) 99.57 ($\downarrow$) Gemini 1.5 Pro 47.10 75.27 92.90 97.63 100.00 LVLM-Count (Gemini 1.5 Pro as LVLM) 45.16 ($\downarrow$) 76.13 ($\uparrow$) 91.61 ($\downarrow$) 95.91 ($\downarrow$) 99.35 ($\downarrow$) Qwen2 VL 72B AWQ 49.03 70.75 87.74 93.33 98.92 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 41.72 ($\downarrow$) 67.10 ($\downarrow$) 85.81 ($\downarrow$) 93.55 ($\uparrow$) 98.49 ($\downarrow$)
Table 11: TallyQA Complex Benchmark. A ($\uparrow$) next to the measured accuracies for LVLM-Count indicates improvement over the base LVLM it uses, while a ($\downarrow$) indicates degradation compared to the corresponding base LVLM.
Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) TFOC [Shi et al., 2024] 1.34 15.44 32.89 46.31 66.44 GroundingRec [Dai et al., 2024] 16.78 28.19 51.68 64.43 85.23 CountGD [Amini-Naieni et al., 2024] 5.37 12.75 31.54 48.99 73.15 GPT4o 29.08 50.34 74.94 86.13 98.21 LVLM-Count (GPT4o as LVLM) 34.68 ($\uparrow$) 55.03 ($\uparrow$) 78.97 ($\uparrow$) 89.71 ($\uparrow$) 96.42 ($\uparrow$) Gemini 1.5 Pro 25.73 25.73 81.66 91.72 98.88 LVLM-Count (Gemini 1.5 Pro as LVLM) 26.62 ($\uparrow$) 51.68 ($\uparrow$) 76.06 ($\downarrow$) 85.23 ($\downarrow$) 94.41 ($\downarrow$) Qwen2 VL 72B AWQ 24.61 46.31 68.01 78.75 94.41 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 30.65 ($\uparrow$) 55.26 ($\uparrow$) 78.30 ($\uparrow$) 86.80 ($\uparrow$) 96.64 ($\uparrow$)
Table 12: Emoji Benchmark. A ($\uparrow$) next to the measured accuracies for LVLM-Count indicates improvement over the base LVLM it uses, while a ($\downarrow$) indicates degradation compared to the corresponding base LVLM.
Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) TFOC [Shi et al., 2024] 0.24 0.72 2.89 4.58 8.43 GroundingRec [Dai et al., 2024] 4.58 8.43 12.29 17.11 24.10 CountGD [Amini-Naieni et al., 2024] 0.48 0.72 0.96 0.96 1.20 GPT4o 1.85 5.54 13.73 21.12 43.94 LVLM-Count (GPT4o as LVLM) 4.98 ($\uparrow$) 17.03 ($\uparrow$) 37.67 ($\uparrow$) 55.90 ($\uparrow$) 76.71 ($\uparrow$) Gemini 1.5 Pro 2.65 7.23 14.14 23.37 42.97 LVLM-Count (Gemini 1.5 Pro as LVLM) 2.25 ($\downarrow$) 5.46 ($\downarrow$) 13.57 ($\downarrow$) 23.61 ($\uparrow$) 57.19 ($\uparrow$) Qwen2 VL 72B AWQ 0.88 2.89 7.55 11.41 19.76 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 3.37 ($\uparrow$) 8.76 ($\uparrow$) 19.04 ($\uparrow$) 29.00 ($\uparrow$) 48.51 ($\uparrow$)
Appendix M Visual Examples of LVLM-Count’s Performance on the FSC-147 Dataset, TallyQA Benchmark, and Emoji-Count Benchmark
This section presents several visual examples showcasing the performance of LVLM-Count on the FSC-147 dataset, the TallyQA benchmark, and the Emoji-Count benchmark. The LVLM used in the pipeline to generate these visual examples is GPT-4o. Figure 23 illustrates both successful and unsuccessful examples of LVLM-Count’s performance on FSC-147. The top two rows in this figure demonstrate strong performance, where the sub-images preserve sufficient context, enabling the LVLM to predict the correct count. The last row depicts a failure case. Here, the task is to count the number of skis. However, LVLM-Count generates sub-images containing only irrelevant objects. Consequently, GPT-4o is misled by these objects and predicts an incorrect non-zero count.
Figure 23: Illustration of good and bad examples for LVLM-Count. These examples are taken from FSC-147. The first column shows the input image. The second column shows the paths found in the object-aware division stage. The third column illustrates the obtained sub-images and the predicted count by LVLM-Count. The top two rows correspond to good examples where the base GPT-4o has a large error, but LVLM-Count provides an answer very close to the ground truth. The bottom row shows a bad example where the base GPT-4o performs better, and LVLM-Count has a large error.
Additionally, Figure 24 displays an example from the TallyQA Simple benchmark, while Figure 25 illustrates an example from the TallyQA Complex benchmark. Additionally, we include five visual examples from the Emoji-Count benchmark, split into two figures for better clarity and presentation. Figure 26 contains the first three examples, and Figure 27 includes the final two examples. In these visual examples, LVLM-Count consistently achieves more accurate results compared to the base GPT-4o.
Figure 24: A visual example of the performance of LVLM-Count on the TallyQA Simple benchmark. The input question is: “How many people are in the image?”.
Figure 25: A visual example of the performance of LVLM-Count on the TallyQA Complex benchmark. The input question is: “How many pieces of the plane are yellow?”.
Figure 26: Three examples of the performance of LVLM-Count on the Emoji-Count benchmark. Top row: The object of interest is “twelve o’clock”. Middle row: The object of interest is “kissing cat”. Bottom row: The object of interest is “woman with red hair”.
Figure 27: Two additional examples of the performance of LVLM-Count on the Emoji-Count benchmark. Top row: The object of interest is “derelict house”. Bottom row: The object of interest is “right arrow curving left”.

### v3

@@
 Abstract
-Counting is a fundamental skill for various visual tasks in real-life applications, requiring both object recognition and robust counting capabilities.
-Despite their advanced visual perception, large vision-language models (LVLMs) struggle with counting tasks, especially when the number of objects exceeds those commonly encountered during training.
-We enhance LVLMs’ counting abilities using a divide-and-conquer approach, breaking counting problems into sub-counting tasks.
-Unlike prior methods, which do not generalize well to counting datasets on which they have not been trained, our method performs well on new datasets without any additional training or fine-tuning.
-We demonstrate that our approach enhances counting capabilities across various datasets and benchmarks.
+Counting is a fundamental operation for various real-world visual tasks, requiring both object recognition and robust counting capabilities.
+Despite their advanced visual perception, large vision-language models (LVLMs) are known to struggle with counting tasks.
+In this work, we evaluate the performance of several recent LVLMs on visual counting tasks across multiple counting and vision datasets.
+We observe that while their performance may be less prone to error for small numbers of objects, they exhibit significant weaknesses as the number of objects increases.
+To alleviate this issue, we propose a simple yet effective baseline method that enhances LVLMs’ counting ability for large numbers of objects using a divide-and-conquer approach.
+Our method decomposes counting problems into sub-tasks.
+Moreover, it incorporates a mechanism to prevent objects from being split during division, which could otherwise lead to repetitive counting—a common issue in a naive divide-and-conquer implementation.
+We demonstrate the effectiveness of this approach across various datasets and benchmarks, establishing it as a valuable reference for evaluating future solutions.
 1 Introduction
@@
 It improves manufacturing, inventory, and quality control, ensures safety in medical settings, and helps manage resources in environmental efforts (Wang and Wang, 2011; Zen et al., 2012; Arteta et al., 2016).
-Recent advancements in prompt-based models enable counting of unlimited object varieties without visual exemplars.
-Although current trained, and text-prompt-based, counting models by Dai et al.
-(2024); Amini-Naieni et al.
-(2024) perform well on the datasets they are trained on, they face the following challenges.
-First, they require fine-tuning on new datasets.
-Second, since the concepts in the counting datasets are limited, these models do not generalize well to counting questions that involve complex reasoning.
-There are also training-free models by Shi et al.
-(2024), but overall they have weaker performance compared to trained models.
-On the other hand, large vision-language models (LVLMs), such as GPT-4o (Achiam et al., 2023), which also do not need dataset specific training, show very good performance in counting low numbers of objects, usually less than 20, across most datasets.
-However, their performance deteriorates for larger numbers of objects, regardless of the dataset.
-We enhance the accuracy of LVLMs to count objects in images by leveraging their reasoning power within a divide-and-conquer method.
-The LVLMs allow us to handle diverse objects and complex counting questions, while our divide-and-conquer method alleviates challenges associated with counting large numbers of objects.
+Figure 1: Illustration of our proposed pipeline.
+First, an expression ($E$) describing the area of interest is extracted from the prompted question ($Q$), such as “brown eggs”.
+The expression is extracted using a large language model (LLM) which is the same as LVLM in our work.
+Then, $E$ and the image are provided as input to a grounding model, such as the one by Liu et al.
+(2023) to detect the area of interest.
+Second, any objects corresponding to $E$ are segmented.
+Third, in the object-aware division step, we use the segmentation masks to divide the detected area of interest without cutting through the objects of interest.
+Finally, the number of objects of interest in each sub-image is computed using an LVLM, and the results are aggregated.
+Recent advancements in the field of large vision-language models (LVLMs), such as GPT-4o (Achiam et al., 2023), coupled with the massive scale of web-scraped data used to train them, have resulted in unprecedented zero-shot capabilities across diverse tasks.
+However, evaluations of LVLMs’ capabilities have revealed weaknesses in numerical reasoning Yin et al.
+(2023); Xu et al.
+(2024); Yang et al.
+(2023).
+In this work, we focus specifically on the visual counting task, one of the most fundamental aspects of numerical reasoning.
+We observe that although LVLMs perform well in counting small numbers of objects—typically fewer than 20—their accuracy deteriorates with larger quantities.
 Inspired by prior work on the rapid and accurate estimation of small quantities by Chattopadhyay et al.
-(2017), we divide an image into sub-images, and prompt the LVLM to count the objects of interest in each sub-image.
-The counts from each sub-image are then aggregated to make the final prediction.
-Our proposed workflow is illustrated in Figure 1.
-Figure 1: Illustration of our proposed pipeline.
-First, the area of interest, such as a set of zebras, is detected by extracting the object of interest ($E$) from a prompt question ($Q$) using a large language model (LLM) which is the same as LVLM in our work.
-Then, $E$ and the image are provided as input to an area detection model, such as the one by Liu et al.
-(2023).
-Second, any objects corresponding to $E$ are segmented.
-Third, in the object-aware division step, we use the segmentation masks to divide the areas of objects of interest without cutting through the objects.
-Finally, the number of objects of interest in each sub-image is computed using an LVLM, and the results are aggregated.
-Initially, in our pipeline, the object of interest is extracted from the input question using an LLM.
-The area containing the object of interest is detected in the image by a grounding model, such as Liu et al.
+(2017) we propose a simple yet effective baseline method to alleviate this issue.
+Leveraging a divide-and-conquer approach, we engineer a simple pipeline that divides an image into sub-images, and prompts the LVLM to count the objects of interest in each sub-image.
+The counts from the sub-images are then aggregated to make the final prediction.
+A key feature of our pipeline is a mechanism that prevents objects of interest from being split by the dividing lines, which could otherwise lead to double-counting.
+To have this mechanism two existing pre-trained detection and segmentation models are employed in the pipeline.
+The workflow is illustrated in Figure 1.
+Initially, in our pipeline, the category name of the object of interest is extracted from the input question using an LLM (We use the LVLM as an LLM for this step).
+The area containing the objects of interest is detected in the image by a grounding model, such as Liu et al.
 (2023), and then cropped.
-The cropping step is important since it removes irrelevant context from the image.
+The cropping step removes irrelevant context from the image.
 Secondly, using an object detection model by Liu et al.
@@
 We call this mechanism object-aware division.
-The division positions can be either set manually or determined automatically using unsupervised and non-parametric methods.
-In the latter case, we automatically select the division positions based on object masks, then we treat the object-aware division as a path-finding problem, avoiding objects as obstacles.
+The division positions are determined automatically using an unsupervised and non-parametric method based on object masks.
+Then we treat the object-aware division as a path-finding problem, avoiding the segmented objects as obstacles.
 A black-white image is built by converting all the masks into black and the rest of the image into white pixels.
@@
   1. 1.
-    We propose LVLM-Count that leverages the strengths of large vision-language models (LVLMs) in visual perception for counting.
-Our method is a prompt-based counting approach that, in addition to simple counting problems, can handle complex cases as well.
-Our method does not require any additional training or fine-tuning on any dataset.
-LVLM-Count outperforms existing state-of-the-art models across various datasets and benchmarks.
+    We evaluate the counting performance of several recent Large Vision-Language Models (LVLMs) on multiple counting and vision datasets.
+We propose a simple yet effective baseline method, LVLM-Count, which enhances the counting performance of LVLMs without requiring additional training.
+Similar to standard counting with LVLMs, our method is a prompt-based approach that retains their zero-shot capabilities while addressing their difficulties in handling large numbers.
+Through experiments, we demonstrate that LVLM-Count improves LVLMs’ counting performance across the evaluated datasets.
   2. 2.
@@
 Accurate division is crucial, as parts of cut objects can lead to over-counting (see Figure 2(a)).
-To the best of our knowledge, this is the first method to divide images without cutting through objects of interest.
+The proposed solution divides images without cutting through objects of interest specified by an arbitrary prompt.
 (a) Naive division
@@
 The results are averaged over three trials.
-As a minor contribution, we create a new benchmark to address two drawbacks of existing benchmarks.
-Prior datasets either feature simple counting tasks, e.g., counting “strawberries”, or include complex questions with small numbers of objects.
-To address both issues, we develop a challenging benchmark for counting emoji icons.
+As a minor contribution, we create a new benchmark to address some of the drawbacks of existing datasets.
+Prior datasets feature simple counting tasks, e.g., counting “strawberries”, and lack intra-class complexity.
+To address these issues, we develop a challenging benchmark for counting emoji icons.
 The subtle variations within emoji classes make this benchmark uniquely difficult.
-Since none of the models have been exposed to it, this benchmark serves as a fair test of the performance of counting methods for various attributes and complex concepts.
 2 Related Work
@@
 However, most existing class-agnostic, or open-world, models require visual exemplars of the target objects (Đukić et al., 2023; Gong et al., 2022; Lin et al., 2022; Liu et al., 2022; Lu et al., 2019; Nguyen et al., 2022; Ranjan et al., 2021; Shi et al., 2022; Yang et al., 2021; You et al., 2023).
-The concept of divide and conquer has also been used in early work(Xiong et al., 2019; Chattopadhyay et al., 2017).
-However, this early work requires training dedicated models to utilize this concept.
-Text-based counting-specific trained models.
-With the advent of vision-language foundation models such as CLIP and GroundingDINO, text-based open-world methods have been proposed.
-Leveraging the rich textual and visual feature extraction capabilities of foundation models, obtained through web-scale training, the text-based counting methods by Amini-Naieni et al.
+Text-based counting models.
+With the advent of vision-language foundation models such as CLIP and GroundingDINO, text-based open-world models have been proposed that are trained specifically for counting.
+Leveraging the rich textual and visual feature extraction capabilities of foundation models, obtained through web-scale training, the text-based counting models by Amini-Naieni et al.
 (2023); Dai et al.
@@
 (2024) have started to demonstrate comparable or superior accuracy.
-GroundingREC by Dai et al.
-(2024) is an open-world model built on top of GroundingDINO (Liu et al., 2023), and introduces an additional task called referring expression counting.
-Concurrently, Amini-Naieni et al.
-(2024) proposed a method that also builds on GroundingDINO but adds an extra image-text fusion module in the input, enabling the model to accept text and/or visual exemplars to determine the target.
-Models without counting-specific training.
-Shi et al.
+In addition, Shi et al.
 (2024) introduce TFOC, a counting model that does not require any counting-specific training.
-Instead, they cast the counting problem as a prompt-based segmentation task, using SAM (Kirillov et al., 2023) to obtain segmentation maps that determine the output number.
-Another group of models that do not need further training to count are LVLMs.
-State-of-the-art models such as GPT-4o (Achiam et al., 2023), Gemini 1.5 Pro (Reid et al., 2024), and Claude 3.5 Sonnet (Anthropic, 2024) show strong performance in counting small numbers of objects, although their performance degrades with larger numbers of objects.
-In comparison to prior work, our method (LVLM-Count) is an open-world, text-based counting approach that does not require any counting-specific training.
-Although we use pre-trained GroundingDINO and SAM in our method, we differ from prior work by not treating the counting task as a segmentation or detection problem.
-Instead, we use LVLMs as a counting tool.
+Instead, they cast the counting problem as a prompt-based segmentation task, using SAM (Kirillov et al., 2023) to obtain segmentation masks that determine the output number.
+Leveraging the concept of divide and conquer for counting.
+The concept of divide and conquer has been used in early work (Chattopadhyay et al., 2017; Xiong et al., 2019; Stahl et al., 2018).
+Chattopadhyay et al.
+(2017) use an image-level divide and conquer approach and train a convolutional neural network (CNN) that can count objects from a predetermined and limited set of categories in sub-images.
+Xiong et al.
+(2019) propose applying the divide step on the convolutional feature map instead of the input image to avoid repeatedly computing convolutional features for each sub-image, thereby improving efficiency.
+However, the CNN in their work is only capable of counting a single object category.
+Similar to Chattopadhyay et al.
+(2017), Stahl et al.
+(2018) also employ image-level division and train a CNN to count objects from a predetermined set of categories.
+Nonetheless, their method does not require local image annotations for training.
+Assessment of LVLMs’ counting performance.
+Several prior works have explored the visual counting capabilities of LVLMs as part of broader evaluations, underscoring the difficulties these models encounter in counting tasks Yin et al.
+(2023); Xu et al.
+(2024); Yang et al.
+(2023).
+However, these studies have not focused on developing solutions to address these challenges.
+In this work, we conduct a comprehensive quantitative assessment of LVLMs’ counting performance across diverse visual counting benchmarks.
+More importantly, we propose a simple yet effective baseline method to improve their ability to count large numbers.
+Our baseline method is not tailored to any specific model and can be used in a plug-and-play manner with different LVLMs.
+Regarding the categories outlined earlier, our method, LVLM-Count, is an open-world, prompt-based counting approach that requires no additional training.
+To the best of our knowledge, we are the first to propose a divide-and-conquer strategy that avoids splitting and potentially double-counting objects of interest specified by an arbitrary text prompt.
 3 LVLM-Count
@@
 First, in the “Area Detection” stage, we localize areas containing relevant objects.
-Second, in the “Target Segmentation” stage, we identify and segment these local areas.
+Second, in the “Target Segmentation” stage, we identify and segment the objects of interest.
 Third, in the “Object-aware Division” stage, we divide the localized areas into sub-images without cutting through the segmented objects.
 Finally, the LVLM counts the target objects in each sub-image and aggregates the results.
-Figure 1 illustrates the workflow of our method, which we will detail in the following subsections.
-Note that in this section, we illustrate each stage using example images.
-These images are are used solely for illustration purposes.
-LVLM-Count may not outperform other methods on these examples.
+Figure 1 illustrates the workflow of our method, which we detail in the following subsections.
 3.1 Area Detection
@@
 By employing an LLM, the expression $E$ is extracted from $Q$.
-For example, let $Q$ be “How many people are in the boat?”.
-$Q$ is given to an LLM, which is prompted to return the expression $E$ “people in the boat”, referring to the objects we want to count.
-After $E$ is extracted, it is given as input to GroundingDINO along with the image.
-The output of GroundingDINO is a set of bounding boxes that have relevance to $E$ beyond a certain threshold.
+For example, let $Q$ be “How many brown eggs are in the image?”.
+$Q$ is given to an LLM, which is prompted to return the expression $E$, “brown eggs”, referring to the objects we want to count.
+After $E$ is extracted, it is provided as input to GroundingDINO along with the image.
+The output of GroundingDINO may be a single bounding box or a set of bounding boxes that have relevance to $E$ beyond a certain threshold.
 These bounding boxes often overlap and typically contain repeated objects.
@@
 We consider the non-overlapping areas as “detected areas”, which are then cropped to be passed to the next stage.
-Note that the area detection stage is important as it extracts the area with the most relevant context for the counting question.
 This process is illustrated in Figure 3.
-Figure 3: Illustration of the area detection step of LVLM-Count.
-For this image, $Q$ is set to “How many people are in the boat”.
-The LLM that is used in this step returns an $E$ which is “people in the boat”.
-$E$ and the original image are given as input to GroundingDINO, which returns some bounding boxes (left and upper right images) that are merged to form the final detected area.
+Figure 3: Illustration of the area detection stage of LVLM-Count.
+For this image, $Q$ is set to “How many brown eggs are in the image”.
+The LLM that is used in this step returns an $E$ which is “brown eggs”.
+$E$ and the original image are given as input to GroundingDINO, which returns a bounding box.
+If the grounding model returns multiple bounding boxes, they are merged to form the final detected area.
 3.2 Target Segmentation
@@
 We illustrate an example of this process in Figure 4.
-(a) Original image
-(b) GroundingDINO output
-(c) SAM output
+(a) GroundingDINO output
+(b) SAM output
 Figure 4: Illustration of the target segmentation step of LVLM-Count.
 The goal is to produce all the instance masks for $E$ set to “brown egg”.
-Figure 4(a), together with $E$, is given as input to GroundingDINO, which produces the output shown in Figure 4(b).
-Figure 4(b) is then given as input to SAM, which produces the output shown in Figure 4(c).
+The cropped detected area from Figure 3, together with $E$, is given as input to GroundingDINO, which produces the output shown in Figure 4(a).
+Figure 4(a) is then given as input to SAM, which produces the output shown in Figure 4(b).
 How to determine the bounding boxes.
-To determine the bounding boxes, we use GroundingDINO and set the bounding box probability threshold to a low value to avoid missing any objects.
-The bounding boxes alone cannot help with the object-aware division of the cropped images due to their rigid structure, which includes redundant areas in the vicinity of the object and, in the worst case, overlaps with other bounding boxes.
+To determine the bounding boxes, we use GroundingDINO and set the bounding box probability threshold to a low value to avoid missing any object of interest.
+The bounding boxes alone cannot help with the object-aware division of the detected areas due to their rigid structure, which includes redundant areas in the vicinity of the object and, in the worst case, overlaps with other bounding boxes.
 Our goal is to precisely locate the pixels of an object of interest.
 How to determine the segmentation masks.
-We use a pre-trained segmentation model, i.e., SAM (Kirillov et al., 2023), for the segmentation task, which accepts bounding boxes as prompts.
-It produces a mask covering the most prominent object within a bounding box.
-We run non-maximum suppression on the masks produced by SAM to remove the masks corresponding to less certain bounding boxes that have large overlaps with other masks.
-The masks produced for each cropped image are then passed to the next stage.
-For GroundingDINO we set the bounding box threshold to a very low value to avoid missing any objects of interest, especially if the objects belong to a rare category that GroundingDINO has encountered less frequently during pre-training.
-As shown in Figure 4, a low threshold value can lead to false positives, making the number of masks unreliable for the counting problem.
-Nonetheless, the false positive masks have no negative effect on the object-aware division in the next stage.
-The only consequence is that, in addition to the target masks, the division paths will not cut through the false positive masks either.
-In other words, extra masks are of no concern as long as all the target masks are identified.
+We employ a pre-trained segmentation model, specifically SAM (Kirillov et al., 2023), for the segmentation task.
+This model accepts bounding boxes as prompts and generates masks covering the most prominent objects within these boxes.
+However, we do not use SAM’s output masks directly, as in crowded scenes or cases with multiple occlusions, the masks often overlap, making it difficult to identify reliable division paths.
+To address this issue, we apply several post-processing steps.
+First, we perform non-maximum suppression on SAM-generated masks to eliminate those with significant overlaps corresponding to less certain bounding boxes.
+Additionally, we apply an erosion function to the segmentations, ensuring adjacent masks maintain a minimum separation of two pixels.
+For a detailed analysis and visual examples of how these post-processing operations enhance robustness in crowded scenes and occlusion cases, see Appendix C.
+The final processed masks for each cropped area are then passed to the next stage of our pipeline.
+Robustness to the Accuracy of Area Detection and Target Segmentation Stages.
+Although we employ GroundingDINO—a state-of-the-art model for grounding and detection—it lacks the flexibility and robustness of LVLMs when generalizing to unseen data and complex concepts.
+Consequently, we minimizes dependence on the accuracy of the detection model used for area detection and target segmentation.
+Rather than prioritizing accuracy, we emphasize: i) not missing any region containing objects of interest during area detection, and ii) not missing segmentation masks for any objects of interest during target segmentation.
+These objectives are easily achieved by setting GroundingDINO’s detection threshold to a very low value in both stages.
+A low threshold may lead to false positives.
+In area detection, this could result in regions that contain no objects of interest.
+However, since counting is performed by the LVLM in the final stage, these regions will be ignored.
+For target segmentation, false positive masks might occur, but the consequence is that some irrelevant objects will be protected from being cut by the division lines, just as the objects of interest are.
+Furthermore, as we will demonstrate in subsequent sections, all masks are removed from the sub-images after division, ensuring no noise propagates from this stage to the LVLM-based counting stage.
+We demonstrate the effectiveness and robustness of our strategy in the early pipeline stages by showing that LVLM-Count significantly improves LVLMs’ performance on one of the most challenging counting datasets—the Penguin Dataset Penguin Research (2016), as detailed in Section 4.2.
+This dataset is particularly difficult due to frequent occlusion, camouflage, and complex backgrounds Arteta et al.
+(2016).
+Furthermore, in our ablation studies, we test a variant of our pipeline that removes the detection model entirely.
+Instead, SAM is configured to generate segmentation masks for all objects in the scene.
+As shown in Table 5, our method still enhances counting performance in this configuration, underscoring its robustness and minimal dependence on the detection model during the initial pipeline stages.
 3.3 Object-aware Division
@@
 How to determine the starting and ending points of the paths.
-We consider the first sub-step as a hyper-parameter, which can be set in various ways depending on the type of dataset.
-We utilize two approaches in our experiments.
-The first approach is unsupervised and non-parametric, where we use a clustering method to obtain the start and end points of the paths.
-We describe the clustering approach in more detail below.
-The second approach is to simply use a pre-determined number of equidistant points.
-This is helpful when we know a priori that objects are uniformly distributed across the image.
-In this case, we fix $k$ equidistant vertical and horizontal coordinates in the image as the points of the dividing paths for the entire dataset.
-In our experiments, we specify which approach we use for each dataset.
-Nonetheless, we will illustrate in the experimental results that even if we do not treat this as a hyperparameter and always use the unsupervised and non-parametric approach, LVLM-Count is quite effective.
+We utilize an unsupervised and non-parametric approach, to obtain the start and end points of the paths.
+A few pixels are sampled from each of the masks.
+To determine the location of the division paths on the $x$-axis, the samples taken from the masks are projected onto the $x$-axis.
+The projected points are automatically clustered using a non-parametric mean-shift algorithm.
+Once the clusters are identified, the point between the point with the highest $x$ value in one cluster and the point with the lowest $x$ value in the subsequent cluster is considered the $x$-coordinate of a division path.
+In effect, knowing the $x$-coordinate of a vertical path means that the coordinates of its endpoints are known.
+In particular, assuming height $h$ for an image crop, we consider $P_{s}=(x,0)$ and $P_{e}=(x,h)$ as the start and end points, respectively.
+Note that using this technique, we obtain the appropriate coordinates for the division paths, as well as the number of paths.
+For example, if there is only one cluster along the $x$-axis, no division is required, and if there are two clusters, one vertical path will divide the image into two parts.
+We illustrate this approach in Figure 5.
 Figure 5: Illustration of the unsupervised and non-parametric method to obtain the division points ($P^{1}_{s},P^{1}_{e}$), and ($P^{2}_{s},P^{2}_{e}$,).
-The object of interest is “horse”.
-Firstly, the target masks are produced.
-Then, a few pixels are sampled (shown as points inside the segmented objects) from the pixels composing each mask.
+A few pixels are sampled (shown as points inside the segmented objects) from the pixels composing target masks.
 The samples are projected onto the $x$-axis.
@@
 The point in the middle of two consecutive clusters is considered a vertical division point.
-The straight vertical lines are obtained using our path-finding method.
-Description of the unsuperivsed and non-parametric approach.
-A few pixels are sampled from each of the masks.
-To determine the points of the vertical division paths, the samples taken from the masks are projected onto the $x$-axis.
-The projected points are automatically clustered using a non-parametric mean-shift algorithm.
-Once the clusters are identified, the point between the point with the highest $x$ value in one cluster and the point with the lowest $x$ value in the subsequent cluster is considered the $x$-coordinate of a vertical division path.
-Using this technique, we obtain the appropriate coordinates for the vertical paths, as well as the number of paths.
-For example, if there is only one cluster, no vertical division is required, and if there are two clusters, one vertical path will divide the image into two parts.
-We illustrate this approach in Figure 5.
-Note that we chose the vertical axis purely for convenience in illustration.
-The same process can be applied along the $y$-axis to obtain the $y$-coordinates of the horizontal division paths.
+Blue lines are solely for illustration
 How to draw the paths.
-In effect, knowing the $x$-coordinate of a vertical path means that the coordinates of its endpoints are known.
-In particular, assuming height $h$ for an image crop, we consider $P_{s}=(x,0)$ and $P_{e}=(x,h)$ as the start and end points, respectively.
-In an ideal case where there are no masks in the path of a straight line connecting the two points, this line will be drawn by connecting all the pixels on the straight path.
+Previous step obtains the endpoints of the division paths.
+Assume $P_{s}=(x,0)$ and $P_{e}=(x,h)$ are the start and end points of a vertical division path, respectively.
+In an ideal case where there are no masks in the path of a straight line connecting the two points, this path will be drawn by connecting all the pixels on the straight line.
 However, there are potential masks that can be considered obstacles blocking the path.
@@
 Figure 6: Illustration of object-aware division.
-The masks for “donut” are obtained and turned into a black-and-white image.
-A dividing path is found by connecting $P_{s}$ to $P_{e}$ using the $A^{*}$ search algorithm in a graph that corresponds to the binary image, where the only nodes in the graph are white pixels, which are fully connected.
-The path is mapped back to the pixel domain.
+The masks are turned into a black-and-white image.
+A dividing path is found by connecting $P_{s}$ to $P_{e}$ using the $A^{*}$ search algorithm in a graph that corresponds to the binary image.
+The only nodes in the graph are white pixels of the black-and-white image, which are connected to all other white pixels in their $3\times 3$ neighborhood.
+The nodes and edges on the obtained paths have been colored red.
 3.4 Target Counting
-At this point, all the sub-images obtained from the cropped areas are gathered.
-Since these are small partitions of the original image, they might lack the desired resolution.
-Thus, using a super-resolution model can alleviate this problem for the sub-images.
-In this work, we use Real-ESRGAN (Wang et al., 2021).
+All the sub-images obtained from the cropped areas are gathered.
 Then, question $Q$ and each sub-image are given as input to an LVLM.
 At the end of the loop, the recorded numbers for the sub-images are aggregated to form the final answer.
-[^1] Note that using super-resolution before counting the objects in the sub-images is also the default behavior of our method.
-For a specific dataset where we observe that the resolution is preserved after division, we deactivate super-resolution, as the cost and usage of LVLM APIs is much higher.
-Finally, for images with a very large number of objects, sometimes LVLMs refuse to count, citing the large number.
-In those cases, an “estimate” prompt is given instead of a “count” prompt..
+For images with a very large number of objects, sometimes LVLMs refuse to count, citing the large number.
+In those cases, a new prompt requires the model to give the closest estimate of the number.
 4 Experiments
-In this section, we present the performance results of our method on a counting-specific dataset, an open-ended counting benchmark with two types of questions, simple and complex, and a challenging counting benchmark that we propose using emoji icons.
-We compare the results to state-of-the-art models which have been specifically trained on counting datasets, and we also compare to state-of-the-art models which have not been trained on counting datasets.
+In this section, we present the performance results of different LVLMs and their enhanced performance results using our method on a counting-specific dataset, a counting benchmark taken from a popular vision datasets, and a challenging counting benchmark that we propose using emoji icons.
+Additionally, we show the success of our design in enhancing counting performance on one of the most challenging counting datasets that features heavy occlusions and complex backgrounds.
 The code to reproduce the experiments can be found here: https://github.com/mrghofrani/lvlm-count.
@@
 The training set contains $3659$ images, with the validation and test sets containing $1286$ and $1190$ images, respectively.
-For each image in the test set, a single category name is given, and the expected output is the number of instances from the category.
-Open-ended Counting Benchmark.
-TallyQA (Acharya et al., 2019) is an open-ended counting dataset that includes complex counting questions involving relationships between objects, attribute identification, reasoning, and more.
-TallyQA is quite a large dataset, with the train set having $249,318$ questions and the test set having $22,991$ simple and $22,991$ complex counting questions.
-Please refer to Appendix C and Figure 9 for more information on simple and complex categorization of the questions.
-The number of objects in each image ranges from 00 to $15$.
-The dataset has a heavy bias towards a low number of objects (see Figure 8 in the Appendix).
-To alleviate this bias and create a benchmark for efficiently measuring the simple and complex open-ended capabilities of a counting model, we randomly sample $10$ questions per ground truth count.
-This results in $155$ simple and $149$ complex open-ended counting questions in total.
-It is important to note that the bias in TallyQA is pronounced; for most ground truth values greater than $10$, there are fewer than $10$ samples available in the entire test set.
+For each image in the test set, a single category name is given, and the expected output is the number of instances.
+PASCAL VOC Benchmark We build a counting benchmark from PASCAL VOC dataset (Everingham et al., 2015).
+Similar to Chattopadhyay et al.
+(2017), we choose PASCAL VOC 2007 among other variants.
+This variant contains a training set of 2501 images, a validation set of 2510 images, and a test set of 4952 images, with 20 object categories that remain consistent across the splits.
+Each image includes annotations for instances of the 20 object categories in the dataset.
+We first create 20 simple counting questions asking for the number of objects from each of the 20 categories for every image in the test set.
+Then, we randomly sample five questions from each ground truth count.
+This process resulted in 102 questions in total.
+Finally, we manually checked the ground truth counts and corrected them if required.
 Emoji-Count.
-Although TallyQA addresses the scarcity of complex counting questions in prior datasets to a certain degree, the range of target objects is limited, spanning from 00 to $15$.
-To our knowledge, no counting benchmark exists for large numbers involving complex reasoning.
+To our knowledge, no counting benchmark exists with large number of objects in a scene involving complex reasoning.
 To this end, we propose a challenging counting benchmark using emoji icons.
@@
 For each of the $82$ classes, an empty $1024\times 1024$ image is first created.
-This image is filled with six categories chosen randomly from the class, with each category having a random count between $30$ and $50$ in the image.
-We illustrate an example of this dataset in Figure 7.
-(a) Q: How many waning gibbous moons are there in the image? Answer: 13.
+This image is filled with up to six categories chosen randomly from the class, with each category having a random count between $30$ and $50$ in the image.
+For each image, we create questions that ask the number of instances of the available categories in the image.
+This results in 415 image-question pairs.
+We illustrate two examples of this dataset in Figure 7.
+(a) Q: How many waning gibbous moons are there in the image? Answer: 18.
 (b) Q: How many clocks at time “two-thirty” are there in the image? Answer: 15.
-Figure 7: Illustration of two challenging cases from Emoji-Count.
-Note that, for convenience in visualization, a smaller version of the real images is depicted here.
-The real images are $1024\times 1024$ and contain many more icons.
-In Figure 7(a), the class name is “Moon Phase”.
-To predict the correct answers for the question, the model needs accurate information about the different moon phases in addition to strong counting capability.
-In Figure 7(b), the class name is “Clock Time”.
-In this case, the model needs to be able to read the time shown on the clocks in addition to strong counting capability.
+Figure 7: Illustration of a smaller version of two challenging cases from Emoji-Count. In Figure 7(a), the class name is “Moon Phase”. In Figure 7(b), the class name is “Clock Time”.
+Penguin Benchmark.
+The challenging Penguin dataset Penguin Research (2016) consistently exhibits heavy occlusion and complex background patterns that can easily be mistaken for penguins Arteta et al.
+(2016).
+The test set of this dataset is quite large, containing thousands of images with penguin counts ranging from 0 to 213.
+To build a manageable benchmark, we randomly sample 100 images while preserving a balanced ground truth range.
+We refer the reader to Appendix D for more details about this benchmark.
 4.2 Results
-The following discusses the numerical results of our experiments with LVLM-Count on each benchmark described in Section 4.1.
-For visual examples of LVLM-Count’s performance on each benchmark, see Appendix M.
-Additionally, for the ablation study and experiments on the PASCAL VOC dataset (Everingham et al., 2015), see Appendix A and Appendix K, respectively.
+The following section presents the numerical results of our experiments with the base LVLMs and their corresponding LVLM-Count on each benchmark described in Section 4.1.
+For visual examples of LVLM-Count’s performance on these benchmarks, see Appendix B.
+The ablation study results are discussed in Appendix A.
+Additionally, Appendix K provides the performance of SOTA counting models on these benchmarks as a reference for evaluating the LVLMs’ performance.
+Also, an inference time analysis is provided in Appendix J.
 FSC-147.
-We compare the performance of LVLM-Count to an extensive list of state-of-the-art counting models on the test set of the FSC-147 dataset.
-We also include two baselines: i) taking the number of target segmentation masks as the final answer, and ii) giving the output of the target segmentation stage to base GPT-4o and asking it to count the masks.
-We run different experiments using GPT-4o, Gemini 1.5 Pro, and an open-source model Qwen2 VL 72B AWQ (Yang et al., 2024) as LVLMs.
-Since objects in the FSC-147 dataset are uniformly distributed across scenes, we deactivate the cluster-based method for finding division points and instead use two vertical and two horizontal object-aware division paths with equidistant start and end points.
-We also deactivate the super-resolution.
-We further run an experiment with GPT-4o as the LVLM that uses the cluster-based approach along both horizontal and vertical axes to automatically determine the start and end points of the division paths.
-This is done to measure the performance of our method when prior knowledge is not used to set the start and end points.
-The rest of the process follows the procedure described in Section 3.
-The expression $E$ used in different stages of our method is the category name provided in the test set.
-A simple $Q$ in the form of “How many $E$ are there?
-If you don’t see any, say zero.” is built and given as a text prompt to the LVLM during the counting stage.
+We evaluate the performance of various LVLMs and their enhanced performance using our method, LVLM-Count, on the test set of the FSC-147 dataset.
+We consider two baselines: (i) the baseline where the number of target segmentation masks is taken as the final answer, and (ii) the baseline where an image of the generated segmentation masks is provided to GPT-4o with a prompt to count the number of masks.
+Our experiments involve GPT-4o (a leading proprietary model), as well as two open-source models: Qwen2 VL 72B AWQ (Yang et al., 2024) and Gemma 3 27B Team et al.
+(2025).
+The expression $E$ used in different stages of our method corresponds to the category name provided in the test set.
+A simple query $Q$ in the form of “How many $E$ are there?” is constructed and used as the text prompt for the LVLM during the counting stage.
+In all experiments, the detection thresholds are set to $0.1$.
+The results are presented in Table 1, reporting the mean absolute error (MAE) and root mean square error (RMSE).
+Our findings indicate that LVLM-Count improves the performance of all three LVLMs in terms of MAE.
+Interestingly, while the base Qwen2 and Gemma 3 models are much less powerful than their commercial counterpart, they outperform the base GPT-4o when integrated into our pipeline.
+Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ Number of target segmentaion masks 44.14 - 154.39 - GPT-4o counting target segmentaion masks 38.45 - 156.11 - GPT-4o 25.17 - 137.86 - LVLM-Count (GPT-4o as LVLM) 18.09 $\downarrow 7.08$ 88.77 $\downarrow 49.09$ Gemma 3 27B 30.23 - 136.43 - LVLM-Count (Gemma 3 27B as LVLM) 19.70 $\downarrow 10.53$ 110.62 $\downarrow 25.81$ Qwen2 VL 72B AWQ 33.45 - 145.90 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 22.67 $\downarrow 10.78$ 124.73 $\downarrow 21.17$
+Table 1: Evaluation on the test set of the FSC-147 dataset.
+In all tables, the results for base LVLMs and LVLM-Count are reported over three trials.
+Also, columns marked with $\Delta$ show the performance difference between the LVLM-Count and the base LVLM it uses.
+Green indicates improvement, while red represents degradation.
+To see the measured accuracy metrics for this dataset and the subsequent benchmarks, refer to Appendix I.
+Additionally, MAE analysis across different intervals of ground truth values for FSC-147 is provided in Appendix H.
+PASCAL VOC Benchmark.
+We evaluate the performance of three LVLMs—GPT-4o, Qwen2, and Gemma 3—on this benchmark, both with and without LVLM-Count.
+As shown in Table 2, LVLM-Count consistently outperforms the base LVLMs.
+Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ Number of the target segmentation masks 4.03 - 7.47 GPT-4o counting target segmentaion masks 6.60 - 15.09 GPT4o 4.46 - 8.35 - LVLM-Count (GPT4o as LVLM) 3.31 $\downarrow$ 1.15 7.11 $\downarrow$ 1.24 Gemma 3 27B 3.37 - 7.49 - LVLM-Count (Gemma 3 27B as LVLM) 2.87 $\downarrow$ 0.5 5.84 $\downarrow$ 1.65 Qwen2 VL 72B AWQ 4.83 - 8.84 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 4.05 $\downarrow$ 0.78 7.72 $\downarrow$ 1.12
+Table 2: Evaluation on the PASCAL VOC counting benchmark.
+Emoji-Count.
+We evaluate the performance of the LVLMs and their performance using LVLM-Count on the Emoji-Count benchmark.
 The results are shown in Table 3.
-For this dataset, the mean absolute error (MAE) and root mean square error (RMSE) are reported.
-We observe that our method enhances the performance of all three LVLMs in terms of MAE.
-With our method, all three LVLMs outperform TFOC, which which is a training-free method.
-Interestingly, although the base Qwen2 VL 72B AWQ is not as powerful as its commercial counterpart, GPT-4o, it performs almost on par with GPT-4o when both use our pipeline.
-The results also show that even without leveraging prior information to set the start and end points of the division paths, our pipeline remains effective.
-Table 1: Evaluation of state-of-the-art models on the test set of the FSC-147 dataset.
-The column “Trained Model” indicates if a model has been trained on FSC-147.
-In all tables, the results for base LVLMs and LVLM-Count are reported over three trials.
-Columns marked with $\Delta$ show the improvement that LVLM-Count brings over the base LVLM that it uses.
-For accuracy metrics, refer to Table 9.
-Additionally, MAE analysis across different intervals of ground truth values is provided in Appendix H.
-Method Trained Model MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC (Shi et al., 2024) 24.79 - 137.15 - VLCounter (Kang et al., 2024) 17.05 - 106.16 - CounTX(Amini-Naieni et al., 2023) 15.88 - 106.29 - $\text{DAVE}_{\text{prm}}$ (Pelhan et al., 2024) 14.90 - 103.42 - CountGD (Amini-Naieni et al., 2024) 14.76 - 120.42 - GroundingREC (Dai et al., 2024) 10.12 - 107.19 - Number of target segmentaion masks 44.14 - 154.39 - Base GPT-4o counting the target segmentaion masks 38.45 - 38.45 - GPT-4o 23.75 - 137.39 - LVLM-Count (GPT-4o as LVLM, Cluster-based on both axes) 17.67 $\downarrow 6.08$ 90.61 $\downarrow 46.78$ LVLM-Count (GPT-4o as LVLM) 16.23 $\downarrow 7.52$ 76.09 $\downarrow 61.3$ Gemini 1.5 Pro 25.20 - 108.76 - LVLM-Count (Gemini 1.5 Pro as LVLM) 14.85 $\downarrow 10.35$ 85.60 $\downarrow 23.16$ Qwen2 VL 72B AWQ 33.45 - 145.90 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 17.77 $\downarrow 15.68$ 113.06 $\downarrow 32.84$
-Table 1 shows that, although we outperform models that have not been trained on FSC-147, the best-performing models are those that have been trained on this dataset.
-However, in subsequent experiments, we demonstrate that, while the best-performing models on FSC-147 show acceptable generalization on benchmarks they have not been trained on, they lack the generalization ability of our method on new datasets, and they are even outperformed by the base LVLM models.
-Open-ended Counting Benchmark.
-We evaluate the performance of LVLM-Count on the simple and complex questions in the open-ended counting benchmark mentioned above.
-For LVLM-Count, we compute the start and end points of the object-aware division paths using the unsupervised and parameter-free clustering approach described in Section 3.3 on $x$-axis.
-This approach is necessary because the images in this benchmark are diverse, and objects are not uniformly distributed across the image.
-Thus, using a fixed number of object-aware division paths with equidistant start and end points does not perform well.
-We also conduct an experiment using GPT-4o as the LVLM, where we use the clustering method along both axes to find the start and end points.
-Since these are special cases of VQA tasks, we report exact accuracy (EA), which is the standard performance metric for VQA, in addition to MAE and RMSE.
-We compare our method against the base GPT-4o, Gemini 1.5 Pro, and Qwen2 VL 72B models, as well as three of the best-performing trained models from Table 1, namely GroundingREC, CountGD, and $\text{DAVE}_{\text{prm}}$ [^2] GroundingREC in Table 2 does not use the same model weights as those trained for FSC-147.
-It was trained on the REC-8K dataset (Dai et al., 2024), which is specifically designed for referring expression counting, making it stronger for complex counting questions.
-CountGD and $\text{DAVE}_{\text{prm}}$ use the same model weights as those used for FSC-147, as the authors have not provided model weights trained on a dataset with referring expressions., and the only prior training-free model, TFOC.
-Note that due to extremely poor performance, we do not provide the original question to these models.
-Instead, we assist them with the $E$ extracted in our pipeline.
-The results are shown in Table 2.
-Our method improves both MAE and EA over the base GPT-4o model on both complex and simple benchmarks.
-In general, however, there is barely any improvement over base LVLMs on simple questions.
-This is because the questions are straightforward, and the ground truths are within [0, 15].
-Based on our observations from Figure 2(c) and Figure 18, we expect the base LVLMs to perform better in such cases.
-However, for complex questions, there is consistent improvement in EA across all three LVLMs.
-Interestingly, despite the considerable gap between the EA of the base Qwen2 VL 72B AWQ and base GPT-4o on complex questions, using LVLM-Count enables this open-source model to achieve a higher EA than the base GPT-4o.
-Moreover, observe that GroundingREC, CountGD, and $\text{DAVE}_{\text{prm}}$ are outperformed by the LVLM-based models on both simple and complex categories.
-This is because they have not been specifically trained on TallyQA.
-The performance gap is more pronounced on complex questions.
-Table 2: Evaluation of models on the TallyQA benchmark.
-Numbers in parentheses under the $\Delta$ signs show the performance difference between LVLM-Count and the base LVLM it uses.
+This is a challenging benchmark, as it requires understanding complex concepts.
+We observe that taking the number of masks as the final count performs particularly poorly, as for any object of interest in the image, the segmentation stage tends to segment all the objects and cannot distinguish between different icons.
+Although GPT-4o and Gemma 3 show reasonable performance, the other open-source model does not perform well.
+Nonetheless, the performance of all three base LVLMs is significantly enhanced by LVLM-Count, especially Qwen2, which performs almost on par with GPT-4o when LVLM-Count is used for it.
+Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ Number of the target segmentation masks 82.47 - 107.98 - GPT-4o counting target segmentaion masks 107.72 - 162.12 - GPT-4o 22.51 - 35.94 - LVLM-Count (GPT-4o as LVLM) 16.29 $\downarrow 6.22$ 32.47 $\downarrow 3.47$ Gemma 3 27B 21.43 - 24.15 - LVLM-Count (Gemma3 27B as LVLM) 15.76 $\downarrow 5.67$ 21.00 $\downarrow 3.15$ Qwen2 VL 72B AWQ 82.41 - 186.32 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 23.74 $\downarrow 58.67$ 40.04 $\downarrow 143.84$
+Table 3: Evaluation on the Emoji-Count benchmark.
+Penguin Benchmark.
+We report the counting performance on the Penguin benchmark in Table 4.
+We evaluate two variants of LVLM-Count.
+The primary variant uses GroundingDINO in both the area detection and target segmentation stages.
+The alternative variant does not use a detection model; instead, SAM is configured to segment all entities in the image.
+Both variants improve MAE across all three LVLMs, highlighting LVLM-Count’s robustness in scenarios with heavy occlusion and complex backgrounds—conditions that pose significant challenges for area detection and target segmentation.
+Additional details and a visual example showing the segmentation masks and division paths generated in both variants for a sample from the Penguin benchmark can be found in Appendix D.
+Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT4o 36.01 - 46.64 - LVLM-Count( Main variant, using GPT-4o ) 26.95 $\downarrow$ 9.06 39.32 $\downarrow$ 7.32 LVLM-Count(SAM-only variant, using GPT-4o) 28.79 $\downarrow$ 7.22 45.71 $\downarrow$ 0.93 Gemma 3 27B 49.87 - 61.06 - LVLM-Count( Main variant, using Gemma 3 ) 36.75 $\downarrow$ 13.12 47.24 $\downarrow$ 13.82 LVLM-Count( SAM-only variant, using Gemma 3 ) 41.97 $\downarrow$ 7.90 54.55 $\downarrow$ 6.51 Qwen2 VL 72B AWQ 44.96 - 66.78 - LVLM-Count( Main variant, using Qwen2 ) 28.67 $\downarrow$ 16.29 43.71 $\downarrow$ 23.07 LVLM-Count(SAM-only, using Qwen2 ) 38.29 $\downarrow$ 6.67 51.47 $\downarrow$ 15.31
+Table 4: Evaluation on the Penguin benchmark.
+Limitations
+In this work, we quantitatively evaluated the visual counting performance of several LVLMs on multiple datasets.
+More importantly, we introduced a simple yet effective baseline method, LVLM-Count, that enhances the visual counting capabilities of LVLMs across the evaluated benchmarks.
+However, like any method, LVLM-Count has limitations.
+One limitation arises in cases where sub-images contain no objects of interest: the LVLM may occasionally predict a non-zero value.
+This reflects a broader weakness in LVLMs, and addressing it will require targeted improvements to ensure their accurate zero prediction in such scenarios.
+Another limitation occurs with open-source models during the target counting stage.
+Ideally, the LVLM’s output for each sub-image should be a numerical value.
+While top proprietary LVLMs, such as GPT-4o, offer functionalities like JSON schema to enforce structured responses (e.g., ensuring a numerical output), open-source models typically lack such features.
+For these models, the prompt must include explicit instructions to format the response correctly.
+For instance, an instruction like "Place the final predicted number inside [[]]" enables the use of regex searches to extract the number for aggregation.
+Appendix Table of Contents
+[…]
+[sections]
+[…]
+[sections]l1
+Appendix A Ablation study
+We examine the effect of each stage in our method: (i) area detection and (ii) object-aware division using masks produced during target segmentation.
+We design different experiments to investigate each stage’s individual contribution, as well as their combined effect in our pipeline.
+To demonstrate our method’s minimal dependence on detection accuracy in the area detection and target segmentation stages, we conduct an experiment without area detection and without using GroundingDINO for segmentation masks.
+Instead, we configure SAM in “segment anything” mode, which generates masks for all entities in the scene.
+The results show that our method remains effective even without any detection model, confirming that LVLM-Count has minimal dependence on the accuracy of these initial stages.
+Additionally, we conduct an experiment where both area detection and segmentation stages are excluded.
+In this case, object-aware division cannot be performed, so we divide images using equidistant straight lines into subimages (which we call “naive division”).
+We also run another experiment with area detection but without segmentation, applying naive division to the detected areas.
+The results demonstrate that unlike object-aware division, the naive approach is unreliable due to repetitive counting of fragmented objects.
+We perform ablation studies using GPT-4o, conducting experiments on the FSC-147 test set. Table 5 presents the results of these ablation experiments.
+Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT-4o 25.17 - 137.86 - GPT-4o + Naive division 33.04 $\uparrow$ 7.87 116.83 $\downarrow$ 21.03 GPT-4o + Area Detection + Naive division 32.69 $\uparrow$ 7.52 102.41 $\downarrow$ 35.45 GPT-4o + Area detection 23.08 $\downarrow$ 2.09 120.06 $\downarrow$ 17.8 GPT-4o + Object-aware division (using SAM without GroundingDINO) 21.01 $\downarrow$ 4.16 135.04 $\downarrow$ 2.82 GPT-4o + Object-aware division 19.17 $\downarrow$ 6.00 120.61 $\downarrow$ 17.25 GPT-4o + Area detection + Object-aware division, (equiv.
+to LVLM-Count) 18.09 $\downarrow$ 7.08 88.77 $\downarrow$ 49.09
+Table 5: An ablation study of LVLM-Count on the FSC-147 test dataset.
+Columns marked with $\Delta$ show the performance difference between LVLM-Count and the base LVLM.
 Green indicates improvement, while red represents degradation.
-For more details about the simple and complex types of questions, please refer to Appendix C and Figure 9.
-For additional accuracy measures, refer to Table 10 and 11.
-Method Simple Questions Complex Questions EA (%$\uparrow$($\Delta$) MAE $\downarrow$($\Delta$) RMSE $\downarrow$($\Delta$) EA (%$\uparrow$($\Delta$) MAE $\downarrow$($\Delta$) RMSE $\downarrow$($\Delta$) TFOC (Shi et al., 2024) 6.45 (-) 8.28 (-) 14.79 (-) 1.34 (-) 12.41 (-) 22.80 (-) $\text{DAVE}_{\text{prm}}$ (Pelhan et al., 2024) 5.81 (-) 16.06 (-) 35.26 (-) 3.36 (-) 28.36 (-) 57.56 (-) GroundingREC (Dai et al., 2024) 23.87 (-) 2.90 (-) 4.71 (-) 16.78 (-) 5.83 (-) 10.13 (-) CountGD (Amini-Naieni et al., 2024) 41.94 (-) 2.37 (-) 4.56 (-) 5.37 (-) 9.78 (-) 17.21 (-) Number of the target segmentation masks 27.31 (-) 2.60 (-) 4.10 (-) 15.44 (-) 4.25 (-) 6.78 (-) Base GPT-4o counting the target segmentation masks 25.38 (-) 2.82 (-) 2.82 (-) 11.41 (-) 4.99 (-) 4.99 (-) GPT-4o 44.30 (-) 1.38 (-) 2.35 (-) 29.08 (-) 2.60 (-) 4.74 (-) LVLM-Count (GPT-4o as LVLM, Cluster-based on both axes) 41.29 ($\downarrow 3.01$) 1.82 ($\uparrow 0.44$) 3.33 ($\uparrow 0.98$) 28.41 ($\downarrow 0.67$) 3.18 ($\uparrow 0.58$) 8.91 ($\uparrow 4.17$) LVLM-Count (GPT-4o as LVLM, Cluster-based on $x$-axis) 44.73 ($\uparrow 0.43$) 1.18 ($\downarrow 0.20$) 2.06 ($\downarrow 0.29$) 34.68 ($\uparrow 5.6$) 2.28 ($\downarrow 0.32$) 4.18 ($\downarrow 0.56$) Gemini 1.5 Pro 47.10 (-) 1.08 (-) 1.87 (-) 25.73 (-) 2.13 (-) 3.36 (-) LVLM-Count (Gemini 1.5 Pro as LVLM, Cluster-based on $x$-axis) 45.16 ($\downarrow 1.94$) 1.43 ($\uparrow 0.35$) 4.72 ($\uparrow 2.85$) 26.62 ($\uparrow 0.89$) 2.79 ($\uparrow 0.66$) 4.70 ($\uparrow 1.34$) Qwen2 VL 72B AWQ 49.03 (-) 1.44 (-) 2.74 (-) 24.61 (-) 3.21 (-) 5.35 (-) LVLM-Count (Qwen2 VL 72B AWQ as LVLM, Cluster-based on $x$-axis) 41.72 ($\downarrow 7.31$) 1.66 ($\uparrow 0.22$) 3.41 ($\uparrow 0.67$) 30.65 ($\uparrow 6.04$) 2.47 ($\downarrow 0.74$) 4.35 ($\downarrow 1$)
-Emoji-Count.
-We evaluate the performance of LVLM-Count on the Emoji-Count benchmark.
-The results are shown in Table 3.
-For LVLM-Count, we deactivate the cluster-based division points and, similar to the FSC-147 experiment, use two vertical and horizontal object-aware division paths with equidistant start and end points.
-We also run an experiment with GPT-4o where LVLM-Count uses the clustering approach along both axes to determine the start and end points of the division paths.
-We report MAE and RMSE and compare against the base LVLMs, TFOC, $\text{DAVE}_{\text{prm}}$, CountGD, and GroundingREC.
-Note that none of the models have been exposed to this benchmark.
-Furthermore, it is a challenging benchmark as it requires understanding complex concepts.
-We observe that prior counting models perform poorly because, for any object of interest in the image, these models tend to count all the objects and cannot distinguish between different icons.
-Nonetheless, the base LVLMs show reasonable performance, but since the number of objects of interest is large in this dataset, all three base LVLMs are outperformed by LVLM-Count.
-The results for the clustering approach show that LVLM-Count is quite effective even when no prior knowledge is used to set the start and end points of the division paths.
-Table 3: Evaluation of state-of-the-art models on the Emoji-Count benchmark.
-Columns marked with $\Delta$ show the improvement LVLM-Count brings over the base LVLM it uses.
-Please refer to Table 12 for more accuracy metrics.
-Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC (Shi et al., 2024) 64.64 - 87.45 - $\text{DAVE}_{\text{prm}}$ (Pelhan et al., 2024) 198.99 - 208.08 - CountGD (Amini-Naieni et al., 2024) 137.93 - 156.80 - GroundingREC (Dai et al., 2024) 36.16 - 51.88 - Number of the target segmentation masks 82.47 - 107.98 - Base GPT-4o counting the target segmentation masks 107.72 - 162.12 - GPT-4o 22.51 - 35.94 - LVLM-Count (GPT-4o as LVLM, Cluster-based on both axes) 11.10 $\downarrow 11.41$ 24.23 $\downarrow 11.71$ LVLM-Count (GPT-4o as LVLM) 9.82 $\downarrow 12.69$ 19.46 $\downarrow 16.48$ Gemini 1.5 Pro 18.17 - 27.83 - LVLM-Count (Gemini 1.5 Pro as LVLM) 14.35 $\downarrow 3.82$ 23.42 $\downarrow 4.41$ Qwen2 VL 72B AWQ 82.41 - 186.32 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 20.67 $\downarrow 61.73$ 34.48 $\downarrow 151.84$
-5 Limitations and Future Work
-A limitation of LVLM-Count is the area detection stage.
-If the cropped areas do not provide enough context, performance may suffer.
-This is especially true for complex questions that require understanding the relationships between all the objects and the background in an image.
-This opens up opportunities for future work on improved area detection methods or the use of a context provider to complement the area detection stage.
-Another limitation arises with images containing thousands of objects.
-After one iteration of division, a significant number of objects may still remain in the sub-images.
-A potential solution is to apply additional iterations of division on the sub-images; however, the low resolution of these sub-images may make this infeasible.
-Developing a solution to maintain resolution is another direction for future work.
-Finally, in some cases, sub-images do not contain any objects of interest.
-The LVLM occasionally predicts a non-zero value in such instances.
-This is a weakness of LVLMs that requires special consideration to improve their performance.
-Appendix A Ablation Study
-We examine the effect of each of the following stages in our method: i) area detection and ii) object-aware division (note that the object-aware division necessitates the inclusion of target segmentation stage).
-The experiments are designed to investigate the effect of each stage individually, as well as when the stages are combined in our pipeline.
-Additionally, we run an experiment for a case where both stages of area-detection and target segmentation are excluded.
-In this case, the object-aware division can not be performed.
-Thus, images are divided by equidistant straight lines into subimages.
-We give the name of naive division to such an approach.
-Moreover, we run another experiment where area detection is in place but the target segmentation is excluded and naive division is applied on the detected areas We run the ablation scenarios for two LVLMs: GPT-4o, and Gemini 1.5 Pro.
-Furthermore, the effect of super resolution at the final stage of LVLM-Count is also investigated for the case that GPT-4o is the LVLM.
-For the experiments, we randomly sample $4$ images from each category in the FSC-147 test set ($29$ categories) and report the performance on the resulting subset containing $116$ samples.
-In Table 4 we show the results of the ablation experiments.
-Table 4: Ablation study for LVLM-Count on 116 samples from the FSC-147 test dataset.
-Columns marked with $\Delta$ show the performance difference between an entry and the base LVLM it uses.
-Green indicates improvement, while red represents degradation.
-Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT-4o 18.58 - 56.84 - GPT-4o + Naive division 33.95 $\uparrow 15.37$ 88.75 $\uparrow 31.91$ GPT-4o + Area Detection + Naive division 34.82 $\uparrow 16.24$ 73.19 $\uparrow 16.35$ GPT-4o + Area detection 19.27 $\uparrow 0.69$ 70.64 $\uparrow 13.8$ GPT-4o + Object-aware division 14.92 $\downarrow 3.66$ 31.56 $\downarrow 25.28$ GPT-4o + Area detection + Object-aware divsion (equiv.
-to LVLM-Count) 11.03 $\downarrow 7.55$ 21.02 $\downarrow 35.82$ GPT-4o + Area detection + Object-aware divsion + Super resolution 12.30 $\downarrow 6.28$ 28.98 $\downarrow 27.86$ Gemini 1.5 Pro 22.07 - 64.34 - Gemini 1.5 Pro + Naive division 41.31 $\uparrow 19.24$ 76.56 $\uparrow 12.22$ Gemini 1.5 Pro + Area Detection + Naive division 43.71 $\uparrow 21.64$ 75.47 $\uparrow 11.13$ Gemini 1.5 Pro + Area detection 23.04 $\uparrow 0.97$ 61.94 $\downarrow 2.4$ Gemini 1.5 Pro + Object-aware division 15.39 $\downarrow 6.68$ 38.15 $\downarrow 26.19$ Gemini 1.5 Pro + area detection + Object-aware division (equiv.
-to LVLM-Count) 13.64 $\downarrow 8.43$ 36.56 $\downarrow 27.78$
-Appendix B Bias to Small Numbers in Datasets and Performance of LVLMs
-In our experiments, LVLMs (GPT-4o) are able to make correct predictions when the number of items to be counted is small, but errors increase as the ground truth number grows.
-Although, we cannot be certain, one likely reason for this behavior in LVLMs is that, during training, the counting questions these models encounter are heavily biased toward small numbers.
-As an example, in Figure 8 we show the distribution of ’How many’ questions in some well-known VQA datasets.
-(a) TallyQA Dataset [Acharya et al., 2019]
-(b) VQAv2 Dataset [Goyal et al., 2017]
-Figure 8: Distribution of answers in two VQA datasets.
-Appendix C Definition and Example of Simple and Complex Counting Tasks
-Acharya et al.
-[2019] were among the first to formally categorize counting questions into simple and complex types.
-They applied a linguistic rule: first, they removed any substrings such as “…in the photo?” or “…in the image?”.
-Then, they used SpaCy to perform part-of-speech tagging on the remaining substring.
-They classified a question as simple if it contained only one noun, no adverbs, and no adjectives; otherwise, they deemed it complex.
-This rule classifies questions such as “How many dogs?” as simple and “How many brown dogs?” as complex.
-Following this rule, they built two splits for the TallyQA dataset: a simple split and a complex split.
-Since we sample our open-ended counting benchmark from the TallyQA simple and complex splits, we adopt the same classification criteria for this benchmark.
-In Figure 9, we provide examples of a simple and a complex VQA question from TallyQA.
-The VQA question in Figure 9(a) is selected from the simple split of TallyQA.
-It is a straightforward counting question, merely asking for all instances of the animal.
-Figure 9(b), on the other hand, is selected from the complex split of TallyQA.
-In addition to counting, it requires detecting context and distinguishing between consonants and vowels.
-(a) How many cows are visible?
-(b) How many consonants are there on the green sign?
-Figure 9: A simple (left image) and a complex (right image) question from TallyQA.
-Appendix D Incorrect Examples in the FSC-147 Dataset
-We observed some incorrect instances in the FSC-147 dataset.
-For example, see Figure 10.
-In these cases, the category names which are provided are incorrect.
-For these instances, the extensive knowledge embedded within the LVLMs employed in our approach proves to be a disadvantage.
-These models detect inconsistencies and provide a count of zero as the output, whereas other methods are misled by superficial similarities and mistakenly count the objects.
-A more thorough study is required to detect all the incorrect examples in FSC-147.
-(a) Category name in FSC-147: apples, Answer in FSC-147: $182$, Ground truth: 00.
-(b) Category name in FSC-147: sunglasses, Answer in FSC-147: $81$, Ground truth: 00.
-Figure 10: Two erroneous examples from FSC-147.
-Figure 10(a) shows a number of crabapples while the category name in the FSC-147 is apples.
-Figure 10(b) features a number of glasses while the category name in FSC-147 is sunglasses.
-Appendix E Alleviating the Problem of Inaccurate Detections at the Area Detection and Target Segmentation Stages of LVLM-Count
-One of the limitations of LVLM-Count is that if the area detection fails to detect a relevant area, the objects of interest in that area will not be counted.
-Another limitation arises when the target segmentation fails to segment all the instances of the target category of objects.
-In such a scenario, non-segmented objects of interest might be cut through by division lines, causing the predicted total number to increase due to being counted multiple times.
-One simple, yet quite effective, approach that we use to overcome this limitation to a significant degree is setting the detection threshold of GroundingDINO, which is used in both of the mentioned stages, to an extremely low value.
-This significantly reduces the chance of missing an area or object of interest, although it might lead to some false positives.
-However, note that false positives do not harm the performance of LVLM-Count, as their only effect is that the object-aware division lines avoid cutting through them.
-To demonstrate the effectiveness of this approach, we evaluate it on the Penguin dataset [Penguin Research, 2016].
+Appendix B Visual Examples of LVLM-Count’s Performance on the FSC-147 Dataset, PASCAL VOC, and Emoji-Count Benchmarks
+This section presents several visual examples showcasing the performance of LVLM-Count on the FSC-147 dataset , PASCAL VOC, and the Emoji-Count benchmarks.
+The LVLM used in the pipeline to generate these visual examples is GPT-4o.
+Figure 8 illustrates three examples of LVLM-Count’s performance on FSC-147.
+Additionally, Figure 9 includes an example from the PASCAL VOC benchmark.
+Finally, we present three visual examples from the Emoji-Count benchmark in Figure 10.
+In these visual examples, LVLM-Count achieves more accurate results compared to the base GPT-4o.
+Figure 8: Illustration of three examples of the performance of LVLM-Count on FSC-147.
+Top row: The object of interest is “strawberry”.
+Middle row: The object of interest is “hot air balloon”.
+Bottom row: The object of interest is “sheep”.
+Figure 9: An example of the performance of LVLM-Count on the PASCAL VOC benchmark. The object of interest is “bottle”.
+Figure 10: Three examples of the performance of LVLM-Count on the Emoji-Count benchmark.
+Top row: The object of interest is “right arrow curving left”.
+Middle row: The object of interest is “woman with red hair”.
+Bottom row: The object of interest is “hibiscus”.
+Appendix C Robustness to Crowded Scenes and Heavy Occlusion
+One of the key strengths of our approach is its robustness in crowded scenes and images with heavy occlusions.
+To achieve this robustness, the masks produced by SAM are not directly used for object-aware division.
+Instead, several important post-processing operations are applied beforehand.
+The most critical of these is non-maximum suppression, which removes lower-confidence masks that overlap with others.
+Additionally, an erosion function is applied to trim the outermost layer of each mask, ensuring at least two pixels of empty space between adjacent masks to allow for reliable placement of division lines.
+A polishing step further refines the masks by smoothing their surfaces and eliminating artifacts, preventing division lines from mistakenly passing through them.
+Figure 11 demonstrates the strong performance of our method on a randomly selected image from the web containing various types of flowers.
+This scene features extreme occlusion; nonetheless, our method effectively handles the division task.
+It is important to note that the strong performance of our method in this challenging scenario is not coincidental.
+Figure 12 illustrates the impact of our post-processing operations.
+The upper row displays the unprocessed SAM masks, while the bottom row shows the results after post-processing, which generates reliable paths for the division lines.
+Figure 11: An example of the performance of LVLM-Count on a random crowded image from the web, involving heavy occlusion. The object of interest is “flower”.
+Figure 12: The effect of post-processing on the masks produced by SAM in making LVLM-Count robust to crowded scenes with heavy occlusions. The object of interest is “flower”.
+To further demonstrate the effectiveness of our method in crowded scenes, we present Table 6, which compares the performance of different LVLMs on FSC-147 dataset samples where the ground truth count is 100 or higher.
+The table contrasts their baseline performance with their performance when using our LVLM-Count method.
+The results show significant improvements when LVLM-Count is employed.
+Furthermore, Figure 13 presents three visual examples of LVLM-Count’s performance on crowded images from FSC-147.
+Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT-4o 89.22 - 291.34 - LVLM-Count (GPT-4o as LVLM) 57.87 $\downarrow 31.35$ 185.29 $\downarrow 106.05$ Gemma 3 27B 125.69 - 330.38 - LVLM-Count (Gemma 3 27B as LVLM) 71.68 $\downarrow 54.01$ 251.48 $\downarrow 78.9$ Qwen2 VL 72B AWQ 102.64 - 302.23 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 79.66 $\downarrow 22.98$ 289.47 $\downarrow 12.76$
+Table 6: Evaluation on samples with ground truth counts equal to or higher than 100 from the test set of the FSC-147 dataset. Showing the effectiveness of our method for crowded scenes.
+Figure 13: Visual examples of the performance of LVLM-Count on three crowded scenes from the FSC-147 dataset.
+Appendix D Additional Information and Visual Examples from the Penguin Benchmark
+To demonstrate the effectiveness of our strategy of setting a low detection threshold to overcome challenges in area detection and target segmentation stages, we evaluate it on a benchmark taken from the Penguin dataset [Penguin Research, 2016].
 The goal in this dataset is to count penguins in images.
-The challenging Penguin dataset consistently exhibits heavy occlusion and complex background patterns that can easily be mistaken for penguins Arteta et al.
+This is challenging since images in the dataset consistently exhibit heavy occlusion and complex background patterns that can easily be mistaken for penguins Arteta et al.
 [2016].
-This dataset consists of two splits: the mixed-site split, in which images from the same camera can appear in both the training and testing sets, and the separate-site split, in which images in each set strictly belong to different cameras.
+This dataset consists of two splits: i) the mixed-site split, in which images from the same camera can appear in both the training and testing sets, and ii) the separate-site split, in which images in each set strictly belong to different cameras.
 Images in this dataset are annotated by multiple annotators, where each annotator might identify a different number of penguins due to the challenges in locating them within the images.
@@
 [2016], we take the maximum number of penguins among the annotations as the ground truth and calculate MAE and RMSE with respect to this value.
-For more details about the dataset and the calculation of metrics, we refer the reader to Arteta et al.
+For additional details regarding the dataset and metric calculations, we direct readers to Arteta et al.
 [2016].
-Since both splits are very large, we randomly select 100 samples from each split.
-To preserve balance, the probability of selecting an image with a certain ground-truth annotation number is inversely proportional to the frequency of that annotation number in the entire split, excluding images with zero penguins.
-We set the area detection and target segmentation thresholds to 0.5, and then we run them with both thresholds set to 0.1.
-The results are presented in Table 5 and Table 6.
-We observe that for both splits, when the detection threshold is high, the MAE generally increases to a value higher than the MAE of the base LVLMs.
-However, setting the thresholds to a low value improves the MAE for all three LVLMs.
-Table 5: Penguin dataset - Mixed sites.
-Columns marked with $\Delta$ show the performance difference between LVLM-Count and the base LVLM it uses.
-Green indicates improvement, while red represents degradation.
-Method MAE (Max)$\downarrow$ $\Delta$ RMSE (Max)$\downarrow$ $\Delta$ GPT-4o 26.45 - 37.53 - LVLM-Count(GPT-4o - Thredshold=0.5) 36.06 $\uparrow$ 9.61 49.23 $\uparrow$ 11.7 LVLM-Count(GPT-4o - Thredshold=0.1) 23.54 $\downarrow$ 2.91 42.24 $\uparrow$ 4.71 Gemini 1.5 Pro 37.23 - 49.15 - LVLM-Count(Gemini 1.5 Pro - Threshold=0.5) 42.42 $\uparrow$ 5.19 50.39 $\uparrow$ 1.24 LVLM-Count(Gemini 1.5 Pro - Threshold=0.1) 20.09 $\downarrow$ 17.14 29.95 $\downarrow$ 19.2 Qwen2 VL 72B AWQ 34.25 - 55.70 - LVLM-Count(Qwen2 VL 72B AWQ - Threshold=0.5) 37.13 $\uparrow$ 2.87 53.71 $\downarrow$ 1.99 LVLM-Count(Qwen2 VL 72B AWQ - Threshold=0.1) 18.22 $\downarrow$ 16.03 28.76 $\downarrow$ 26.94
-Table 6: Penguin dataset - Separate sites.
-Columns marked with $\Delta$ show the performance difference between LVLM-Count and the base LVLM it uses.
-Green indicates improvement, while red represents degradation.
-Method MAE (Max)$\downarrow$ $\Delta$ RMSE (Max)$\downarrow$ $\Delta$ GPT4o 36.01 - 46.64 - LVLM-Count(GPT-4o - Thredshold=0.5) 43.25 $\uparrow$ 7.24 52.17 $\uparrow$ 5.53 LVLM-Count(GPT-4o - Thredshold=0.1) 27.11 $\downarrow$ 8.9 36.72 $\downarrow$ 9.92 Gemini 1.5 Pro 46.67 - 57.26 - LVLM-Count(Gemini 1.5 Pro - Threshold=0.5) 33.59 $\uparrow$ 13.08 46.23 $\uparrow$ 11.03 LVLM-Count(Gemini 1.5 Pro - Threshold=0.1) 28.50 $\downarrow$ 18.17 41.39 $\downarrow$ 15.87 Qwen2 VL 72B AWQ 44.96 - 66.78 - LVLM-Count(Qwen2 VL 72B AWQ - Threshold=0.5) 51.17 $\uparrow$ 6.21 75.49 $\uparrow$ 8.71 LVLM-Count(Qwen2 VL 72B AWQ - Threshold=0.1) 24.83 $\downarrow$ 20.13 38.92 $\downarrow$ 27.86
-In order to also provide a visual insight into this limitation, we chose an image from the dataset where distinguishing penguins from the background features is challenging for GroundingDINO.
-Figure 11 shows the performance of our method on this image, with the area detection and target segmentation thresholds both set to 0.5.
-We can observe that the failure of the target segmentation stage in segmenting all instances of penguins in Figure 11(a) has resulted in division paths cutting through several penguins inside the areas surrounded by yellow shapes in Figure 11(b).
-Consequently, these penguins are counted twice, causing a larger error.
-On the other hand, Figure 12 shows the performance of LVLM-Count when the area detection and target segmentation thresholds are set to 0.1.
-Figure 12(a) illustrates that, while a low threshold has helped to segment all instances of the penguins, it has also caused some false positive areas to be segmented.
-Nonetheless, it can be observed in Figure 12(b) that the only effect of false positive segmentation masks is that the division paths avoid cutting through those regions as well.
-Note that for both images, the cluster-based approach was used to automatically find the start and end points of the division paths based on the arrangement of the segmentation masks in the scene.
-Comparing Figure 11 and Figure 12, we can clearly see that reducing the detection threshold is a quite successful technique to overcome limitations in the initial stages such as target segmentation.
-(a) With a threshold of 0.5, target segmentation stage has failed to segment all penguins
-(b) Areas surrounded by yellow shapes contain several penguins bisected by the dividing lines.
-Figure 11: In this figure, the detection threshold in the target segmentation stage is set to 0.5.
-For this example, the cluster-based approach has been used to find the start and end points of the division paths.
-Ground truth (Max) = 106, LVLM-Count prediction = 131.
-(a) With a low threshold such as 0.1, some false positive areas have also been segmented.
-(b) The division paths avoid cutting through the false positive areas as well.
-Figure 12: In this figure, the detection threshold in the target segmentation stage is set to 0.1.
-For this example, a cluster-based approach was used to find the start and end points of the division paths.
-Ground truth (Max) = 106, LVLM-Count prediction = 110
-Additionally, Figure 13 shows how using a low detection threshold can help alleviate the limitations in the area detection stage for a challenging image taken in unfavorable weather conditions.
-Figure 13(a) shows the output of the area detection stage when the threshold is set to 0.5.
-The output contains two detected areas that, in combination, fail to cover all the penguins, resulting in error.
-However, Figure 13(b) shows the output of the area detection stage when the threshold is set to 0.1, successfully giving the area containing all the penguins.
-(a) The area detection threshold set to 0.5
-(b) The area detection threshold set to 0.1
-Figure 13: This figure shows that setting the detection threshold in the area detection stage to a very low value (Figure 13(b)) significantly alleviates the area detection limitations at the area detection stage.
-Appendix F LVLM-Count’s Power in Handling Multiple Object Categories in the Same Image
-LVLM-Count is a highly effective method for handling counting tasks that involve multiple objects in the same image.
-Its strength in such scenarios stems from the capabilities of LVLMs to answer numerous visual questions about an image and its objects.
-Depending on the given text prompt, it can count instances of a single object category among others or instances of multiple object categories simultaneously.
-In this section, we demonstrate how LVLM-Count performs in counting different objects of interest, determined simply by a prompt, using an image with multiple object categories.
-The image in Figure 14 contains three object categories: person, cow, and horse.
-In the top row, the object of interest is “cow.” We prompt LVLM-Count to count the cows.
-First, the masks are produced through the initial stages of our pipeline, and then the cluster-based approach is used to automatically determine the start and end points of the division paths.
-It can be observed that horses have also been masked as cows.
-Nonetheless, this does not negatively impact the final answer; it merely causes the division lines to avoid cutting through the horses as well.
-The counting in LVLM-Count is performed by an LVLM (GPT-4o in this figure) and does not rely on the masks.
-We observe that GPT-4o successfully counts the number of cows in the resulting subimages, leading to the correct final answer.
-In the middle row of Figure 14, the object of interest is “person.” LVLM-Count again successfully counts the number of people accurately.
-A more interesting case is the bottom row of Figure 14, where both cows and persons are objects of interest.
-We prompt LVLM-Count to count the number of “cows and persons.” Similar to the first row of the figure, there are false positive masks here as well.
-However, LVLM-Count successfully counts the number of instances from both categories combined since the counting is ultimately performed by the LVLM.
-Note that the number of objects in this image is limited, and GPT-4o might answer these questions correctly without the need for the LVLM-Count pipeline.
-This image has been chosen to illustrate LVLM-Count’s power in handling multiple objects in a counting task rather than for comparison with the baseline LVLM.
-Figure 14: Illustration of the ability of LVLM-Count in counting an object of interest determined my a prompt when multiple object categories exist in a single image.
-Top row: Object of interest is “person”.
-Middle row: Object of interest is “cow”.
-Bottom row: Object of interest is “person and cow”
-Appendix G Real-world Application of LVLM-Count
+We construct our benchmark using the separate-site split.
+Given the dataset’s substantial size, which comprises tens of thousands of images, we randomly select 100 samples from the chosen split.
+To ensure balance, the sampling probability for an image with a specific ground-truth annotation count is inversely proportional to the frequency of that count in the entire split.
+We evaluate two variants of our pipeline.
+The main variant employs GroundingDINO with a low threshold for area detection and target segmentation, while the alternative variant does not use GroundingDINO at all.
+In this variant, masks are obtained by configuring SAM to segment any entity in the image.
+The results presented in Table 4 show that both variants significantly improve the counting performance of the LVLMs, demonstrating the robustness of the initial stages in our pipeline.
+Figure 14 demonstrates the segmentation masks and division paths for the two variants on a sample from the Penguin benchmark.
+In Figure 14 (a), GroundingDINO with a low detection threshold is employed.
+It correctly identifies all penguin instances, though there are some false positives.
+These detections are subsequently segmented by SAM.
+Figure 14 (c) depicts the division paths found using these masks.
+In Figure 14 (b), GroundingDINO is not used, and SAM is configured to segment all entities in the scene.
+As a result, in addition to the penguins, a large number of other objects are segmented.
+Nevertheless, Figure 14 (d) shows that the division paths derived from these masks successfully partition the image without cutting any penguins.
+Note that some excessively large masks generated by SAM have been removed due to post-processing in our pipeline.
+Figure 14: Illustration of robustness in detection during the initial pipeline stages: an example from the challenging Penguin benchmark.
+(a) GroundingDINO with a low threshold successfully detects all penguins, though a few false positives remain.
+These detections are then segmented by SAM.
+(b) Masks produced without a detection model: Here, SAM is set to segment all entities in the image.
+Excessively large masks are filtered out during post-processing in our pipeline.
+(c) Division paths are derived using the masks generated by GroundingDINO and SAM.
+(d) Division paths based on SAM-only segmentation effectively partition the image without cutting through any penguins.
+Note that the slight difference in the area of the images on the left is due to the existence of area detection stage.
+Appendix E Real-world Application of LVLM-Count
 As stated in Section 1, counting has numerous real-world applications, including but not limited to biology, health, industry, warehousing, and environmental monitoring.
@@
 The image is sourced from [Penguin Research, 2016], an initiative to monitor the penguin population in Antarctica, with the objects of interest being “penguins.”
+Appendix F LVLM-Count’s Power in Handling Multiple Object Categories in the Same Image
+LVLM-Count is a highly effective method for handling counting tasks that involve multiple objects in the same image.
+Its strength in such scenarios stems from the capabilities of LVLMs to answer numerous visual questions about an image and its objects.
+Depending on the given text prompt, it can count instances of a single object category among others or instances of multiple object categories simultaneously.
+In this section, we demonstrate how LVLM-Count performs in counting different objects of interest, determined simply by a prompt, using an image with multiple object categories.
+The image in Figure 18 contains three object categories: person, cow, and horse.
+In the top row, the object of interest is “cow.” We prompt LVLM-Count to count the cows.
+First, the masks are produced through the initial stages of our pipeline, and then the cluster-based approach is used to automatically determine the start and end points of the division paths.
+It can be observed that horses have also been masked as cows.
+Nonetheless, this does not negatively impact the final answer; it merely causes the division lines to avoid cutting through the horses as well.
+The counting in LVLM-Count is performed by an LVLM (GPT-4o in this figure) and does not rely on the masks.
+We observe that GPT-4o successfully counts the number of cows in the resulting subimages, leading to the correct final answer.
+In the middle row of Figure 18, the object of interest is “person.” LVLM-Count again successfully counts the number of people accurately.
+A more interesting case is the bottom row of Figure 18, where both cows and persons are objects of interest.
+We prompt LVLM-Count to count the number of “cows and persons.” Similar to the first row of the figure, there are false positive masks here as well.
+However, LVLM-Count successfully counts the number of instances from both categories combined since the counting is ultimately performed by the LVLM.
+Note that the number of objects in this image is limited, and GPT-4o might answer these questions correctly without the need for the LVLM-Count pipeline.
+This image has been chosen to illustrate LVLM-Count’s power in handling multiple objects in a counting task rather than for comparison with the baseline LVLM.
+Figure 18: Illustration of the ability of LVLM-Count in counting an object of interest determined by a prompt when multiple object categories exist in a single image.
+Top row: Object of interest is “cow”.
+Middle row: Object of interest is “person”.
+Bottom row: Object of interest is “person and cow”
+Appendix G False Positive Masks at the Target Segmentation Stage
+One of the reasons we task an LVLM to count the objects in the subimages instead of using the number of generated masks at the target segmentation stage as the final count of the objects of interest is the existence of false positive masks.
+The GroundingDINO model is responsible for detecting the objects of interest, determined by expression $E$, and passing the output bounding boxes to SAM for producing segmentation masks.
+Nonetheless, GroundingDINO is not as strong as an LVLM in understanding expressions extracted from complex questions.
+Thus, it often returns bounding boxes for all instances of the object category mentioned in the expression, even if those instances do not satisfy other conditions in the expression.
+For example, in the top row of Figure 19, $E=$brown egg”.
+However, all the eggs have been segmented regardless of their color.
+Thus, counting the masks results in a significant error.
+Interestingly, as we can see, the false positive masks do not negatively affect LVLM-Count’s final answer, as the counting is done by an LVLM at the final stage, which is much stronger than GroundingDINO at understanding referring expressions.
+The only effect is that the white eggs have not been cut through by the division lines either.
+In the bottom row, we have chosen an image from the challenging Emoji-Count benchmark.
+The image contains icons, all of which have an arrow but point in different directions.
+However, the objects of interest are only “right arrows curving left.” Similar to the eggs example, taking the masks used for object-aware division results in a significant error.
+Figure 19: Top row: The object of interest is “brown egg.” However, all the eggs have been segmented because of the limitation of the GroundingDINO model in understanding complex referring expressions.
+Regardless, LVLM-Count provides a significantly more accurate number compared to the number of masks.
+Bottom row: The object of interest is “right arrows curving left.” Similar to the image of the eggs, counting the number of masks results in a very large error, while LVLM-Count provides a much more accurate number.
 Appendix H Performance Analysis of LVLM-Count for different ground truth ranges on FSC-147 dataset
-To further investigate the performance of our pipeline, we divide the ground truth values in the FSC-147 test set into intervals and plot the MAE for the base GPT-4o and Gemini 1.5 Pro models, alongside the results from LVLM-Count using each model, as shown in Figure 18.
-The first interval contains relatively small ground truth values, a range where LVLMs already perform well.
+To further investigate the performance of our pipeline, we divide the ground truth values in the FSC-147 test set into intervals and plot the MAE for the base GPT-4o, Qwen2, and Gemma 3 models alongside the results from LVLM-Count using each model, as shown in Figure 20.
+The first interval contains relatively small ground truth values, a range in which LVLMs already perform well.
 As the ground truth values increase, the base models exhibit increasingly larger errors compared to LVLM-Count, with the margin growing rapidly.
@@
 (a) LVLM-Count using GPT-4o
-(b) LVLM-Count using Gemini 1.5 Pro
-Figure 18: Performance analysis of our method, LVLM-Count, on the FSC-147 test set using GPT-4o (Figure 18(a)) and Gemini 1.5 Pro (Figure 18(b)).
-In the first interval, both base LVLMs exhibit a lower MAE.
-However, in intervals with higher ground truth values, LVLM-Count achieves a lower MAE compared to the base LVLMs, and this difference increases rapidly.
-Appendix I Illustration of the Workflow for the Zebra Image in Figure 1
-In this paper, we choose the most appropriate images for simple and understandable illustrations of the execution of each stage in our pipeline.
-Inevitably, this led to the selection of different images for each stage.
-However, in this section, for the sake of consistency, we demonstrate the same concepts illustrated in Figures 3, 4, 5, and 6 for the zebra image used in Figure 1.
+(b) LVLM-Count using Gemma 3
+(c) LVLM-Count using Qwen2
+Figure 20: Performance analysis of our method, LVLM-Count, on the FSC-147 test set using GPT-4o (Figure 20(a)), Gemma 3 (Figure 20(b)), and Qwen2 (Figure 20(c)).
+In the first interval, all base LVLMs exhibit a lower MAE.
+However, in intervals with higher ground truth values, LVLM-Count achieves a lower MAE compared to the base LVLMs.
+Note that as the ground truth values grow higher, the margin of improvement increases rapidly.
+Appendix I Report of Various Accuracy Metrics for the Performance of LVLM-Count on the FSC-147 Dataset, PASCAL VOC, Emoji-Count, and Penguin Benchmarks
+This section presents various accuracy measures for the experiments reported in Tables 1, 2, 3, and 4.
+The accuracy metrics are defined in Table 7.
+Following these definitions, Table 8 shows the measured accuracies for the FSC-147 test set.
+Similarly, Tables 9, 10, and 11 present the accuracy metrics for the PASCAL VOC, Emoji-Count, and Penguin benchmarks, respectively.
+The results in the tables demonstrate that LVLM-Count generally achieves a higher accuracy than the base LVLM it employs.
+Metric Definition Acc Percentage of answers such that $answer=GT$ Acc±$\pm 1$ Percentage of answers such that $|answer-GT|\leq 1$ Acc±$\pm 3$ Percentage of answers such that $|answer-GT|\leq 3$ Acc±$\pm 5$ Percentage of answers such that $|answer-GT|\leq 5$ Acc±$\pm 10$ Percentage of answers such that $|answer-GT|\leq 10$
+Table 7: Definitions of Various Accuracy Metrics. GT denotes the ground truth number.
+Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) GPT4o 12.24 26.22 42.10 52.41 66.58 LVLM-Count (GPT4o as LVLM) 14.45 ($\uparrow$) 28.99 ($\uparrow$) 47.25 ($\uparrow$) 57.03 ($\uparrow$) 70.45 ($\uparrow$) Gemma 3 27B 8.71 18.94 32.55 41.57 55.15 LVLM-Count (Gemma 3 as LVLM) 11.37 ($\uparrow$) 25.46 ($\uparrow$) 41.51 ($\uparrow$) 52.07 ($\uparrow$) 67.09 ($\uparrow$) Qwen2 VL 72B AWQ 9.19 20.81 36.83 46.95 62.07 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 9.61 ($\uparrow$) 23.95 ($\uparrow$) 43.05 ($\uparrow$) 53.78 ($\uparrow$) 69.08 ($\uparrow$)
+Table 8: FSC-147 Dataset.
+A ($\uparrow$) next to the measured accuracies for LVLM-Count indicates improvement over the base LVLM it uses, while a ($\downarrow$) indicates degradation compared to the corresponding base LVLM.
+Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) GPT4o 30.39 47.06 61.11 71.90 89.87 LVLM-Count (GPT4o as LVLM) 32.35 ($\uparrow$) 51.96 ($\uparrow$) 73.20 ($\uparrow$) 82.35 ($\uparrow$) 93.46 ($\uparrow$) Gemma 3 27B 29.41 54.58 75.49 85.62 92.16 LVLM-Count (Gemma 3 27B as LVLM) 30.39 ($\uparrow$) 50.98 ($\downarrow$) 78.43 ($\uparrow$) 85.62 (-) 94.77 ($\uparrow$) Qwen2 VL 72B AWQ 24.18 42.48 60.13 72.22 86.60 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 28.76 ($\uparrow$) 46.73 ($\uparrow$) 67.32 ($\uparrow$) 78.43 ($\uparrow$) 89.87 ($\uparrow$)
+Table 9: PASCAL VOC Benchmark.
+A ($\uparrow$) next to the measured accuracies for LVLM-Count indicates improvement over the base LVLM it uses, while a ($\downarrow$) indicates degradation compared to the corresponding base LVLM.
+Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) GPT4o 1.85 5.54 13.73 21.12 43.94 LVLM-Count (GPT4o as LVLM) 2.81 ($\uparrow$) 9.88 ($\uparrow$) 23.53 ($\uparrow$) 36.22 ($\uparrow$) 60.24 ($\uparrow$) Gemma 3 27B 0.80 1.77 4.10 7.07 16.14 LVLM-Count (Gemma 3 27B as LVLM) 2.49 ($\uparrow$) 5.30 ($\uparrow$) 13.33 ($\uparrow$) 19.92 ($\uparrow$) 38.63 ($\uparrow$) Qwen2 VL 72B AWQ 0.88 2.89 7.55 11.41 19.76 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 2.33 ($\uparrow$) 6.75 ($\uparrow$) 16.22 ($\uparrow$) 25.70 ($\uparrow$) 43.86 ($\uparrow$)
+Table 10: Emoji-Count Benchmark.
+A ($\uparrow$) next to the measured accuracies for LVLM-Count indicates improvement over the base LVLM it uses, while a ($\downarrow$) indicates degradation compared to the corresponding base LVLM.
+Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) GPT4o 2.00 3.67 8.33 12.00 22.00 LVLM-Count (GPT4o as LVLM) 2.67 ($\uparrow$) 6.33 ($\uparrow$) 14.33 ($\uparrow$) 18.67 ($\uparrow$) 33.00 ($\uparrow$) Gemma 3 27B 1.33 1.67 7.00 8.67 16.00 LVLM-Count (Gemma 3 27B as LVLM) 1.00 ($\downarrow$) 3.67 ($\uparrow$) 9.33 ($\uparrow$) 12.67 ($\uparrow$) 23.67 ($\uparrow$) Qwen2 VL 72B AWQ 1.67 2.00 6.33 10.33 16.67 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 1.67 (-) 5.33 ($\uparrow$) 11.67 ($\uparrow$) 18.00 ($\uparrow$) 31.00 ($\uparrow$)
+Table 11: Penguin Benchmark.
+A ($\uparrow$) next to the measured accuracies for LVLM-Count indicates improvement over the base LVLM it uses, while a ($\downarrow$) indicates degradation compared to the corresponding base LVLM.
+Appendix J Inference Time Comparison
+LVLM-Count is a pipeline consisting of multiple stages.
+However, in the following section, we will demonstrate that the dominant portion of time during inference with a GPT-4o model is spent querying the LVLM to count objects, while other stages are negligible by comparison.
+Consequently, LVLM-Count can achieve inference times reasonably close to that of the base LVLM used.
+In the four stages of the pipeline, the following main steps are taken: making a call to an LLM/LVLM to extract the object of interest from $Q$, calling GroundingDINO for area detection, calling GroundingDINO and SAM for target segmentation, performing some post-processing on the masks, running the mean-shift clustering algorithm, running the $A^{*}$ search, and querying the LVLM for the sub-images.
+From the listed steps, the running time of the post-processing on masks is negligible.
+The time required for the rest of the steps in an optimized pipeline is demonstrated in Table 12.
+To obtain the inference time for each step, we evaluated it on 1,000 images from the FSC-147 dataset and calculated the average.
+footnotetext: This is for a case where sub-image queries to the LVLM are made in parallel
+Step Extract $E$ from $Q$ Area Detection Target Segmentation Clustering $A^{*}$ Sub-image Analysis Total Model Called GPT-4o GroundingDINO GroundingDINO SAM - - GPT-4o Time 0.56 0.10 0.10 0.33 0.24 0.03 1.65J 3.03
+Table 12: Inference time breakdown of the LVLM-Count pipeline (Time in seconds)
+As shown in Table 12, the dominant term in the inference time comes from calling GPT-4o to count objects of interest in the sub-images, which occupies more than half of the total inference time.
+However, a similar step occurs when querying the base LVLM to count objects of interest in the main image.
+The additional steps present in LVLM-Count but absent in the base LVLM constitute a smaller portion of the time.
+Table 13 presents a comparison between the total inference time of LVLM-Count and the base GPT-4o.
+This increase in inference time may be acceptable for most applications, given the substantial improvements in counting accuracy when using LVLM-Count.
+Method Inference (s)
+Base GPT4o 1.68
+LVLM-Count (Using GPT-4o) 3.03
+Table 13: Comparison of the inference time between the base GPT-4o and LVLM-Count
+Appendix K Comparison of Base LVLMs and LVLM-Count’s Counting Performance with SOTA Counting Models as a Reference Point
+LVLM-Count is designed to enhance the counting performance of LVLMs.
+Tables 1, 2, 3, and 4 in the main paper present extensive experimental results demonstrating LVLM-Count’s accuracy improvements over the base LVLMs.
+In this section, we include the results of state-of-the-art (SOTA) dedicated counting models for the experiments reported in the main paper.
+Specifically, we compare the performance of GroundingREC Dai et al.
+[2024], CountGD Amini-Naieni et al.
+[2024], and $\text{DAVE}_{\text{prm}}$ Pelhan et al.
+[2024], as well as TFOC Shi et al.
+[2024].
+The first three models have counting-specific training, while the latter, though a dedicated counting model, does not require counting-specific training.
+Reporting these results provides additional insights into the progress of counting performance in LVLMs and their enhanced performance with LVLM-Count.
+Table 14 presents the results of dedicated counting models alongside LVLMs and their corresponding LVLM-Count variants on the FSC-147 test set.
+Note that GroundingREC, CountGD, and $\text{DAVE}_{\text{prm}}$ are the top three highest-performing text-based trained models on FSC-147 in the literature.
+Moreover, TFOC is the best-performing training-free model on FSC-147.
+We observe that, in general, the dedicated counting models (henceforth referred to simply as counting models) outperform the base LVLMs.
+However, the use of LVLM-Count significantly narrows this performance gap.
+Method Trained Model MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC [Shi et al., 2024] 24.79 - 137.15 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 14.90 - 103.42 - CountGD [Amini-Naieni et al., 2024] 14.76 - 120.42 - GroundingREC [Dai et al., 2024] 10.12 - 107.19 - GPT-4o 25.17 - 137.86 - LVLM-Count (GPT-4o as LVLM) 18.09 $\downarrow 7.08$ 88.77 $\downarrow 49.09$ Gemma 3 27B 30.23 - 136.43 - LVLM-Count (Gemma 3 27B as LVLM) 19.70 $\downarrow 10.53$ 110.62 $\downarrow 25.81$ Qwen2 VL 72B AWQ 33.45 - 145.90 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 22.67 $\downarrow 10.78$ 124.73 $\downarrow 21.17$
+Table 14: Evaluation of SOTA counting models on the FSC-147 test set serves as a reference point for assessing the performance of LVLMs and their corresponding LVLM-Count variants.
+The column “Trained Model” indicates whether a model has been trained on FSC-147.
+Columns marked with $\Delta$ show the performance difference between LVLM-Count and its base LVLM.
+Improvements are shown in green, while degradations are shown in red.
+Results on the PASCAL VOC dataset are reported in Table 15.
+For training-based counting models, this is considered a cross-dataset evaluation, meaning that they are trained on the training set of FSC-147 and tested on the PASCAL VOC benchmark.
+We observe that some trained counting models still outperform LVLMs; however, this is likely due to the significant overlap between the categories in PASCAL VOC and FSC-147.
+Nevertheless, LVLM-Count is very effective in narrowing the performance gap.
+Additionally, the results on the Emoji-Count benchmark are reported in Table 16.
+This is also a cross-dataset evaluation.
+In contrast to PASCAL VOC, Emoji-Count consists of complex concepts and categories that do not overlap with FSC-147.
+As a result, LVLMs, which have far stronger generalization power for unseen data and complex concepts, outperform counting models by a large margin.
+Employing LVLM-Count widens this gap even further.
+Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TrainingFree [Shi et al., 2024] 12.03 - 18.18 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 12.39 - 22.81 - CountGD [Amini-Naieni et al., 2024] 2.81 - 7.01 - GroundingRec [Dai et al., 2024] 4.05 - 7.80 - GPT4o 4.46 - 8.35 - LVLM-Count (GPT4o as LVLM) 3.31 $\downarrow$ 1.15 7.11 $\downarrow$ 1.24 Gemma 3 27B 3.37 - 7.49 - LVLM-Count (Gemma 3 27B as LVLM) 2.87 $\downarrow$ 0.5 5.84 $\downarrow$ 1.65 Qwen2 VL 72B AWQ 4.83 - 8.84 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 4.05 $\downarrow$ 0.78 7.72 $\downarrow$ 1.12
+Table 15: Evaluation of SOTA counting models on the PASCAL VOC counting benchmark as a reference point for the performance of LVLMs and their corresponding LVLM-Count performance.
+Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC [Shi et al., 2024] 64.64 - 87.45 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 198.99 - 208.08 - CountGD [Amini-Naieni et al., 2024] 137.93 - 156.80 - GroundingREC [Dai et al., 2024] 143.22 - 158.74 - GPT-4o 22.51 - 35.94 - LVLM-Count (GPT-4o as LVLM) 16.29 $\downarrow 6.22$ 32.47 $\downarrow 3.47$ Gemma 3 27B 21.43 - 24.15 - LVLM-Count (Gemma3 27B as LVLM) 15.76 $\downarrow 5.67$ 21.00 $\downarrow 3.15$ Qwen2 VL 72B AWQ 82.41 - 186.32 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 23.74 $\downarrow 58.67$ 40.04 $\downarrow 146.28$
+Table 16: Evaluation of SOTA counting models on the Emoji-Count benchmark, serving as a reference point for the performance of LVLMs and their corresponding LVLM-Count performance.
+Finally, Table 17 presents the results of prior counting models on the Penguin benchmark alongside the counting performance of LVLMs.
+We observe that prior training-based counting models significantly outperform LVLMs on this benchmark.
+However, similar to other datasets, LVLM-Count narrows this performance gap.
+Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC [Shi et al., 2024] 59.34 - 74.40 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 22.29 - 31.38 - CountGD [Amini-Naieni et al., 2024] 17.07 - 25.08 - GroundingREC [Dai et al., 2024] 22.04 - 26.75 - GPT4o 36.01 - 46.64 - LVLM-Count (GPT-4o as LVLM) 26.95 $\downarrow$ 9.06 39.32 $\downarrow$ 7.32 Gemma 3 27B 49.87 - 61.06 - LVLM-Count (Gemma 3 27B as LVLM) 36.75 $\downarrow$ 13.12 47.24 $\downarrow$ 13.82 Qwen2 VL 72B AWQ 44.96 - 66.78 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 28.67 $\downarrow$ 16.29 43.71 $\downarrow$ 23.07
+Table 17: Evaluation of SOTA counting models on the Penguin benchmark, serving as a reference point for the performance of LVLMs and their corresponding LVLM-Count performance.
+Appendix L LVLM-Count’s Ability to Handle Visual Exemplars
+Exemplar-based counting methods face challenges due to the difficulty of acquiring representative exemplars.
+Additionally, they often struggle with intra-class variability, as objects within the same category may exhibit diverse appearances, leading to noisy matches and reduced accuracy.
+In contrast, text-based methods offer greater flexibility, as textual descriptions are easily provided, modifiable, and capable of encoding abstract or fine-grained concepts.
+This adaptability makes text-based approaches particularly suitable for dynamic or open-set scenarios, where predefined exemplars are impractical.
+Our focus in this work, similar to many recent works Liu et al.
+[2023], Amini-Naieni et al.
+[2023], is on text-based counting due to the above-mentioned reasons.
+Nevertheless, LVLM-Count is capable of working with visual exemplars alone as well as combinations of text prompts and visual exemplars, in addition to its text-only capabilities.
+The adaptation is straightforward, as most LVLMs also accept visual prompts.
+In Table 18, we evaluate the base GPT-4o and LVLM-Count (which uses GPT-4o) on Emoji-Count with text-only prompts, visual exemplars, and combinations of text and visual exemplars.
+For the visual-exemplar version, a single image of the target emoji icon is provided, and the LVLM is required to count similar instances.
+For the combined text-and-visual version, an image of the target emoji icon alongside its name is provided to the model.
+The results show higher accuracy when combining text and visual exemplars, while the text-only and exemplar-only versions achieve approximately the same performance.
+Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT4o 22.51 - 35.94 - LVLM-Count (GPT4o as LVLM, text) 16.29 $\downarrow$ 6.22 32.47 $\downarrow$ 3.47 LVLM-Count (GPT4o as LVLM, visual exemplars) 15.66 $\downarrow$ 6.85 31.69 $\downarrow$ 4.25 LVLM-Count (GPT4o as LVLM, text + visual exemplars) 13.46 $\downarrow$ 9.05 24.35 $\downarrow$ 11.59
+Table 18: Evaluation of LVLM-Count’s performance using text-only inputs, exemplar-only inputs, and a combination of text and exemplars on the Emoji-Count dataset.
+Appendix M Chain of Thought Prompting for LVLM-Count
+In this work, our primary focus is on developing a pipeline to enhance the counting ability of LVLMs for counting prompts in general, rather than optimizing for a specific prompting method.
+Nevertheless, we investigated the impact of Chain-of-Thought (CoT) prompting Wei et al.
+[2022] on visual counting task.
+For our experiments, we appended the phrase "think step-by-step" to the counting prompt.
+As shown in Table 19, the results on the PASCAL VOC benchmark reveal that CoT negatively affects counting performance—not only for LVLM-Count but also for the base LVLM.
+Method MAE $\downarrow$ RMSE $\downarrow$ GPT4o 4.46 8.35 GPT4o with CoT prompting 5.69 10.10 LVLM-Count (GPT4o as LVLM) 3.31 7.11 LVLM-Count with CoT prompting (GPT4o as LVLM) 3.89 7.70
+Table 19: Evaluation of CoT prompting for visual counting on PASCAL VOC.
+Appendix N Illustration of the Complete Workflow of the LVLM-Count for an Additional Image
+In this section, we demonstrate the same steps illustrated for the example image of eggs in Figures 3, 4, 5, and 6 for an image of zebras drinking water.
 The zebra image is passed to the pipeline along with the question $Q=$“how many zebras are in the image?”.
@@
 First, the masks are used in the cluster-based approach to find the location of the start and end points of the division paths, i.e., ($P^{1}_{s},P^{1}_{e}$) and ($P^{2}_{s},P^{2}_{e}$).
-Then these masks are turned into black-and-white images, which, in turn, are mapped to a graph.
+Then these masks are turned into a black-and-white image, which, in turn, is mapped to a graph.
 The division paths are then found by connecting each start point to its corresponding end point by running the $A^{*}$ search algorithm on the graph.
@@
 The masks from Figure 22(b) are turned into a black-and-white image.
-A dividing path is found by connecting $P_{s}$ to $P_{e}$ using the $A^{*}$ search algorithm in a graph that corresponds to the binary image, where the only nodes in the graph are white pixels, which are fully connected.
+A dividing path is found by connecting $P_{s}$ to $P_{e}$ using the $A^{*}$ search algorithm in a graph that corresponds to the binary image, where the only nodes in the graph are white pixels, which are connected to all of their white neighboring pixels.
 The path is mapped back to the pixel domain.
-Appendix J False Positive Masks at the Target Segmentation Stage
-One of the reasons we task an LVLM to count the objects in the subimages instead of using the number of generated masks at the target segmentation stage as the final count of the objects of interest is the existence of false positive masks.
-The GroundingDINO model is responsible for detecting the objects of interest, determined by expression $E$, and passing the output bounding boxes to SAM for producing segmentation masks.
-Nonetheless, GroundingDINO is not as strong as an LVLM in understanding expressions extracted from complex questions.
-Thus, it often returns bounding boxes for all instances of the object category mentioned in the expression, even if those instances do not satisfy other conditions in the expression.
-For example, in the top row of Figure 22, $E=$brown egg”.
-However, all the eggs have been segmented regardless of their color.
-Thus, counting the masks results in a significant error.
-Interestingly, as we can see, the false positive masks do not negatively affect LVLM-Count’s final answer, as the counting is done by an LVLM at the final stage, which is much stronger than GroundingDINO at understanding referring expressions.
-The only effect is that the white eggs have not been cut through by the division lines either.
-In the bottom row, we have chosen an image from the challenging Emoji-Count benchmark.
-The image contains icons, all of which have an arrow but point in different directions.
-However, the objects of interest are only “right arrows curving left.” Similar to the eggs example, taking the masks used for object-aware division results in a significant error.
-Figure 22: Top row: The object of interest is “brown egg.” However, all the eggs have been segmented because of the limitation of the GroundingDINO model in understanding complex referring expressions.
-Regardless, LVLM-Count provides a significantly more accurate number compared to the number of masks.
-Bottom row: The object of interest is “right arrows curving left.” Similar to the image of the eggs, counting the number of masks results in a very large error, while LVLM-Count provides a much more accurate number.
-Appendix K LVLM-Count’s Performance on a Benchmark Sampled from the PASCAL VOC Dataset
-The PASCAL VOC dataset [Everingham et al., 2015] is a well-known dataset in the field of computer vision, depicting everyday objects in everyday scenes.
-It is primarily used for tasks like object detection, classification, and segmentation.
-To provide more experimental results for the performance of LVLM-Count, we adapt a small subset of this dataset into the form of a counting benchmark.
-Similar to Chattopadhyay et al.
-[2017], we choose PASCAL VOC 2007 among other variants.
-This variant contains a training set of 2501 images, a validation set of 2510 images, and a test set of 4952 images, with 20 object categories that remain consistent across the splits.
-Each image includes annotations for instances of the 20 object categories in the dataset.
-The most frequent count per object category (as one would expect in everyday scenes) is 0.
-Moreover, there is a clear bias toward lower count values.
-To address this bias and obtain a balanced benchmark, we first create 20 simple counting questions asking for the number of objects from each of the 20 categories for every image in the test set.
-Then, we randomly sample five questions for each ground truth count.
-Note that the ground truth counts for this sampling are based on the original annotations of the dataset.
-The bias toward lower numbers is so significant that for ground truth counts larger than 16, there are fewer than five questions for the entire test set.
-In such cases, we simply use the available samples.
-This process resulted in 102 questions in total.
-Finally, we manually checked the ground truth counts and corrected them if required.
-We call the resulting counting questions the PASCAL VOC counting benchmark.
-We evaluate the performance of LVLM-Count on this benchmark using three different LVLMs: GPT-4o, Gemini 1.5 Pro, and Qwen2 VL 72B AWQ.
-Note that we use a clustering approach to automatically find the start and end points of the division paths along the $x$-axis.
-Table 7 shows the performance of LVLM-Count in comparison with state-of-the-art counting models and the base LVLMs.
-We observe that LVLM-Count improves upon the base LVLMs it uses and outperforms all prior counting methods, except CountGD.
-Upon further inspection, we noticed that the 20 object categories in PASCAL VOC have a high overlap with the object categories of the dataset used to train CountGD.
-Table 7: Evaluation of state-of-the-art models on the PASCAL VOC counting benchmark.
-The results for base LVLMs and LVLM-Count are reported over three trials.
-Columns marked with $\Delta$ show the improvement brought by our method over the base LVLM that it uses.
-Method EA (%$\uparrow$ $\Delta$ MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TrainingFree [Shi et al., 2024] 2.94 - 12.03 - 18.18 - GroundingRec [Dai et al., 2024] 19.61 - 5.05 - 8.44 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 5.88 - 12.39 - 22.81 - CountGD [Amini-Naieni et al., 2024] 26.47 - 2.81 - 7.01 - GPT4o 30.39 - 4.46 - 8.35 - LVLM-Count (GPT4o as LVLM) 31.37 $\uparrow$ 0.98 3.55 $\downarrow$ 0.91 7.18 $\downarrow$ 1.17 Gemini 1.5 Pro 33.01 - 3.24 - 6.62 - LVLM-Count (Gemini 1.5 Pro as LVLM) 39.22 $\uparrow$ 6.21 3.00 $\downarrow$ 0.24 6.30 $\downarrow$ 0.32 Qwen2 VL 72B AWQ 24.18 - 4.83 - 8.84 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 29.08 $\uparrow$ 4.90 4.12 $\downarrow$ 0.71 7.59 $\downarrow$ 1.25
-Appendix L Report of Various Accuracy Metrics for the Performance of LVLM-Count on the FSC-147 dataset, TallyQA Benchmark, and Emoji-Count Benchmark
-This section presents various accuracy measures for the experiments reported in Tables Table 1, 2, and 3.
-The accuracy metrics are defined in Table 8.
-The observations for each table can be summarized as follows:
-i) FSC-147: For this dataset, similar to the results in Table 1, LVLM-Count achieves higher accuracy metrics compared to the prior training-free method and each of the base LVLMs it uses.
-However, models specifically trained on this dataset generally achieve higher accuracy.
-ii) TallyQA Simple Benchmark: For this benchmark, LVLM-Count achieves higher accuracy compared to all prior counting models.
-However, it does not surpass the accuracy of the base LVLMs used.
-This is because the questions are straightforward, and the ground truth values are limited to numbers between 0 and 15—a range where the base LVLMs excel.
-This observation aligns with those in Figure 2 and Figure 18.
-iii) TallyQA Complex Benchmark: For this benchmark, LVLM-Count demonstrates significantly higher accuracies compared to prior counting models and, more importantly, consistent accuracy improvements over the base LVLMs used.
-iv) Emoji-Count: This is a challenging benchmark due to high object counts.
-LVLM-Count achieves substantially higher accuracies than both the base LVLMs and prior counting models, particularly for metrics like Acc±$\pm k$ where $k\geq 1$.
-Table 8: Definitions of Various Accuracy Metrics. GT denotes the ground truth number.
-Metric Definition
-Acc Percentage of answers such that $answer=GT$
-Acc±$\pm 1$ Percentage of answers such that $|answer-GT|\leq 1$
-Acc±$\pm 3$ Percentage of answers such that $|answer-GT|\leq 3$
-Acc±$\pm 5$ Percentage of answers such that $|answer-GT|\leq 5$
-Acc±$\pm 10$ Percentage of answers such that $|answer-GT|\leq 10$
-Table 9: FSC-147 Dataset.
-A ($\uparrow$) next to the measured accuracies for LVLM-Count indicates improvement over the base LVLM it uses, while a ($\downarrow$) indicates degradation compared to the corresponding base LVLM.
-Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) TFOC [Shi et al., 2024] 9.33 20.17 34.20 44.37 61.60 GroundingRec [Dai et al., 2024] 34.03 51.68 67.65 75.04 85.13 CountGD [Amini-Naieni et al., 2024] 31.85 47.90 63.78 73.03 82.61 GPT4o 12.24 26.22 42.10 52.41 66.58 LVLM-Count (GPT4o as LVLM) 14.26 ($\uparrow$) 28.10 ($\uparrow$) 47.42 ($\uparrow$) 58.01 ($\uparrow$) 72.91 ($\uparrow$) Gemini 1.5 Pro 12.97 26.58 41.57 51.71 63.78 LVLM-Count (Gemini 1.5 Pro as LVLM) 13.92 ($\uparrow$) 27.96 ($\uparrow$) 47.65 ($\uparrow$) 58.85 ($\uparrow$) 75.07 ($\uparrow$) Qwen2 VL 72B AWQ 9.19 20.81 36.83 46.95 62.07 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 9.80 ($\uparrow$) 23.14 ($\uparrow$) 40.95 ($\uparrow$) 51.04 ($\uparrow$) 68.54 ($\uparrow$)
-Table 10: TallyQA Simple Benchmark.
-A ($\uparrow$) next to the measured accuracies for LVLM-Count indicates improvement over the base LVLM it uses, while a ($\downarrow$) indicates degradation compared to the corresponding base LVLM.
-Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) TFOC [Shi et al., 2024] 6.45 17.42 34.84 59.35 79.35 GroundingRec [Dai et al., 2024] 23.87 49.03 73.55 86.45 93.55 CountGD [Amini-Naieni et al., 2024] 41.94 63.23 78.06 85.16 94.84 GPT4o 44.30 69.89 87.96 94.41 100.00 LVLM-Count (GPT4o as LVLM) 44.73 ($\uparrow$) 73.33 ($\uparrow$) 91.83 ($\uparrow$) 96.99 ($\uparrow$) 99.57 ($\downarrow$) Gemini 1.5 Pro 47.10 75.27 92.90 97.63 100.00 LVLM-Count (Gemini 1.5 Pro as LVLM) 45.16 ($\downarrow$) 76.13 ($\uparrow$) 91.61 ($\downarrow$) 95.91 ($\downarrow$) 99.35 ($\downarrow$) Qwen2 VL 72B AWQ 49.03 70.75 87.74 93.33 98.92 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 41.72 ($\downarrow$) 67.10 ($\downarrow$) 85.81 ($\downarrow$) 93.55 ($\uparrow$) 98.49 ($\downarrow$)
-Table 11: TallyQA Complex Benchmark.
-A ($\uparrow$) next to the measured accuracies for LVLM-Count indicates improvement over the base LVLM it uses, while a ($\downarrow$) indicates degradation compared to the corresponding base LVLM.
-Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) TFOC [Shi et al., 2024] 1.34 15.44 32.89 46.31 66.44 GroundingRec [Dai et al., 2024] 16.78 28.19 51.68 64.43 85.23 CountGD [Amini-Naieni et al., 2024] 5.37 12.75 31.54 48.99 73.15 GPT4o 29.08 50.34 74.94 86.13 98.21 LVLM-Count (GPT4o as LVLM) 34.68 ($\uparrow$) 55.03 ($\uparrow$) 78.97 ($\uparrow$) 89.71 ($\uparrow$) 96.42 ($\uparrow$) Gemini 1.5 Pro 25.73 25.73 81.66 91.72 98.88 LVLM-Count (Gemini 1.5 Pro as LVLM) 26.62 ($\uparrow$) 51.68 ($\uparrow$) 76.06 ($\downarrow$) 85.23 ($\downarrow$) 94.41 ($\downarrow$) Qwen2 VL 72B AWQ 24.61 46.31 68.01 78.75 94.41 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 30.65 ($\uparrow$) 55.26 ($\uparrow$) 78.30 ($\uparrow$) 86.80 ($\uparrow$) 96.64 ($\uparrow$)
-Table 12: Emoji Benchmark.
-A ($\uparrow$) next to the measured accuracies for LVLM-Count indicates improvement over the base LVLM it uses, while a ($\downarrow$) indicates degradation compared to the corresponding base LVLM.
-Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) TFOC [Shi et al., 2024] 0.24 0.72 2.89 4.58 8.43 GroundingRec [Dai et al., 2024] 4.58 8.43 12.29 17.11 24.10 CountGD [Amini-Naieni et al., 2024] 0.48 0.72 0.96 0.96 1.20 GPT4o 1.85 5.54 13.73 21.12 43.94 LVLM-Count (GPT4o as LVLM) 4.98 ($\uparrow$) 17.03 ($\uparrow$) 37.67 ($\uparrow$) 55.90 ($\uparrow$) 76.71 ($\uparrow$) Gemini 1.5 Pro 2.65 7.23 14.14 23.37 42.97 LVLM-Count (Gemini 1.5 Pro as LVLM) 2.25 ($\downarrow$) 5.46 ($\downarrow$) 13.57 ($\downarrow$) 23.61 ($\uparrow$) 57.19 ($\uparrow$) Qwen2 VL 72B AWQ 0.88 2.89 7.55 11.41 19.76 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 3.37 ($\uparrow$) 8.76 ($\uparrow$) 19.04 ($\uparrow$) 29.00 ($\uparrow$) 48.51 ($\uparrow$)
-Appendix M Visual Examples of LVLM-Count’s Performance on the FSC-147 Dataset, TallyQA Benchmark, and Emoji-Count Benchmark
-This section presents several visual examples showcasing the performance of LVLM-Count on the FSC-147 dataset, the TallyQA benchmark, and the Emoji-Count benchmark.
-The LVLM used in the pipeline to generate these visual examples is GPT-4o.
-Figure 23 illustrates both successful and unsuccessful examples of LVLM-Count’s performance on FSC-147.
-The top two rows in this figure demonstrate strong performance, where the sub-images preserve sufficient context, enabling the LVLM to predict the correct count.
-The last row depicts a failure case.
-Here, the task is to count the number of skis.
-However, LVLM-Count generates sub-images containing only irrelevant objects.
-Consequently, GPT-4o is misled by these objects and predicts an incorrect non-zero count.
-Figure 23: Illustration of good and bad examples for LVLM-Count.
-These examples are taken from FSC-147.
-The first column shows the input image.
-The second column shows the paths found in the object-aware division stage.
-The third column illustrates the obtained sub-images and the predicted count by LVLM-Count.
-The top two rows correspond to good examples where the base GPT-4o has a large error, but LVLM-Count provides an answer very close to the ground truth.
-The bottom row shows a bad example where the base GPT-4o performs better, and LVLM-Count has a large error.
-Additionally, Figure 24 displays an example from the TallyQA Simple benchmark, while Figure 25 illustrates an example from the TallyQA Complex benchmark.
-Additionally, we include five visual examples from the Emoji-Count benchmark, split into two figures for better clarity and presentation.
-Figure 26 contains the first three examples, and Figure 27 includes the final two examples.
-In these visual examples, LVLM-Count consistently achieves more accurate results compared to the base GPT-4o.
-Figure 24: A visual example of the performance of LVLM-Count on the TallyQA Simple benchmark. The input question is: “How many people are in the image?”.
-Figure 25: A visual example of the performance of LVLM-Count on the TallyQA Complex benchmark. The input question is: “How many pieces of the plane are yellow?”.
-Figure 26: Three examples of the performance of LVLM-Count on the Emoji-Count benchmark.
-Top row: The object of interest is “twelve o’clock”.
-Middle row: The object of interest is “kissing cat”.
-Bottom row: The object of interest is “woman with red hair”.
-Figure 27: Two additional examples of the performance of LVLM-Count on the Emoji-Count benchmark.
-Top row: The object of interest is “derelict house”.
-Bottom row: The object of interest is “right arrow curving left”.
+Appendix O Bias to Small Numbers in Datasets and Counting Performance of LVLMs
+In our experiments, LVLMs are able to make correct predictions when the number of items to be counted is small, but errors increase as the ground truth number grows.
+Although, we cannot be certain, one likely reason for this behavior in LVLMs is that, during training, the counting questions these models encounter are heavily biased toward small numbers.
+As an example, in Figure 24 we show the distribution of ’How many’ questions in some well-known VQA datasets.
+(a) TallyQA Dataset [Acharya et al., 2019]
+(b) VQAv2 Dataset [Goyal et al., 2017]
+Figure 24: Distribution of answers in two VQA datasets.

### v4

@@
 Despite their advanced visual perception, large vision-language models (LVLMs) are known to struggle with counting tasks.
-In this work, we evaluate the performance of several recent LVLMs on visual counting tasks across multiple counting and vision datasets.
+In this work, we evaluate the performance of several LVLMs on visual counting tasks across multiple counting and vision datasets.
 We observe that while their performance may be less prone to error for small numbers of objects, they exhibit significant weaknesses as the number of objects increases.
@@
 Finally, the number of objects of interest in each sub-image is computed using an LVLM, and the results are aggregated.
-Recent advancements in the field of large vision-language models (LVLMs), such as GPT-4o (Achiam et al., 2023), coupled with the massive scale of web-scraped data used to train them, have resulted in unprecedented zero-shot capabilities across diverse tasks.
-However, evaluations of LVLMs’ capabilities have revealed weaknesses in numerical reasoning Yin et al.
-(2023); Xu et al.
-(2024); Yang et al.
-(2023).
-In this work, we focus specifically on the visual counting task, one of the most fundamental aspects of numerical reasoning.
+These applications often require distinguishing between objects of the same class with subtle variations, as well as recognizing complex concepts.
+Models trained solely on counting datasets struggle to generalize to such scenarios due to the limited availability of annotated data for fine-grained counting.
+Recent advancements in large vision-language models (LVLMs), such as GPT-4o (Achiam et al., 2023), combined with the massive scale of web-scraped training data, have enabled unprecedented zero-shot recognition capabilities, making them a promising candidate for handling complex and fine-grained counting tasks.
+However, evaluations of LVLMs have also revealed notable weaknesses in their numerical reasoning (Yin et al., 2023; Xu et al., 2024; Yang et al., 2023).
+In this work, we focus specifically on the visual counting, one of the most fundamental aspects of numerical reasoning.
 We observe that although LVLMs perform well in counting small numbers of objects—typically fewer than 20—their accuracy deteriorates with larger quantities.
@@
 (2017) we propose a simple yet effective baseline method to alleviate this issue.
-Leveraging a divide-and-conquer approach, we engineer a simple pipeline that divides an image into sub-images, and prompts the LVLM to count the objects of interest in each sub-image.
+Leveraging a divide-and-conquer approach, we engineer a simple pipeline that divides an image into carefully cut sub-images, and prompts the LVLM to count the objects of interest in each sub-image.
 The counts from the sub-images are then aggregated to make the final prediction.
@@
 To have this mechanism two existing pre-trained detection and segmentation models are employed in the pipeline.
-The workflow is illustrated in Figure 1.
+The workflow is illustrated in Figure˜1.
 Initially, in our pipeline, the category name of the object of interest is extracted from the input question using an LLM (We use the LVLM as an LLM for this step).
@@
     We propose a solution for object-aware division.
-Accurate division is crucial, as parts of cut objects can lead to over-counting (see Figure 2(a)).
+Accurate division is crucial, as parts of cut objects can lead to over-counting (see Figure˜2(a)).
 The proposed solution divides images without cutting through objects of interest specified by an arbitrary prompt.
@@
 The objects of interest are the circles.
-In Figure 2(a), we illustrate a naive division of the input image, which is divided into equally sized sub-images with straight lines.
-In Figure 2(b), we illustrate the object-aware division, which avoids cutting through circles.
-In Figure 2(c), we illustrate the counting error of GPT-4o for images with randomly positioned circles.
+In Figure˜2(a), we illustrate a naive division of the input image, which is divided into equally sized sub-images with straight lines.
+In Figure˜2(b), we illustrate the object-aware division, which avoids cutting through circles.
+In Figure˜2(c), we illustrate the counting error of GPT-4o for images with randomly positioned circles.
 The absolute counting error is the absolute difference between the ground truth and the number predicted by GPT-4o.
@@
 Instead, they cast the counting problem as a prompt-based segmentation task, using SAM (Kirillov et al., 2023) to obtain segmentation masks that determine the output number.
+Despite this progress, these models remain constrained by two principal limitations.
+First, their performance can degrade significantly under considerable distribution shift in the input samples.
+Second, the prompts they interpret are typically limited to simple object categories or referring expressions; as task complexity increases, these models often fail.
+Employing LVLMs for counting offers a potential pathway to mitigate these issues.
 Leveraging the concept of divide and conquer for counting.
@@
 We consider the non-overlapping areas as “detected areas”, which are then cropped to be passed to the next stage.
-This process is illustrated in Figure 3.
+This process is illustrated in Figure˜3.
 Figure 3: Illustration of the area detection stage of LVLM-Count.
@@
 The bounding boxes are then given as input to a segmentation model, which returns segmentation masks for the objects within each bounding box.
-We illustrate an example of this process in Figure 4.
+We illustrate an example of this process in Figure˜4.
 (a) GroundingDINO output
@@
 The goal is to produce all the instance masks for $E$ set to “brown egg”.
-The cropped detected area from Figure 3, together with $E$, is given as input to GroundingDINO, which produces the output shown in Figure 4(a).
-Figure 4(a) is then given as input to SAM, which produces the output shown in Figure 4(b).
+The cropped detected area from Figure˜3, together with $E$, is given as input to GroundingDINO, which produces the output shown in Figure˜4(a).
+Figure˜4(a) is then given as input to SAM, which produces the output shown in Figure˜4(b).
 How to determine the bounding boxes.
@@
 Additionally, we apply an erosion function to the segmentations, ensuring adjacent masks maintain a minimum separation of two pixels.
-For a detailed analysis and visual examples of how these post-processing operations enhance robustness in crowded scenes and occlusion cases, see Appendix C.
+For a detailed analysis and visual examples of how these post-processing operations enhance robustness in crowded scenes and occlusion cases, see Appendix˜C.
 The final processed masks for each cropped area are then passed to the next stage of our pipeline.
@@
 Furthermore, as we will demonstrate in subsequent sections, all masks are removed from the sub-images after division, ensuring no noise propagates from this stage to the LVLM-based counting stage.
-We demonstrate the effectiveness and robustness of our strategy in the early pipeline stages by showing that LVLM-Count significantly improves LVLMs’ performance on one of the most challenging counting datasets—the Penguin Dataset Penguin Research (2016), as detailed in Section 4.2.
+We demonstrate the effectiveness and robustness of our strategy in the early pipeline stages by showing that LVLM-Count significantly improves LVLMs’ performance on one of the most challenging counting datasets—the Penguin Dataset Penguin Research (2016), as detailed in Section˜4.2.
 This dataset is particularly difficult due to frequent occlusion, camouflage, and complex backgrounds Arteta et al.
@@
 Instead, SAM is configured to generate segmentation masks for all objects in the scene.
-As shown in Table 5, our method still enhances counting performance in this configuration, underscoring its robustness and minimal dependence on the detection model during the initial pipeline stages.
+As shown in Table˜5, our method still enhances counting performance in this configuration, underscoring its robustness and minimal dependence on the detection model during the initial pipeline stages.
 3.3 Object-aware Division
@@
 To determine the location of the division paths on the $x$-axis, the samples taken from the masks are projected onto the $x$-axis.
-The projected points are automatically clustered using a non-parametric mean-shift algorithm.
+The projected points are automatically clustered using a non-parametric mean-shift algorithm [^1] In our implementation, we employ the MeanShift algorithm from the Scikit-learn library (Pedregosa et al., 2011), utilizing its default heuristic for automatic bandwidth estimation.
+(Comaniciu and Meer, 2002).
 Once the clusters are identified, the point between the point with the highest $x$ value in one cluster and the point with the lowest $x$ value in the subsequent cluster is considered the $x$-coordinate of a division path.
@@
 For example, if there is only one cluster along the $x$-axis, no division is required, and if there are two clusters, one vertical path will divide the image into two parts.
-We illustrate this approach in Figure 5.
+We illustrate this approach in Figure˜5.
 Figure 5: Illustration of the unsupervised and non-parametric method to obtain the division points ($P^{1}_{s},P^{1}_{e}$), and ($P^{2}_{s},P^{2}_{e}$,).
@@
 We use the $A^{*}(G,P_{s},P_{e},g)$ search algorithm to find a path that connects $P_{s}$ to $P_{e}$, where the heuristic $g$ is set to be Manhattan distance.
-The output of $A^{*}$ is a set of connected pixels that go around the obstacles and connect $P_{s}$ to $P_{e}$, creating an object-aware division path, as shown in Figure 6.
+The output of $A^{*}$ is a set of connected pixels that go around the obstacles and connect $P_{s}$ to $P_{e}$, creating an object-aware division path, as shown in Figure˜6.
 The path-finding algorithm is run for all division coordinates.
 Finally, we draw the image contours based on these division paths and take the area surrounded by each contour as a resulting sub-image.
+Note that although not part of the main LVLM-Count pipeline, the above two steps can be applied to the $y$-axis in the same manner to obtain horizontal division paths.
+For further discussion, the reader may refer to Appendix˜S.
 Figure 6: Illustration of object-aware division.
@@
 Additionally, we show the success of our design in enhancing counting performance on one of the most challenging counting datasets that features heavy occlusions and complex backgrounds.
-The code to reproduce the experiments can be found here: https://github.com/mrghofrani/lvlm-count.
+The code to reproduce the experiments is available at our GitHub repository  [^2] Our GitHub repository is available at: https://github.com/MuhammadFetrat/lvlm-count.
 4.1 Datasets and Benchmarks
@@
 This results in 415 image-question pairs.
-We illustrate two examples of this dataset in Figure 7.
+We illustrate two examples of this dataset in Figure˜7.
 (a) Q: How many waning gibbous moons are there in the image? Answer: 18.
 (b) Q: How many clocks at time “two-thirty” are there in the image? Answer: 15.
-Figure 7: Illustration of a smaller version of two challenging cases from Emoji-Count. In Figure 7(a), the class name is “Moon Phase”. In Figure 7(b), the class name is “Clock Time”.
+Figure 7: Illustration of a smaller version of two challenging cases from Emoji-Count. In Figure˜7(a), the class name is “Moon Phase”. In Figure˜7(b), the class name is “Clock Time”.
 Penguin Benchmark.
@@
 To build a manageable benchmark, we randomly sample 100 images while preserving a balanced ground truth range.
-We refer the reader to Appendix D for more details about this benchmark.
+We refer the reader to Appendix˜D for more details about this benchmark.
 4.2 Results
-The following section presents the numerical results of our experiments with the base LVLMs and their corresponding LVLM-Count on each benchmark described in Section 4.1.
-For visual examples of LVLM-Count’s performance on these benchmarks, see Appendix B.
-The ablation study results are discussed in Appendix A.
-Additionally, Appendix K provides the performance of SOTA counting models on these benchmarks as a reference for evaluating the LVLMs’ performance.
-Also, an inference time analysis is provided in Appendix J.
+This section presents the numerical results of our experiments with the base LVLMs and their corresponding LVLM-Count on each benchmark described in Section˜4.1.
+In the experiments, we consider two baselines: (i) a baseline where the number of target segmentation masks is taken as the final answer, and (ii) a baseline where an image of the generated segmentation masks is provided to GPT-4o along with a prompt to count them.
+For reference, we also report the performance of three state-of-the-art (SOTA) text-based counting-specific models: GroundingREC (Dai et al., 2024), CountGD (Amini-Naieni et al., 2024), and $\text{DAVE}_{\text{prm}}$ (Pelhan et al., 2024).
+It is important to note that the primary aim of this work is to propose a method that enhances the counting ability of an arbitrary large vision-language model (LVLM) across diverse datasets.
+The inclusion of these specialized models in our comparison should not be interpreted as an attempt to replace or compete with them on specific counting tasks—particularly those involving simple counting.
+These models are trained on dedicated counting datasets and perform efficiently on test samples from the same distribution or datasets with high category overlap.
+However, their performance may degrade significantly when applied to out-of-distribution samples.
+Furthermore, these models are constrained by the type and complexity of the input prompts they can process; they typically only accept simple category names or referring expressions, and often fail when presented with complex phrases or sentences.
+In contrast, our method is designed to be comprehensive.
+Although it may not match the efficiency of task-specific models on datasets they were trained on, it performs robustly across a wide range of test distributions, despite not being explicitly trained on any of the tasks.
+Moreover, it is capable of handling arbitrary prompt types and structures, regardless of complexity.
+We also report the performance of TFOC (Shi et al., 2024), a SOTA training-free model.
+In essence, TFOC is a segmentation-based counting approach, similar to our baselines, but employs a more sophisticated methodology.
+The above discussion is reflected in our experimental results.
+For the trained counting models, we report performance using weights obtained from training on the FSC-147 training set.
+These models perform well on the FSC-147 dataset and on the PASCAL VOC benchmark, which has high category overlap with FSC-147.
+However, on the Emoji-Count dataset—which exhibits a different distribution and requires complex understanding—their performance drops considerably.
+In contrast, our method, LVLM-Count, maintains high performance even on challenging benchmarks such as Emoji-Count.
+Another experiment on the TallyQA-Complex benchmark, which contains real-world scenes with complex counting questions, further reveals the limited complex reasoning capability of conventional trained models.
+LVLMs and our proposed LVLM-Count, however, demonstrate a superior ability to handle such challenging cases.
+The corresponding results on the TallyQA-Complex benchmark are provided in Appendix˜T.
+We also conduct an ablation study on the effectiveness of different components of our pipeline in Appendix˜A.
+In summary, the ablation experiments support three conclusions: (i) each component has a positive effect over the baseline LVLM, and when combined, the full pipeline in Figure˜1 achieves the best performance compared to different variants; (ii) naive division with straight lines cannot replace object-aware division; and (iii) LVLM-Count is robust to the accuracy of the target segmentation stage.
+For visual examples of LVLM-Count’s performance on the benchmarks introduced in Section˜4.1, see Appendix˜B.
+An inference time analysis is provided in Appendix˜J.
+This analysis shows that the largest portion of the time is spent querying the LVLM, which is a step common to both the base LVLM and LVLM-Count.
+The other steps contribute a much smaller share to the overall inference time.
 FSC-147.
@@
 In all experiments, the detection thresholds are set to $0.1$.
-The results are presented in Table 1, reporting the mean absolute error (MAE) and root mean square error (RMSE).
+The results are presented in Table˜1, reporting the mean absolute error (MAE) and root mean square error (RMSE).
 Our findings indicate that LVLM-Count improves the performance of all three LVLMs in terms of MAE.
 Interestingly, while the base Qwen2 and Gemma 3 models are much less powerful than their commercial counterpart, they outperform the base GPT-4o when integrated into our pipeline.
-Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ Number of target segmentaion masks 44.14 - 154.39 - GPT-4o counting target segmentaion masks 38.45 - 156.11 - GPT-4o 25.17 - 137.86 - LVLM-Count (GPT-4o as LVLM) 18.09 $\downarrow 7.08$ 88.77 $\downarrow 49.09$ Gemma 3 27B 30.23 - 136.43 - LVLM-Count (Gemma 3 27B as LVLM) 19.70 $\downarrow 10.53$ 110.62 $\downarrow 25.81$ Qwen2 VL 72B AWQ 33.45 - 145.90 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 22.67 $\downarrow 10.78$ 124.73 $\downarrow 21.17$
 Table 1: Evaluation on the test set of the FSC-147 dataset.
-In all tables, the results for base LVLMs and LVLM-Count are reported over three trials.
-Also, columns marked with $\Delta$ show the performance difference between the LVLM-Count and the base LVLM it uses.
+In all the tables of this section, the results for base LVLMs and LVLM-Count are reported over six trials.
+In the tables, columns marked with $\Delta$ show the performance difference between the LVLM-Count and the base LVLM it uses.
 Green indicates improvement, while red represents degradation.
-To see the measured accuracy metrics for this dataset and the subsequent benchmarks, refer to Appendix I.
-Additionally, MAE analysis across different intervals of ground truth values for FSC-147 is provided in Appendix H.
+Also, the 95% confidence intervals for the MAEs are reported in the column named MAE 95% CI.
+To see the measured accuracy metrics for this dataset and the subsequent benchmarks, refer to Appendix˜I.
+Additionally, MAE analysis across different intervals of ground truth values for FSC-147 is provided in Appendix˜H.
+Method Trained Model MAE $\downarrow$ $\Delta$ MAE 95% CI RMSE $\downarrow$ $\Delta$ TFOC (Shi et al., 2024) 24.79 - - 137.15 - $\text{DAVE}_{\text{prm}}$ (Pelhan et al., 2024) 14.90 - - 103.42 - CountGD (Amini-Naieni et al., 2024) 14.76 - - 120.42 - GroundingREC (Dai et al., 2024) 10.12 - - 107.19 - Number of target segmentaion masks 44.14 - - 154.39 - GPT-4o counting target segmentaion masks 38.45 - - 156.11 - GPT-4o 25.57 - [24.74, 26.39] 137.26 - LVLM-Count (GPT-4o as LVLM) 17.86 $\downarrow 7.71$ [16.96, 18.77] 91.71 $\downarrow 45.55$ Gemma 3 27B 30.59 - [30.22, 30.97] 132.61 - LVLM-Count (Gemma 3 27B as LVLM) 20.25 $\downarrow 10.34$ [19.62, 20.89] 110.45 $\downarrow 22.16$ Qwen2 VL 72B AWQ 34.18 - [32.85, 35.51] 149.49 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 22.29 $\downarrow 11.89$ [21.63, 22.96] 119.46 $\downarrow 30.44$
 PASCAL VOC Benchmark.
 We evaluate the performance of three LVLMs—GPT-4o, Qwen2, and Gemma 3—on this benchmark, both with and without LVLM-Count.
-As shown in Table 2, LVLM-Count consistently outperforms the base LVLMs.
-Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ Number of the target segmentation masks 4.03 - 7.47 GPT-4o counting target segmentaion masks 6.60 - 15.09 GPT4o 4.46 - 8.35 - LVLM-Count (GPT4o as LVLM) 3.31 $\downarrow$ 1.15 7.11 $\downarrow$ 1.24 Gemma 3 27B 3.37 - 7.49 - LVLM-Count (Gemma 3 27B as LVLM) 2.87 $\downarrow$ 0.5 5.84 $\downarrow$ 1.65 Qwen2 VL 72B AWQ 4.83 - 8.84 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 4.05 $\downarrow$ 0.78 7.72 $\downarrow$ 1.12
+As shown in Table˜2, LVLM-Count consistently outperforms the base LVLMs.
 Table 2: Evaluation on the PASCAL VOC counting benchmark.
+Method MAE $\downarrow$ $\Delta$ MAE 95% CI RMSE $\downarrow$ $\Delta$ TFOC (Shi et al., 2024) 12.03 - - 18.18 - $\text{DAVE}_{\text{prm}}$ (Pelhan et al., 2024) 12.39 - - 22.81 - CountGD (Amini-Naieni et al., 2024) 2.81 - - 7.01 - GroundingRec (Dai et al., 2024) 4.05 - - 7.80 - Number of the target segmentation masks 4.03 - - 7.47 - GPT-4o counting target segmentaion masks 6.60 - - 15.09 - GPT4o 4.64 - [4.45, 4.82] 8.56 - LVLM-Count (GPT4o as LVLM) 3.42 $\downarrow$ 1.22 [3.25, 3.59] 7.17 $\downarrow$ 1.39 Gemma 3 27B 3.40 - [3.24 ,3.55] 7.54 - LVLM-Count (Gemma 3 27B as LVLM) 2.97 $\downarrow$ 0.43 [2.82, 3.11] 6.16 $\downarrow$ 1.38 Qwen2 VL 72B AWQ 4.80 - [4.61, 4.99] 8.71 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 4.12 $\downarrow$ 0.68 [3.87, 4.37] 7.76 $\downarrow$ 0.95
 Emoji-Count.
 We evaluate the performance of the LVLMs and their performance using LVLM-Count on the Emoji-Count benchmark.
-The results are shown in Table 3.
+The results are shown in Table˜3.
 This is a challenging benchmark, as it requires understanding complex concepts.
@@
 Nonetheless, the performance of all three base LVLMs is significantly enhanced by LVLM-Count, especially Qwen2, which performs almost on par with GPT-4o when LVLM-Count is used for it.
-Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ Number of the target segmentation masks 82.47 - 107.98 - GPT-4o counting target segmentaion masks 107.72 - 162.12 - GPT-4o 22.51 - 35.94 - LVLM-Count (GPT-4o as LVLM) 16.29 $\downarrow 6.22$ 32.47 $\downarrow 3.47$ Gemma 3 27B 21.43 - 24.15 - LVLM-Count (Gemma3 27B as LVLM) 15.76 $\downarrow 5.67$ 21.00 $\downarrow 3.15$ Qwen2 VL 72B AWQ 82.41 - 186.32 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 23.74 $\downarrow 58.67$ 40.04 $\downarrow 143.84$
 Table 3: Evaluation on the Emoji-Count benchmark.
+Method MAE $\downarrow$ $\Delta$ MAE 95% CI RMSE $\downarrow$ $\Delta$ TFOC (Shi et al., 2024) 64.64 - - 87.45 - $\text{DAVE}_{\text{prm}}$ (Pelhan et al., 2024) 198.99 - - 208.08 - CountGD (Amini-Naieni et al., 2024) 137.93 - - 156.80 - GroundingREC (Dai et al., 2024) 143.22 - - 158.74 - Number of the target segmentation masks 82.47 - - 107.98 - GPT-4o counting target segmentaion masks 107.72 - - 162.12 - GPT-4o 23.57 - [22.37, 24.76] 36.97 - LVLM-Count (GPT-4o as LVLM) 16.57 $\downarrow 7$ [16.25, 16.90] 33.11 $\downarrow 3.86$ Gemma 3 27B 21.39 - [21.32, 21.45] 24.04 - LVLM-Count (Gemma3 27B as LVLM) 16.16 $\downarrow 5.23$ [15.76, 16.56] 21.27 $\downarrow 2.77$ Qwen2 VL 72B AWQ 78.05 - [72.17, 83.93] 159.22 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 24.43 $\downarrow 53.62$ [23.65, 25.21] 43.38 $\downarrow 115.84$
 Penguin Benchmark.
-We report the counting performance on the Penguin benchmark in Table 4.
+We report the counting performance on the Penguin benchmark in Table˜4.
 We evaluate two variants of LVLM-Count.
@@
 Both variants improve MAE across all three LVLMs, highlighting LVLM-Count’s robustness in scenarios with heavy occlusion and complex backgrounds—conditions that pose significant challenges for area detection and target segmentation.
-Additional details and a visual example showing the segmentation masks and division paths generated in both variants for a sample from the Penguin benchmark can be found in Appendix D.
-Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT4o 36.01 - 46.64 - LVLM-Count( Main variant, using GPT-4o ) 26.95 $\downarrow$ 9.06 39.32 $\downarrow$ 7.32 LVLM-Count(SAM-only variant, using GPT-4o) 28.79 $\downarrow$ 7.22 45.71 $\downarrow$ 0.93 Gemma 3 27B 49.87 - 61.06 - LVLM-Count( Main variant, using Gemma 3 ) 36.75 $\downarrow$ 13.12 47.24 $\downarrow$ 13.82 LVLM-Count( SAM-only variant, using Gemma 3 ) 41.97 $\downarrow$ 7.90 54.55 $\downarrow$ 6.51 Qwen2 VL 72B AWQ 44.96 - 66.78 - LVLM-Count( Main variant, using Qwen2 ) 28.67 $\downarrow$ 16.29 43.71 $\downarrow$ 23.07 LVLM-Count(SAM-only, using Qwen2 ) 38.29 $\downarrow$ 6.67 51.47 $\downarrow$ 15.31
+Additional details and a visual example showing the segmentation masks and division paths generated in both variants for a sample from the Penguin benchmark can be found in Appendix˜D.
 Table 4: Evaluation on the Penguin benchmark.
-Limitations
+Method MAE $\downarrow$ $\Delta$ MAE 95% CI RMSE $\downarrow$ $\Delta$ GPT4o 35.18 - [33.13, 37.23] 45.76 - LVLM-Count( Main variant, using GPT-4o ) 26.76 $\downarrow$ 8.42 [25.86, 27.66] 38.60 $\downarrow$ 7.16 LVLM-Count(SAM-only variant, using GPT-4o) 29.02 $\downarrow$ 6.16 [28.21, 29.84] 44.89 $\downarrow$ 0.87 Gemma 3 27B 49.44 - [48.94, 49.95] 60.20 - LVLM-Count( Main variant, using Gemma 3 ) 34.13 $\downarrow$ 15.31 [31.72, 36.53] 44.11 $\downarrow$ 16.09 LVLM-Count( SAM-only variant, using Gemma 3 ) 41.09 $\downarrow$ 8.35 [39.92, 42.25] 53.01 $\downarrow$ 7.19 Qwen2 VL 72B AWQ 44.02 - [40.45, 47.60] 65.59 - LVLM-Count( Main variant, using Qwen2 ) 28.21 $\downarrow$ 15.81 [26.48, 29.94] 40.40 $\downarrow$ 25.19 LVLM-Count(SAM-only, using Qwen2 ) 33.33 $\downarrow$ 10.69 [30.45, 36.21] 47.55 $\downarrow$ 18.04
+5 Limitations
 In this work, we quantitatively evaluated the visual counting performance of several LVLMs on multiple datasets.
@@
 For instance, an instruction like "Place the final predicted number inside [[]]" enables the use of regex searches to extract the number for aggregation.
+Acknowledgments
+K.
+Fountoulakis would like to acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC).
+Cette recherche a été financée par le Conseil de recherches en sciences naturelles et en génie du Canada (CRSNG), [RGPIN-2019-04067, DGECR-2019-00147].
 Appendix Table of Contents
-[…]
-[sections]
-[…]
-[sections]l1
 Appendix A Ablation Study
@@
 The results demonstrate that unlike object-aware division, the naive approach is unreliable due to repetitive counting of fragmented objects.
-We perform ablation studies using GPT-4o, conducting experiments on the FSC-147 test set. Table 5 presents the results of these ablation experiments.
-Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT-4o 25.17 - 137.86 - GPT-4o + Naive division 33.04 $\uparrow$ 7.87 116.83 $\downarrow$ 21.03 GPT-4o + Area Detection + Naive division 32.69 $\uparrow$ 7.52 102.41 $\downarrow$ 35.45 GPT-4o + Area detection 23.08 $\downarrow$ 2.09 120.06 $\downarrow$ 17.8 GPT-4o + Object-aware division (using SAM without GroundingDINO) 21.01 $\downarrow$ 4.16 135.04 $\downarrow$ 2.82 GPT-4o + Object-aware division 19.17 $\downarrow$ 6.00 120.61 $\downarrow$ 17.25 GPT-4o + Area detection + Object-aware division, (equiv.
-to LVLM-Count) 18.09 $\downarrow$ 7.08 88.77 $\downarrow$ 49.09
+We perform ablation studies using GPT-4o, conducting experiments on the FSC-147 test set. Table˜5 presents the results of these ablation experiments.
 Table 5: An ablation study of LVLM-Count on the FSC-147 test dataset.
@@
 Green indicates improvement, while red represents degradation.
+Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT-4o 25.57 - 137.26 - GPT-4o + Naive division 33.04 $\uparrow$ 7.47 116.83 $\downarrow$ 20.43 GPT-4o + Area Detection + Naive division 32.69 $\uparrow$ 7.12 102.41 $\downarrow$ 34.85 GPT-4o + Area detection 23.08 $\downarrow$ 2.49 120.06 $\downarrow$ 17.2 GPT-4o + Object-aware division (using SAM without GroundingDINO) 21.01 $\downarrow$ 4.56 135.04 $\downarrow$ 2.22 GPT-4o + Object-aware division 19.17 $\downarrow$ 6.40 120.61 $\downarrow$ 16.65 GPT-4o + Area detection + Object-aware division, (equiv.
+to LVLM-Count) 17.86 $\downarrow$ 7.71 91.71 $\downarrow$ 45.55
 Appendix B Visual Examples of LVLM-Count’s Performance on the FSC-147 Dataset, PASCAL VOC, and Emoji-Count Benchmarks
@@
 The LVLM used in the pipeline to generate these visual examples is GPT-4o.
-Figure 8 illustrates three examples of LVLM-Count’s performance on FSC-147.
-Additionally, Figure 9 includes an example from the PASCAL VOC benchmark.
-Finally, we present three visual examples from the Emoji-Count benchmark in Figure 10.
+Figure˜8 illustrates three examples of LVLM-Count’s performance on FSC-147.
+Additionally, Figure˜9 includes an example from the PASCAL VOC benchmark.
+Finally, we present three visual examples from the Emoji-Count benchmark in Figure˜10.
 In these visual examples, LVLM-Count achieves more accurate results compared to the base GPT-4o.
@@
 A polishing step further refines the masks by smoothing their surfaces and eliminating artifacts, preventing division lines from mistakenly passing through them.
-Figure 11 demonstrates the strong performance of our method on a randomly selected image from the web containing various types of flowers.
+Figure˜11 demonstrates the strong performance of our method on a randomly selected image from the web containing various types of flowers.
 This scene features extreme occlusion; nonetheless, our method effectively handles the division task.
 It is important to note that the strong performance of our method in this challenging scenario is not coincidental.
-Figure 12 illustrates the impact of our post-processing operations.
+Figure˜12 illustrates the impact of our post-processing operations.
 The upper row displays the unprocessed SAM masks, while the bottom row shows the results after post-processing, which generates reliable paths for the division lines.
@@
 Figure 12: The effect of post-processing on the masks produced by SAM in making LVLM-Count robust to crowded scenes with heavy occlusions. The object of interest is “flower”.
-To further demonstrate the effectiveness of our method in crowded scenes, we present Table 6, which compares the performance of different LVLMs on FSC-147 dataset samples where the ground truth count is 100 or higher.
+To further demonstrate the effectiveness of our method in crowded scenes, we present Table˜6, which compares the performance of different LVLMs on FSC-147 dataset samples where the ground truth count is 100 or higher.
 The table contrasts their baseline performance with their performance when using our LVLM-Count method.
 The results show significant improvements when LVLM-Count is employed.
-Furthermore, Figure 13 presents three visual examples of LVLM-Count’s performance on crowded images from FSC-147.
+Furthermore, Figure˜13 presents three visual examples of LVLM-Count’s performance on crowded images from FSC-147.
+Table 6: Evaluation on samples with ground truth counts equal to or higher than 100 from the test set of the FSC-147 dataset. Showing the effectiveness of our method for crowded scenes.
 Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT-4o 89.22 - 291.34 - LVLM-Count (GPT-4o as LVLM) 57.87 $\downarrow 31.35$ 185.29 $\downarrow 106.05$ Gemma 3 27B 125.69 - 330.38 - LVLM-Count (Gemma 3 27B as LVLM) 71.68 $\downarrow 54.01$ 251.48 $\downarrow 78.9$ Qwen2 VL 72B AWQ 102.64 - 302.23 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 79.66 $\downarrow 22.98$ 289.47 $\downarrow 12.76$
-Table 6: Evaluation on samples with ground truth counts equal to or higher than 100 from the test set of the FSC-147 dataset. Showing the effectiveness of our method for crowded scenes.
 Figure 13: Visual examples of the performance of LVLM-Count on three crowded scenes from the FSC-147 dataset.
@@
 In this variant, masks are obtained by configuring SAM to segment any entity in the image.
-The results presented in Table 4 show that both variants significantly improve the counting performance of the LVLMs, demonstrating the robustness of the initial stages in our pipeline.
-Figure 14 demonstrates the segmentation masks and division paths for the two variants on a sample from the Penguin benchmark.
-In Figure 14 (a), GroundingDINO with a low detection threshold is employed.
+The results presented in Table˜4 show that both variants significantly improve the counting performance of the LVLMs, demonstrating the robustness of the initial stages in our pipeline.
+Figure˜14 demonstrates the segmentation masks and division paths for the two variants on a sample from the Penguin benchmark.
+In Figure˜14 (a), GroundingDINO with a low detection threshold is employed.
 It correctly identifies all penguin instances, though there are some false positives.
 These detections are subsequently segmented by SAM.
-Figure 14 (c) depicts the division paths found using these masks.
-In Figure 14 (b), GroundingDINO is not used, and SAM is configured to segment all entities in the scene.
+Figure˜14 (c) depicts the division paths found using these masks.
+In Figure˜14 (b), GroundingDINO is not used, and SAM is configured to segment all entities in the scene.
 As a result, in addition to the penguins, a large number of other objects are segmented.
-Nevertheless, Figure 14 (d) shows that the division paths derived from these masks successfully partition the image without cutting any penguins.
+Nevertheless, Figure˜14 (d) shows that the division paths derived from these masks successfully partition the image without cutting any penguins.
 Note that some excessively large masks generated by SAM have been removed due to post-processing in our pipeline.
@@
 Appendix E Real-world Application of LVLM-Count
-As stated in Section 1, counting has numerous real-world applications, including but not limited to biology, health, industry, warehousing, and environmental monitoring.
+As stated in Section˜1, counting has numerous real-world applications, including but not limited to biology, health, industry, warehousing, and environmental monitoring.
 Below, we demonstrate the performance of LVLM-Count on examples from the following areas: i) biology/health, ii) industry/warehousing, and iii) environmental monitoring.
@@
 Note that in all examples, the cluster-based approach automatically determines the start and end points of the division paths.
-In Figure 15, images of two laboratory samples are analyzed using LVLM-Count.
+In Figure˜15, images of two laboratory samples are analyzed using LVLM-Count.
 The first row shows an image from a dataset introduced by Lempitsky and Zisserman [2010], which contains simulated bacterial cells from fluorescence-light microscopy, created by Lehmussola et al.
@@
 The standard staining procedure highlights the nuclei of various cell types in blue, while other cellular components appear in shades of pink and red [Paul Cohen et al., 2017].
-As observed, LVLM-Count achieves much higher accuracy in counting bacterial cells and bone marrow nuclei in the top and bottom rows of Figure 15, respectively, compared to the base LVLM, particularly for the bone marrow nuclei.
-In Figure 16, two images from industrial scenes are analyzed using LVLM-Count.
+As observed, LVLM-Count achieves much higher accuracy in counting bacterial cells and bone marrow nuclei in the top and bottom rows of Figure˜15, respectively, compared to the base LVLM, particularly for the bone marrow nuclei.
+In Figure˜16, two images from industrial scenes are analyzed using LVLM-Count.
 The top row shows a sectional image of a stockpile of tree logs, and the bottom row shows an image from an industrial area containing barrels of various colors.
@@
 In both cases, LVLM-Count’s predictions are significantly closer to the ground truth values than those of the base LVLM, particularly for the tree logs, where the ground truth number is too large for the base LVLM to estimate accurately.
-Figure 17 shows an image sourced from a dataset [Penguin Research, 2016] created as part of an ongoing initiative to monitor the penguin population in Antarctica.
+Figure˜17 shows an image sourced from a dataset [Penguin Research, 2016] created as part of an ongoing initiative to monitor the penguin population in Antarctica.
 This dataset comprises images captured hourly by a network of fixed cameras installed at over 40 locations.
@@
 In this section, we demonstrate how LVLM-Count performs in counting different objects of interest, determined simply by a prompt, using an image with multiple object categories.
-The image in Figure 18 contains three object categories: person, cow, and horse.
+The image in Figure˜18 contains three object categories: person, cow, and horse.
 In the top row, the object of interest is “cow.” We prompt LVLM-Count to count the cows.
@@
 We observe that GPT-4o successfully counts the number of cows in the resulting subimages, leading to the correct final answer.
-In the middle row of Figure 18, the object of interest is “person.” LVLM-Count again successfully counts the number of people accurately.
-A more interesting case is the bottom row of Figure 18, where both cows and persons are objects of interest.
+In the middle row of Figure˜18, the object of interest is “person.” LVLM-Count again successfully counts the number of people accurately.
+A more interesting case is the bottom row of Figure˜18, where both cows and persons are objects of interest.
 We prompt LVLM-Count to count the number of “cows and persons.” Similar to the first row of the figure, there are false positive masks here as well.
@@
 Thus, it often returns bounding boxes for all instances of the object category mentioned in the expression, even if those instances do not satisfy other conditions in the expression.
-For example, in the top row of Figure 19, $E=$brown egg”.
+For example, in the top row of Figure˜19, E=E=brown egg”.
 However, all the eggs have been segmented regardless of their color.
@@
 Appendix H Performance Analysis of LVLM-Count for different ground truth ranges on FSC-147 dataset
-To further investigate the performance of our pipeline, we divide the ground truth values in the FSC-147 test set into intervals and plot the MAE for the base GPT-4o, Qwen2, and Gemma 3 models alongside the results from LVLM-Count using each model, as shown in Figure 20.
+To further investigate the performance of our pipeline, we divide the ground truth values in the FSC-147 test set into intervals and plot the MAE for the base GPT-4o, Qwen2, and Gemma 3 models alongside the results from LVLM-Count using each model, as shown in Figure˜20.
 The first interval contains relatively small ground truth values, a range in which LVLMs already perform well.
 As the ground truth values increase, the base models exhibit increasingly larger errors compared to LVLM-Count, with the margin growing rapidly.
-This behavior is consistent with our observations of counting errors on the blue circles in Figure 2.
+This behavior is consistent with our observations of counting errors on the blue circles in Figure˜2.
 (a) LVLM-Count using GPT-4o
@@
 (c) LVLM-Count using Qwen2
-Figure 20: Performance analysis of our method, LVLM-Count, on the FSC-147 test set using GPT-4o (Figure 20(a)), Gemma 3 (Figure 20(b)), and Qwen2 (Figure 20(c)).
+Figure 20: Performance analysis of our method, LVLM-Count, on the FSC-147 test set using GPT-4o (Figure˜20(a)), Gemma 3 (Figure˜20(b)), and Qwen2 (Figure˜20(c)).
 In the first interval, all base LVLMs exhibit a lower MAE.
@@
 This section presents various accuracy measures for the experiments reported in Tables 1, 2, 3, and 4.
-The accuracy metrics are defined in Table 7.
-Following these definitions, Table 8 shows the measured accuracies for the FSC-147 test set.
+The accuracy metrics are defined in Table˜7.
+Following these definitions, Table˜8 shows the measured accuracies for the FSC-147 test set.
 Similarly, Tables 9, 10, and 11 present the accuracy metrics for the PASCAL VOC, Emoji-Count, and Penguin benchmarks, respectively.
 The results in the tables demonstrate that LVLM-Count generally achieves a higher accuracy than the base LVLM it employs.
-Metric Definition Acc Percentage of answers such that $answer=GT$ Acc±$\pm 1$ Percentage of answers such that $|answer-GT|\leq 1$ Acc±$\pm 3$ Percentage of answers such that $|answer-GT|\leq 3$ Acc±$\pm 5$ Percentage of answers such that $|answer-GT|\leq 5$ Acc±$\pm 10$ Percentage of answers such that $|answer-GT|\leq 10$
 Table 7: Definitions of Various Accuracy Metrics. GT denotes the ground truth number.
-Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) GPT4o 12.24 26.22 42.10 52.41 66.58 LVLM-Count (GPT4o as LVLM) 14.45 ($\uparrow$) 28.99 ($\uparrow$) 47.25 ($\uparrow$) 57.03 ($\uparrow$) 70.45 ($\uparrow$) Gemma 3 27B 8.71 18.94 32.55 41.57 55.15 LVLM-Count (Gemma 3 as LVLM) 11.37 ($\uparrow$) 25.46 ($\uparrow$) 41.51 ($\uparrow$) 52.07 ($\uparrow$) 67.09 ($\uparrow$) Qwen2 VL 72B AWQ 9.19 20.81 36.83 46.95 62.07 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 9.61 ($\uparrow$) 23.95 ($\uparrow$) 43.05 ($\uparrow$) 53.78 ($\uparrow$) 69.08 ($\uparrow$)
+[…]
 Table 8: FSC-147 Dataset.
 A ($\uparrow$) next to the measured accuracies for LVLM-Count indicates improvement over the base LVLM it uses, while a ($\downarrow$) indicates degradation compared to the corresponding base LVLM.
-Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) GPT4o 30.39 47.06 61.11 71.90 89.87 LVLM-Count (GPT4o as LVLM) 32.35 ($\uparrow$) 51.96 ($\uparrow$) 73.20 ($\uparrow$) 82.35 ($\uparrow$) 93.46 ($\uparrow$) Gemma 3 27B 29.41 54.58 75.49 85.62 92.16 LVLM-Count (Gemma 3 27B as LVLM) 30.39 ($\uparrow$) 50.98 ($\downarrow$) 78.43 ($\uparrow$) 85.62 (-) 94.77 ($\uparrow$) Qwen2 VL 72B AWQ 24.18 42.48 60.13 72.22 86.60 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 28.76 ($\uparrow$) 46.73 ($\uparrow$) 67.32 ($\uparrow$) 78.43 ($\uparrow$) 89.87 ($\uparrow$)
+[…]
 Table 9: PASCAL VOC Benchmark.
 A ($\uparrow$) next to the measured accuracies for LVLM-Count indicates improvement over the base LVLM it uses, while a ($\downarrow$) indicates degradation compared to the corresponding base LVLM.
-Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) GPT4o 1.85 5.54 13.73 21.12 43.94 LVLM-Count (GPT4o as LVLM) 2.81 ($\uparrow$) 9.88 ($\uparrow$) 23.53 ($\uparrow$) 36.22 ($\uparrow$) 60.24 ($\uparrow$) Gemma 3 27B 0.80 1.77 4.10 7.07 16.14 LVLM-Count (Gemma 3 27B as LVLM) 2.49 ($\uparrow$) 5.30 ($\uparrow$) 13.33 ($\uparrow$) 19.92 ($\uparrow$) 38.63 ($\uparrow$) Qwen2 VL 72B AWQ 0.88 2.89 7.55 11.41 19.76 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 2.33 ($\uparrow$) 6.75 ($\uparrow$) 16.22 ($\uparrow$) 25.70 ($\uparrow$) 43.86 ($\uparrow$)
+[…]
 Table 10: Emoji-Count Benchmark.
 A ($\uparrow$) next to the measured accuracies for LVLM-Count indicates improvement over the base LVLM it uses, while a ($\downarrow$) indicates degradation compared to the corresponding base LVLM.
-Method Acc (%) Acc±$\pm$1 (%) Acc±$\pm$3 (%) Acc±$\pm$5 (%) Acc±$\pm$10 (%) GPT4o 2.00 3.67 8.33 12.00 22.00 LVLM-Count (GPT4o as LVLM) 2.67 ($\uparrow$) 6.33 ($\uparrow$) 14.33 ($\uparrow$) 18.67 ($\uparrow$) 33.00 ($\uparrow$) Gemma 3 27B 1.33 1.67 7.00 8.67 16.00 LVLM-Count (Gemma 3 27B as LVLM) 1.00 ($\downarrow$) 3.67 ($\uparrow$) 9.33 ($\uparrow$) 12.67 ($\uparrow$) 23.67 ($\uparrow$) Qwen2 VL 72B AWQ 1.67 2.00 6.33 10.33 16.67 LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 1.67 (-) 5.33 ($\uparrow$) 11.67 ($\uparrow$) 18.00 ($\uparrow$) 31.00 ($\uparrow$)
+[…]
 Table 11: Penguin Benchmark.
 A ($\uparrow$) next to the measured accuracies for LVLM-Count indicates improvement over the base LVLM it uses, while a ($\downarrow$) indicates degradation compared to the corresponding base LVLM.
+[…]
 Appendix J Inference Time Comparison
 LVLM-Count is a pipeline consisting of multiple stages.
-However, in the following section, we will demonstrate that the dominant portion of time during inference with a GPT-4o model is spent querying the LVLM to count objects, while other stages are negligible by comparison.
-Consequently, LVLM-Count can achieve inference times reasonably close to that of the base LVLM used.
+However, in this section, we demonstrate that the dominant portion of the time during inference with the LVLMs tested in this paper, i.e.
+GPT4o, Gemma 3, and Qwen2, is spent querying the LVLM to count objects, and the additional steps in the pipeline comprise a small portion in comparison.
 In the four stages of the pipeline, the following main steps are taken: making a call to an LLM/LVLM to extract the object of interest from $Q$, calling GroundingDINO for area detection, calling GroundingDINO and SAM for target segmentation, performing some post-processing on the masks, running the mean-shift clustering algorithm, running the $A^{*}$ search, and querying the LVLM for the sub-images.
 From the listed steps, the running time of the post-processing on masks is negligible.
-The time required for the rest of the steps in an optimized pipeline is demonstrated in Table 12.
+The time required for the rest of the steps in an optimized pipeline is demonstrated in Table˜12.
 To obtain the inference time for each step, we evaluated it on 1,000 images from the FSC-147 dataset and calculated the average.
 footnotetext: This is for a case where sub-image queries to the LVLM are made in parallel
-Step Extract $E$ from $Q$ Area Detection Target Segmentation Clustering $A^{*}$ Sub-image Analysis Total Model Called GPT-4o GroundingDINO GroundingDINO SAM - - GPT-4o Time 0.56 0.10 0.10 0.33 0.24 0.03 1.65J 3.03
-Table 12: Inference time breakdown of the LVLM-Count pipeline (Time in seconds)
-As shown in Table 12, the dominant term in the inference time comes from calling GPT-4o to count objects of interest in the sub-images, which occupies more than half of the total inference time.
+Table 12: Inference time breakdown of the LVLM-Count pipeline for different LVLMs (Time in seconds)
+Method Extract $E$ from $Q$ Area Detection Target Segmentation Clustering $A^{*}$ Subimage AnalysisJ Total LVLM-Count (GPT4o) 0.11 0.10 0.43 0.24 0.03 1.65 2.56 LVLM-Count (Gemma 3 27B) 0.11 1.11 2.02 LVLM-Count (Qwen2 VL 72B AWQ) 0.10 1.17 2.07
+As shown in Table˜12, the dominant term in the inference time comes from calling the LVLM to count objects of interest in the sub-images.
+This occupies more than half of the total inference time.
 However, a similar step occurs when querying the base LVLM to count objects of interest in the main image.
 The additional steps present in LVLM-Count but absent in the base LVLM constitute a smaller portion of the time.
-Table 13 presents a comparison between the total inference time of LVLM-Count and the base GPT-4o.
+Table˜13 presents a comparison between the total inference time of LVLM-Count and the base LVLMs.
 This increase in inference time may be acceptable for most applications, given the substantial improvements in counting accuracy when using LVLM-Count.
+Table 13: Comparison of the inference time between the base LVLM and LVLM-Count
 Method Inference (s)
 Base GPT4o 1.68
-LVLM-Count (Using GPT-4o) 3.03
-Table 13: Comparison of the inference time between the base GPT-4o and LVLM-Count
+LVLM-Count (Using GPT-4o) 2.56
+Base Gemma 3 27B 1.52
+LVLM-Count (Using Gemma 3 27B) 2.01
+Base Qwen2 VL 72B AWQ 1.46
+LVLM-Count (Using Qwen2 VL 72B AWQ) 2.07
 Appendix K Comparison of Base LVLMs and LVLM-Count’s Counting Performance with SOTA Counting Models as a Reference Point
@@
 Reporting these results provides additional insights into the progress of counting performance in LVLMs and their enhanced performance with LVLM-Count.
-Table 14 presents the results of dedicated counting models alongside LVLMs and their corresponding LVLM-Count variants on the FSC-147 test set.
+Table˜14 presents the results of dedicated counting models alongside LVLMs and their corresponding LVLM-Count variants on the FSC-147 test set.
 Note that GroundingREC, CountGD, and $\text{DAVE}_{\text{prm}}$ are the top three highest-performing text-based trained models on FSC-147 in the literature.
@@
 However, the use of LVLM-Count significantly narrows this performance gap.
-Method Trained Model MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC [Shi et al., 2024] 24.79 - 137.15 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 14.90 - 103.42 - CountGD [Amini-Naieni et al., 2024] 14.76 - 120.42 - GroundingREC [Dai et al., 2024] 10.12 - 107.19 - GPT-4o 25.17 - 137.86 - LVLM-Count (GPT-4o as LVLM) 18.09 $\downarrow 7.08$ 88.77 $\downarrow 49.09$ Gemma 3 27B 30.23 - 136.43 - LVLM-Count (Gemma 3 27B as LVLM) 19.70 $\downarrow 10.53$ 110.62 $\downarrow 25.81$ Qwen2 VL 72B AWQ 33.45 - 145.90 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 22.67 $\downarrow 10.78$ 124.73 $\downarrow 21.17$
 Table 14: Evaluation of SOTA counting models on the FSC-147 test set serves as a reference point for assessing the performance of LVLMs and their corresponding LVLM-Count variants.
@@
 Improvements are shown in green, while degradations are shown in red.
-Results on the PASCAL VOC dataset are reported in Table 15.
+Method Trained Model MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC [Shi et al., 2024] 24.79 - 137.15 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 14.90 - 103.42 - CountGD [Amini-Naieni et al., 2024] 14.76 - 120.42 - GroundingREC [Dai et al., 2024] 10.12 - 107.19 - GPT-4o 25.57 - 137.26 - LVLM-Count (GPT-4o as LVLM) 17.86 $\downarrow 7.71$ 91.71 $\downarrow 45.55$ Gemma 3 27B 30.59 - 132.61 - LVLM-Count (Gemma 3 27B as LVLM) 20.25 $\downarrow 10.34$ 110.45 $\downarrow 22.16$ Qwen2 VL 72B AWQ 34.48 - 149.49 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 22.29 $\downarrow 11.89$ 119.46 $\downarrow 30.44$
+Results on the PASCAL VOC dataset are reported in Table˜15.
 For training-based counting models, this is considered a cross-dataset evaluation, meaning that they are trained on the training set of FSC-147 and tested on the PASCAL VOC benchmark.
@@
 Nevertheless, LVLM-Count is very effective in narrowing the performance gap.
-Additionally, the results on the Emoji-Count benchmark are reported in Table 16.
+Additionally, the results on the Emoji-Count benchmark are reported in Table˜16.
 This is also a cross-dataset evaluation.
@@
 Employing LVLM-Count widens this gap even further.
-Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TrainingFree [Shi et al., 2024] 12.03 - 18.18 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 12.39 - 22.81 - CountGD [Amini-Naieni et al., 2024] 2.81 - 7.01 - GroundingRec [Dai et al., 2024] 4.05 - 7.80 - GPT4o 4.46 - 8.35 - LVLM-Count (GPT4o as LVLM) 3.31 $\downarrow$ 1.15 7.11 $\downarrow$ 1.24 Gemma 3 27B 3.37 - 7.49 - LVLM-Count (Gemma 3 27B as LVLM) 2.87 $\downarrow$ 0.5 5.84 $\downarrow$ 1.65 Qwen2 VL 72B AWQ 4.83 - 8.84 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 4.05 $\downarrow$ 0.78 7.72 $\downarrow$ 1.12
 Table 15: Evaluation of SOTA counting models on the PASCAL VOC counting benchmark as a reference point for the performance of LVLMs and their corresponding LVLM-Count performance.
-Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC [Shi et al., 2024] 64.64 - 87.45 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 198.99 - 208.08 - CountGD [Amini-Naieni et al., 2024] 137.93 - 156.80 - GroundingREC [Dai et al., 2024] 143.22 - 158.74 - GPT-4o 22.51 - 35.94 - LVLM-Count (GPT-4o as LVLM) 16.29 $\downarrow 6.22$ 32.47 $\downarrow 3.47$ Gemma 3 27B 21.43 - 24.15 - LVLM-Count (Gemma3 27B as LVLM) 15.76 $\downarrow 5.67$ 21.00 $\downarrow 3.15$ Qwen2 VL 72B AWQ 82.41 - 186.32 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 23.74 $\downarrow 58.67$ 40.04 $\downarrow 146.28$
+Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TrainingFree [Shi et al., 2024] 12.03 - 18.18 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 12.39 - 22.81 - CountGD [Amini-Naieni et al., 2024] 2.81 - 7.01 - GroundingRec [Dai et al., 2024] 4.05 - 7.80 - GPT4o 4.64 - 8.56 - LVLM-Count (GPT4o as LVLM) 3.42 $\downarrow$ 1.22 7.17 $\downarrow$ 1.39 Gemma 3 27B 3.40 - 7.54 - LVLM-Count (Gemma 3 27B as LVLM) 2.97 $\downarrow$ 0.0.43 6.16 $\downarrow$ 1.38 Qwen2 VL 72B AWQ 4.80 - 8.71 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 4.12 $\downarrow$ 0.68 7.76 $\downarrow$ 0.95
 Table 16: Evaluation of SOTA counting models on the Emoji-Count benchmark, serving as a reference point for the performance of LVLMs and their corresponding LVLM-Count performance.
-Finally, Table 17 presents the results of prior counting models on the Penguin benchmark alongside the counting performance of LVLMs.
+Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC [Shi et al., 2024] 64.64 - 87.45 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 198.99 - 208.08 - CountGD [Amini-Naieni et al., 2024] 137.93 - 156.80 - GroundingREC [Dai et al., 2024] 143.22 - 158.74 - GPT-4o 23.57 - 36.97 - LVLM-Count (GPT-4o as LVLM) 16.57 $\downarrow 7$ 33.11 $\downarrow 3.86$ Gemma 3 27B 21.39 - 24.04 - LVLM-Count (Gemma3 27B as LVLM) 16.16 $\downarrow 5.23$ 21.27 $\downarrow 2.77$ Qwen2 VL 72B AWQ 78.05 - 159.22 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 24.43 $\downarrow 53.62$ 43.38 $\downarrow 115.84$
+Finally, Table˜17 presents the results of prior counting models on the Penguin benchmark alongside the counting performance of LVLMs.
 We observe that prior training-based counting models significantly outperform LVLMs on this benchmark.
 However, similar to other datasets, LVLM-Count narrows this performance gap.
-Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC [Shi et al., 2024] 59.34 - 74.40 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 22.29 - 31.38 - CountGD [Amini-Naieni et al., 2024] 17.07 - 25.08 - GroundingREC [Dai et al., 2024] 22.04 - 26.75 - GPT4o 36.01 - 46.64 - LVLM-Count (GPT-4o as LVLM) 26.95 $\downarrow$ 9.06 39.32 $\downarrow$ 7.32 Gemma 3 27B 49.87 - 61.06 - LVLM-Count (Gemma 3 27B as LVLM) 36.75 $\downarrow$ 13.12 47.24 $\downarrow$ 13.82 Qwen2 VL 72B AWQ 44.96 - 66.78 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 28.67 $\downarrow$ 16.29 43.71 $\downarrow$ 23.07
 Table 17: Evaluation of SOTA counting models on the Penguin benchmark, serving as a reference point for the performance of LVLMs and their corresponding LVLM-Count performance.
+Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC [Shi et al., 2024] 59.34 - 74.40 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 22.29 - 31.38 - CountGD [Amini-Naieni et al., 2024] 17.07 - 25.08 - GroundingREC [Dai et al., 2024] 22.04 - 26.75 - GPT4o 35.18 - 45.76 - LVLM-Count (GPT-4o as LVLM) 26.76 $\downarrow$ 8.42 38.60 $\downarrow$ 7.16 Gemma 3 27B 49.44 - 60.20 - LVLM-Count (Gemma 3 27B as LVLM) 34.13 $\downarrow$ 15.31 44.11 $\downarrow$ 16.09 Qwen2 VL 72B AWQ 44.02 - 65.59 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 28.21 $\downarrow$ 15.81 40.40 $\downarrow$ 25.19
 Appendix L LVLM-Count’s Ability to Handle Visual Exemplars
@@
 The adaptation is straightforward, as most LVLMs also accept visual prompts.
-In Table 18, we evaluate the base GPT-4o and LVLM-Count (which uses GPT-4o) on Emoji-Count with text-only prompts, visual exemplars, and combinations of text and visual exemplars.
+In Table˜18, we evaluate the base GPT-4o and LVLM-Count (which uses GPT-4o) on Emoji-Count with text-only prompts, visual exemplars, and combinations of text and visual exemplars.
 For the visual-exemplar version, a single image of the target emoji icon is provided, and the LVLM is required to count similar instances.
@@
 The results show higher accuracy when combining text and visual exemplars, while the text-only and exemplar-only versions achieve approximately the same performance.
+Table 18: Evaluation of LVLM-Count’s performance using text-only inputs, exemplar-only inputs, and a combination of text and exemplars on the Emoji-Count dataset.
 Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT4o 22.51 - 35.94 - LVLM-Count (GPT4o as LVLM, text) 16.29 $\downarrow$ 6.22 32.47 $\downarrow$ 3.47 LVLM-Count (GPT4o as LVLM, visual exemplars) 15.66 $\downarrow$ 6.85 31.69 $\downarrow$ 4.25 LVLM-Count (GPT4o as LVLM, text + visual exemplars) 13.46 $\downarrow$ 9.05 24.35 $\downarrow$ 11.59
-Table 18: Evaluation of LVLM-Count’s performance using text-only inputs, exemplar-only inputs, and a combination of text and exemplars on the Emoji-Count dataset.
 Appendix M Chain of Thought Prompting for LVLM-Count
@@
 For our experiments, we appended the phrase "think step-by-step" to the counting prompt.
-As shown in Table 19, the results on the PASCAL VOC benchmark reveal that CoT negatively affects counting performance—not only for LVLM-Count but also for the base LVLM.
+As shown in Table˜19, the results on the PASCAL VOC benchmark reveal that CoT negatively affects counting performance—not only for LVLM-Count but also for the base LVLM.
+Table 19: Evaluation of CoT prompting for visual counting on PASCAL VOC.
 Method MAE $\downarrow$ RMSE $\downarrow$ GPT4o 4.46 8.35 GPT4o with CoT prompting 5.69 10.10 LVLM-Count (GPT4o as LVLM) 3.31 7.11 LVLM-Count with CoT prompting (GPT4o as LVLM) 3.89 7.70
-Table 19: Evaluation of CoT prompting for visual counting on PASCAL VOC.
 Appendix N Illustration of the Complete Workflow of the LVLM-Count for an Additional Image
 In this section, we demonstrate the same steps illustrated for the example image of eggs in Figures 3, 4, 5, and 6 for an image of zebras drinking water.
-The zebra image is passed to the pipeline along with the question $Q=$“how many zebras are in the image?”.
-First, $E=$“zebra” is extracted using the LLM.
+The zebra image is passed to the pipeline along with the question Q=Q=“how many zebras are in the image?”.
+First, E=E=“zebra” is extracted using the LLM.
 Then the zebra image is passed to the area detection stage, where the prompt given to GroundingDINO is “zebras”.
-The output bounding boxes are merged, and the resulting area is cropped, as illustrated in Figure 21.
+The output bounding boxes are merged, and the resulting area is cropped, as illustrated in Figure˜21.
 The cropped area is then passed to the target segmentation stage.
 At this stage, GroundingDINO detects the objects of interest defined by $E$ as the input prompt.
-SAM then uses the output bounding boxes to produce segmentation masks for the zebras, as shown in Figure 22.
+SAM then uses the output bounding boxes to produce segmentation masks for the zebras, as shown in Figure˜22.
 Figure 21: Illustration of the area detection step of LVLM-Count for the zebra image.
@@
 The goal is to produce all the instance masks for $E$ set to “zebra”.
-The cropped image from Figure 21, together with $E$, is given as input to GroundingDINO, which produces the output shown in Figure 22(a).
-Figure 22(a) is then given as input to SAM, which produces the output shown in Figure 22(b).
+The cropped image from Figure˜21, together with $E$, is given as input to GroundingDINO, which produces the output shown in Figure˜22(a).
+Figure˜22(a) is then given as input to SAM, which produces the output shown in Figure˜22(b).
 After the target segmentation stage, the masks are passed to the object-aware division stage.
@@
 The division paths are then found by connecting each start point to its corresponding end point by running the $A^{*}$ search algorithm on the graph.
-The found paths are mapped back into the image domain and drawn in red, as depicted in Figure 23.
+The found paths are mapped back into the image domain and drawn in red, as depicted in Figure˜23.
 The image contours are determined based on the drawn red paths, and each contour’s interior is masked out independently to obtain the subimages.
@@
 Right: Illustration of object-aware division.
-The masks from Figure 22(b) are turned into a black-and-white image.
+The masks from Figure˜22(b) are turned into a black-and-white image.
 A dividing path is found by connecting $P_{s}$ to $P_{e}$ using the $A^{*}$ search algorithm in a graph that corresponds to the binary image, where the only nodes in the graph are white pixels, which are connected to all of their white neighboring pixels.
 The path is mapped back to the pixel domain.
-Appendix O Bias to Small Numbers in Datasets and Counting Performance of LVLMs
-In our experiments, LVLMs are able to make correct predictions when the number of items to be counted is small, but errors increase as the ground truth number grows.
-Although, we cannot be certain, one likely reason for this behavior in LVLMs is that, during training, the counting questions these models encounter are heavily biased toward small numbers.
-As an example, in Figure 24 we show the distribution of ’How many’ questions in some well-known VQA datasets.
-(a) TallyQA Dataset [Acharya et al., 2019]
-(b) VQAv2 Dataset [Goyal et al., 2017]
-Figure 24: Distribution of answers in two VQA datasets.
+Appendix O Emoji-Count Benchmark Details
+In this section, we provide further details about the Emoji-Count benchmark.
+The full benchmark, including all images and corresponding annotations, is publicly available in our GitHub repository.
+The emoji icons used in daily chats can be placed in classes closely related to a specific topic.
+While the icons within these groups often appear very similar, they possess subtle yet important differences.
+Identifying a specific emoji among others based on these fine distinctions requires a complex level of understanding from the counting model.
+Our dataset was built from the 1,816 standard Apple emojis available at the time of its creation.
+Each emoji has a unique descriptive name.
+We first grouped these emojis into 82 non-overlapping classes based on their name’s topic.
+To prevent ambiguity for a counting model, we then removed confusing icons, such as those with names that could also refer to a more specific emoji in the class.
+This cleaning process left us with 1,197 icons.
+From each class, we randomly selected six icons for inclusion; classes with fewer than six retained all their icons.
+Figure˜25(b) shows the six selected icons from two example classes.
+To generate the benchmark images, we began with an empty 1024 × 1024 pixel white canvas for each class.
+We then populated each canvas with instances of the icons from that class.
+Note that each icon may also be referred to as an object category.
+The procedure was as follows: for each of the six icons in a class, we generated a random number between 30 and 50.
+A number of instances equal to this random value were then placed at random positions on the canvas, ensuring that the icons did not overlap and remained completely within the canvas borders.
+Each image is annotated with the count for every icon type it contains.
+The benchmark comprises 415 questions in total.
+Table˜20 provides additional statistics for the benchmark, while Figure˜25(a) shows the distributions of the ground truth counts.
+Table 20: Statistics of the Emoji-Count benchmark dataset.
+Metric Value Total Questions 415 Total Classes 82 Total Images 82 Total Categories 415 Average Questions per Image 5.06 Answer Range 30-50 Average Answer Value $39.87\pm 6.33$
+Figure 24: This pie chart illustrates the 82 classes in the Emoji-Count benchmark.
+Inside each class there are 3 to 6 type of emoji icons that are related to the topic of the class.
+For example, the emoji icons inside the plants and hands classes, have been demonstrated in Figure˜25(b)
+(a) Emoji-Count ground truth count distribution
+(b) Two class examples from Figure˜24
+Figure 25: The figure illustrates: (a) the distribution of ground truth counts in the images of Emoji-Count benchmark, and (b) the object categories of two example classes, plants and hands
+Appendix P Parsing Strategy for the Output of Open-source LVLMs
+Unlike GPT-4o, which provides the user with the ability to format the output answer using a JSON schema, the open-source models used in this work, i.e., Qwen2 VL 72B AWQ and Gemma 3 27B, do not have such an option.
+When we prompt these models to count the number of instances of an object of interest, the answer is often accompanied by descriptive text that does not follow a specific format, making the extraction of the predicted object count difficult.
+To solve this issue, we include a sentence in the counting prompt given to the LVLM that asks the model to put its final prediction inside double brackets.
+The template prompt is: “How many object-of-interest are in the image?
+Report your final answer with only one number inside [[Double Brackets]]”.
+After the model generates an answer, we use Python’s regular expression package to extract the predicted number from within the double brackets.
+Often, this number is a word rather than a numeral.
+Therefore, after extraction, we use Python’s text2num package to convert the word into its numerical form.
+The following code snippet shows this number extraction process.
+Details unrelated to the regular expression have been omitted.
+1import re
+2from text_to_num import alpha2digit
+3
+4text_prompt = ("How many object-of-interest are in the image. "
+5 "Report your final answer with only one number inside [[Double Brackets]]")
+6while True:
+7 try:
+8 output_text = LVLM( text_prompt, image)
+9 answer = re.search("\[\[(.+)\]\]", output_text[0]).group(1)
+10 count = alpha2digit(answer, "en")
+11 count = int(count)
+12 break
+13 except Exception:
+14 pass
+Despite the format instruction in the prompt, the models might occasionally fail to produce an answer in the required format.
+This results in an error, which is handled by the code above.
+The model is then asked to generate a new output until a successful prediction is extracted.
+Let us refer to an unsuccessful attempt to extract the predicted number as a failure.
+We measure the average failure rate per image and report it for each of the datasets in the main paper for the two open-source models that we use, i.e., Qwen2 VL 72B AWQ and Gemma 3 27B.
+These results are reported in Table˜21.
+Table 21: Average failure rate of extracting the predicted number of open-source models across different benchmarks
+Method Failure Rate
+FSC-147 PASCAL VOC Emoji-Count Penguin
+Base Gemma 3 27B $\approx 0$ 0 $\approx 0$ $\approx 0$
+LVLM-Count (Using Gemma 3 27B) $\approx 0$ 0 $\approx 0$ 0
+Base Qwen2 VL 72B AWQ 0.07 0.06 0.09 0.13
+LVLM-Count (Using Qwen2 VL 72B AWQ) 0.07 0.08 0.01 0.12
+Appendix Q Non-zero Prediction of LVLMs for Subimages with No Objects of Interest
+As mentioned, LVLMs occasionally predict non-zero values for images that contain no object of interest.
+In our pipeline, it is very rare for subimages generated from images that contain a non-zero number of objects of interest to not contain any objects of interest.
+The area detection stage ensures that areas containing objects of interest are cropped, and the clustering step in the object-aware division stage finds the division points so that the distribution of potential objects of interest in different subimages is balanced.
+However, we provide quantitative information on this phenomenon for our PASCAL VOC benchmark.
+In total, for only $9.80\%$ of the images in the benchmark, at least one of the subimages generated through our pipeline contains no instance of the queried objects.
+Half of these cases are images that have a zero ground truth count to begin with.
+Excluding those cases, the percentage drops to $4.90\%$.
+We report the average non-zero prediction rate for the subimages with a zero ground truth count and the average predicted values in Table˜22.
+For each LVLM, we report two cases.
+In one case, the input text is simply a counting prompt; in the other case, the input text includes the following sentence in addition to the counting prompt: “Say 0 if you do not see any.” For the sake of brevity in the table entries, let us refer to this simple technique as Zero-aware prompt.
+As observed in the table, this simple technique is quite effective in reducing the error.
+Note that the standard LVLM-Count uses this technique.
+Table 22: Average non-zero prediction rate and average predicted values for the subimages with zero ground truth count in PASCAL VOC benchmark.
+Method Average Non-zero Prediction Rate $\downarrow$ Average Predicted Value $\downarrow$ LVLM-Count (GPT4o w/o Zero-aware prompt) 0.39 0.78 LVLM-Count (GPT4o) 0.15 0.15 LVLM-Count (Gemma 3 27B w/o Zero-aware prompt) 0.45 2.24 LVLM-Count (Gemma 3 27B) 0.18 0.69 LVLM-Count (Qwen2 VL 72B AWQ w/o Zero-aware prompt) 0.42 0.51 LVLM-Count (Qwen2 VL 72B AWQ) 0.24 0.27
+Appendix R Reproducibility Table for Implementation Parameter
+This section provides Table˜23, which lists the values of key parameters used in our implementation.
+  • Detection threshold for area detection: This threshold controls the acceptance of bounding boxes proposed by GroundingDINO, based on the probability of each box during the area detection stage.
+  • Detection threshold for target segmentation: This parameter is defined similarly but applies to the target segmentation stage.
+  • Mask NMS IoU threshold: This sets the maximum acceptable Intersection over Union (IoU) value between two masks in the NMS algorithm.
+If two masks exceed this IoU, the one with the lower probability is removed.
+  • Mask erosion thickness: This determines the size of a square kernel (cv2.erode) used to erode the outer layer of masks.
+This prevents obstructions for division paths when two adjacent masks overlap.
+  • Mask refinement thickness: This defines the size of a square kernel (cv2.morphologyEx) used to refine the surface of potentially pitted masks.
+Table 23: Reproducibility table for implementation parameters
+Parameter Value Detection threshold for area detection 0.1 Detection threshold for target segmentation 0.1 Mask NMS IoU threshold 0.4 Mask erosion thickness 2 pixels Mask refinement thickness 3 pixels
+Appendix S Dividing Images Along Both the Horizontal and Vertical Axes
+As described in Section˜3.3, in the object-aware division stage of LVLM-Count, the divisions are vertical.
+This is to keep the pipeline simple.
+Nonetheless, it is possible to perform the same steps to obtain horizontal division paths.
+In Table˜24, we report the performance of a variant of LVLM-Count that leverages horizontal divisions in addition to vertical divisions on the FSC-147 dataset.
+Additionally, Figure˜26 shows a visual example of this variant’s performance on a sample from the FSC-147 dataset.
+Table 24: Evaluation of a version of LVLM-Count that leverages divisions along both vertical and horizontal axes on the FSC-147 dataset.
+Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ GPT-4o 24.48 - 125.38 - LVLM-Count (GPT-4o as LVLM, vertical & horizontal divisions) 17.67 $\downarrow 6.81$ 90.61 $\downarrow 34.47$ Gemma 3 27B 30.59 - 132.61 - LVLM-Count (Gemma 3 27B as LVLM, vertical & horizontal divisions) 18.44 $\downarrow 12.15$ 103.91 $\downarrow 28.7$ Qwen2 VL 72B AWQ 34.18 - 149.49 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM, vertical & horizontal divisions) 18.19 $\downarrow 15.99$ 104.12 $\downarrow 45.37$
+Figure 26: A visual example of the performance of LVLM-Count on a sample from FSC-147 when it leverages both horizontal and vertical divisions
+Appendix T Evaluation of the counting methods on TallyQA-Complex
+TallyQA [Acharya et al., 2019] is an open-world counting dataset that includes complex counting questions involving relationships between objects, attribute identification, reasoning, and more.
+It is a fairly large dataset, with the training set containing $249318$ questions and the test set having $22991$ simple and $22991$ complex counting questions.
+The number of objects in each image ranges from 0 to $15$.
+We randomly sampled $10$ questions per ground truth count from the complex counting questions in the test set.
+This resulted in $149$ complex counting questions in total, since some ground truth values have fewer than $10$ samples available.
+We refer to this benchmark as the TallyQA-Complex benchmark.
+We compare our method, using GPT-4o and Qwen2 VL 72B AWQ as the LVLM, against their corresponding base models.
+For reference, we also report the cross-domain performance of three of the best-performing trained counting models—GroundingREC, CountGD, and $\text{DAVE}_{\text{prm}}$—as well as the only prior training-free model, TFOC.
+The results are shown in Table˜25.
+Note that CountGD and $\text{DAVE}_{\text{prm}}$ use weights trained on FSC-147, while GroundingREC uses weights trained on the REC-8K dataset [Dai et al., 2024], which its authors introduced alongside the model.
+Our method improves the MAE over the base LVLMs.
+Moreover, note that SOTA counting models are outperformed by the base LVLM models, further confirming that LVLMs possess better generalization and complex reasoning capabilities for counting tasks.
+Table 25: Evaluation of the performance of LVLMs and their corresponding LVLM-Count as well as SOTA counting models on the TallyQA-Complex counting benchmark.
+Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ TFOC [Shi et al., 2024] 12.41 - 22.80 - $\text{DAVE}_{\text{prm}}$ [Pelhan et al., 2024] 28.36 - 57.56 - CountGD [Amini-Naieni et al., 2024] 9.78 - 17.21 - GroundingRec [Dai et al., 2024] 5.83 - 10.13 - GPT4o 2.60 - 4.74 - LVLM-Count (GPT4o as LVLM) 2.28 $\downarrow$ 0.32 4.18 $\downarrow$ 0.56 Qwen2 VL 72B AWQ 3.21 - 5.35 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 2.47 $\downarrow$ 0.74 4.35 $\downarrow$ 1.00